### v1

# The dimension and Bose distance of some BCH codes of length $\frac{q^{m}-1}{\lambda}$

Abstract
BCH codes are important error correction codes, widely utilized due to their robust algebraic structure, multi-error correcting capability, and efficient decoding algorithms. Despite their practical importance and extensive study, their parameters, including dimension, minimum distance and Bose distance, remain largely unknown in general. This paper addresses this challenge by investigating the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$ over the finite field $\mathbb{F}_{q}$, where $\lambda$ is a positive divisor of q1q-1. Specifically, for narrow-sense BCH codes of this length with $m\geq 4$, we derive explicit formulas for their dimension for designed distance $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}+1$. We also provide explicit formulas for their Bose distance in the range $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}$. These ranges for $\delta$ are notably larger than the previously known results for this class of BCH codes. Furthermore, we extend these findings to determine the dimension and Bose distance for certain non-narrow-sense BCH codes of the same length. Applying our results, we identify several BCH codes with good parameters.
October 2, 2025
I Introduction
Throughout this paper, let $q$ be a prime power and $\mathbb{F}_{q}$ be the finite field of order $q$. Let $\mathbb{F}_{q}^{n}$ denote the $n$-dimensional linear space over $\mathbb{F}_{q}.$ A code of length $n$ over $\mathbb{F}_{q}$ is defined as a nonempty subset of $\mathbb{F}_{q}^{n}$. In particular, an $[n,k,d]$ linear code $\mathcal{C}$ over $\mathbb{F}_{q}$ is defined as a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum distance $d$. A linear code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is said to be cyclic if $(c_{0},c_{1},\ldots,c_{n-1})\in\mathcal{C}$ implies that $(c_{n-1},c_{0},\ldots,c_{n-2})\in\mathcal{C}.$ Identify each vector $(c_{0},c_{1},\ldots,c_{n-1})\in\mathbb{F}_{q}^{n}$ with its polynomial representation
$c_{0}+c_{1}x+\cdots+c_{n-1}x^{n-1}\in\mathbb{F}_{q}[x]/(x^{n}-1),$
and each code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ with a subset of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ In this way, a code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is a cyclic code if and only if it is an ideal of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ Note that each ideal of $\mathbb{F}_{q}[x]/(x^{n}-1)$ is principal. Therefore, a cyclic code $\mathcal{C}\subseteq\mathbb{F}_{q}[x]/(x^{n}-1)$ can be generated by a monic polynomial $g(x)$, denoted as $\mathcal{C}=\left\langle g(x)\right\rangle.$ Moreover, the polynomial $g(x)$ is a divisor of $x^{n}-1$. The generator $g(x)$ is called the generator polynomial of $\mathcal{C}$, and $h(x)=(x^{n}-1)/{g(x)}$ is called the parity-check polynomial of $\mathcal{C}$.
Suppose that $n$ is an integer such that $\mathrm{gcd}(n,q)=1$. Denote $m=\mathrm{ord}_{n}(q)$, i.e., the smallest integer such that $q^{m}\equiv 1\ (\mathrm{mod}\ n).$ Let $\alpha$ be a primitive element of $\mathbb{F}_{q^{m}}$. Then $\beta=\alpha^{\frac{q^{m}-1}{n}}$ is a primitive $n$-th root of unity. This leads to the factorization $x^{n}-1=\prod\limits_{i=0}^{n-1}(x-\beta^{i}).$ For each integer $i\in[0,n-1]$, we denote by $m_{i}(x)$ the minimal polynomial of $\beta^{i}$ over $\mathbb{F}_{q}$. A cyclic code of length $n$ over $\mathbb{F}_{q}$ is called a BCH code with designed distance $\delta$ if its generator polynomial takes the form
$\mathrm{lcm}(m_{b}(x),m_{b+1}(x),\ldots,m_{b+\delta-2}(x))$
for some integers $b$ and $2\leq\delta\leq n$, where $\mathrm{lcm}$ denotes the least common multiple of the polynomials. We denote by $\mathcal{C}_{(q,n,\delta,b)}$ such a BCH code. If $b=1$, it is called a narrow-sense BCH code, simply denoted by $\mathcal{C}_{(q,n,\delta)}$. If $n=q^{m}-1$, then it is called a primitive BCH code. Note that $\mathcal{C}_{(q,n,\delta,b)}$ and $\mathcal{C}_{(q,n,\delta,^{{}^{\prime}}b)}$ may be the identical for distinct $\delta$ and $\delta^{{}^{\prime}}$. The Bose distance of $\mathcal{C}_{(q,n,\delta,b)}$, denoted by $d_{B}$ or $d_{B}(\mathcal{C}_{(q,n,\delta,b)})$, is the largest integer such that $\mathcal{C}_{(q,n,\delta,b)}=\mathcal{C}_{(q,n,d_{B},b)}$.
BCH codes were first independently discovered by Hocquenghem [16] and by Bose and Ray-Chaudhuri [1, 2]. They occupy a central place in coding theory due to their remarkable properties. First, they offer great flexibility in the choice of code parameters, enabling error correction capabilities to be tailored to specific applications. In addition, for block lengths up to a few hundred bits, many BCH codes are among the most powerful codes known for given length and dimension. In addition, efficient encoding and decoding algorithms have been developed, which make BCH codes highly practical for real-world applications.
Despite their widespread use and extensive study in the literature [4, 9, 6, 13, 15, 21, 20, 18, 19, 22, 23, 26, 29, 37, 3, 33, 12, 32, 34, 11, 31, 38], several open problems persist regarding BCH codes, especially concerning the precise determination of their dimension, minimum distance, and Bose distance. These parameters are crucial indicators of a BCH code’s performance. Specifically, a BCH code over $\mathbb{F}_{q}$ with dimension $k$ and minimum distance $d$ can transmit $k$ $q$-ary information symbols and correct up to $\left\lfloor\frac{d-1}{2}\right\rfloor$ $q$-ary symbol errors. Furthermore, the Bose distance $d_{B}$ provides a fundamental lower bound on the minimum distance, as established by the BCH bound [16, 2]. Notably, Charpin [3] conjectured that for a narrow-sense primitive BCH code, $d\leq d_{B}+4$. Hence, determining the Bose distance is also invaluable for a deeper understanding of BCH codes and their capabilities. However, as noted by Charpin [5] and Ding [10], the general determination of these parameters remains a challenging problem.
This paper is dedicated to the investigation of the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$, denoted by $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta,b)}$, where $\lambda$ is a positive divisor of q1q-1. This class of BCH codes includes primitive BCH codes when $\lambda=1$. To date, these fundamental parameters are precisely known only for limited cases. Most existing results focus on the case where $\lambda=1$ and $b=1$, which corresponds to narrow-sense primitive BCH codes. For a comprehensive overview of the parameters of such codes, readers are referred to [39, 27, 14, 28, 7, 8]. In contrast, for cases when $\lambda\neq 1$, the understanding is much more limited. Even for certain specific cases, such as $\lambda=2$ [24, 36, 40] and $\lambda=q-1$ [25, 22], the dimension and Bose distance are known for only a few designed distances. For general divisors $\lambda$ of q1q-1, Zhu et al. [40] determined the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$. Recently, Sun [30] determined the dimension and minimum distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for other specific designed distances. Readers may refer to [10] for an excellent survey on known results regarding the parameters of BCH codes.
A deep understanding of $q$-cyclotomic cosets, particularly their sizes and coset leaders, is crucial for determining the dimension and Bose distance of BCH codes. Indeed, the main challenge in determining these parameters of BCH codes often stems from the irregular distribution of coset leaders. Our previous work [39] explored the distribution of coset leaders modulo $q^{m}-1$ within the range $1,q^{\lfloor(2m-1)/3\rfloor+1}]$ and the sizes of corresponding $q$-cyclotomic cosets, enabling us to determine the Bose distance and dimension of narrow-sense primitive BCH codes $\mathcal{C}_{(q,q^{m}-1,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}$.
To extend these results to BCH codes of length $\frac{q^{m}-1}{\lambda}$, we need to investigate $q$-cyclotomic cosets modulo $\frac{q^{m}-1}{\lambda}$. A key and useful observation in this regard is that an integer $a$ is a coset leader modulo $\frac{q^{m}-1}{\lambda}$ if and only if $\lambda a$ is a coset leader modulo $q^{m}-1$, and the size of the $q$-cyclotomic coset modulo $\frac{q^{m}-1}{\lambda}$ of $a$ is equal to the size of the $q$-cyclotomic coset modulo $q^{m}-1$ of $\lambda a$. Consequently, the problem of finding coset leaders and sizes of $q$-cyclotomic coset modulo $\frac{q^{m}-1}{\lambda}$ can be reduced to identifying integers divisible by $\lambda$ that are coset leaders modulo $q^{m}-1$ and determining the sizes of their corresponding $q$-cyclotomic coset modulo $q^{m}-1$.
Building upon this crucial observation and our prior analysis of $q$-cyclotomic cosets modulo $q^{m}-1$, we successfully generalize the results on primitive BCH codes in [39] to BCH codes of length $\frac{q^{m}-1}{\lambda}$. Specifically, for any positive divisor $\lambda$ of q1q-1 and positive integer $m\geq 4$, this paper determines:
  • the dimension of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$;
  • the Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$.
It is important to note that the existing knowledge of the dimension for narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ only covers designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$ and some specific cases.
Our results significantly extend this range, as evidenced by the inequality $\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}\geq\frac{q^{\lfloor(m+1)/2\rfloor}-1}{\lambda}\cdot q^{\lceil(m-4)/6\rceil}$. This implies that the range of $\delta$ for which we provide the dimension of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ is substantially larger than previously established. Additionally, we extend these results to some non-narrow-sense BCH codes of length $\frac{q^{m}-1}{\lambda}$. As illustrations of our main results, we also provide some explicit formulas determining the dimension and Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $\lambda\delta=aq^{h+k}+b$, where $h=\lfloor m/2\rfloor$, $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$, and $q\nmid b$.
This paper is organized as follows. In Sections II and III, we provide essential preliminaries and review our previous results concerning $q$-cyclotomic cosets modulo $q^{m}-1$. Section IV presents several auxiliary lemmas that will be employed in subsequent sections. Based on the theoretical foundations established in Sections II through IV, we then determine the dimension of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$ in Section V and the Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta,b)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ in Section VI, both by providing explicit formulas. Utilizing these formulas, we present some examples of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ and compare them with the tables of the best known linear codes maintained by Markus Grassl at http://www.codetables.de, which is called Database later in this paper. Furthermore, as an illustration of our main results, Section VII applies these formulas to compute the dimension and Bose distance of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ specifically when $\lambda\delta$ takes the form $aq^{h+k}+b$. Following this, Section VIII extends our analysis to determine the dimension and Bose distance of certain non-narrow-sense BCH codes. Moreover, we identify some non-narrow-sense BCH codes that possess optimal parameters. Finally, Section IX concludes the paper.
II Preliminaries
Let $n$ be an integer such that $\gcd(n,q)=1$. For each integer $a\in[0,n-1]$, the $q$-cyclotomic coset of $a$ modulo $n$ is defined as
$C_{n}(a)=\{aq^{i}\bmod n\mid i=0,\ldots,\ell_{a}-1\},$ (1)
where $\ell_{a}$ is the smallest positive integer such that $aq^{\ell_{a}}\equiv a\pmod{n}$. It is clear that $|C_{n}(a)|=\ell_{a}$, where $|\cdot|$ denotes the size of a set. Moreover, $|C_{a}|$ is a divisor of $m=\mathrm{ord}_{n}(q)$. The smallest integer in the coset $C_{n}(a)$ is called the coset leader of $C_{n}(a)$. For convenience, we occasionally refer to it simply as a coset leader modulo $n$. Let $\ell$ be a positive integer, and let $c,d\in[0,n-1]$ be two integers such that $c\leq d$. We define
$\mathcal{L}^{n}(c,d)=\{a\in[c,d]:a\text{ is the coset leader of }C_{n}(a)\}$
and
[…]
In particular, when (c,d)=(0,n1)(c,d)=(0,n-1), we simply denote $\mathcal{L}_{\ell}(c,d)$ by $\mathcal{L}_{\ell}$.
By [17, Theorem 4.1.1], the minimal polynomial $m_{a}(x)$ of $\beta^{a}$ over $\mathbb{F}_{q}$ can be given by
$m_{a}(x)=\prod\limits_{i\in C_{n}(a)}(x-\beta^{i}).$
Hence, the generator polynomial $g(x)$ of the narrow-sense BCH code $\mathcal{C}_{(q,n,\delta)}$ can be expressed as
$g(x)=\prod\limits_{i\in\mathcal{G}}(x-\beta^{i}),\quad\mathcal{G}=\bigcup\limits_{a=1}^{\delta-1}C_{n}(a).$
Consequently, the dimension of $\mathcal{C}_{(q,n,\delta)}$ can be given by
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=n-\left|\bigcup\limits_{a=1}^{\delta-1}C_{n}(a)\right|=n-\sum\limits_{a\in\mathcal{L}^{n}(1,\delta-1)}|C_{n}(a)|.$ (2)
Thus, determining the coset leaders modulo $n$ within $1,\delta-1]$ and the sizes of corresponding cyclotomic cosets allows us to compute the dimension of $\mathcal{C}_{(q,n,\delta)}$. Additionally, for $\delta^{{}^{\prime}}>\delta$, one can easily verify that
$\bigcup\limits_{a=1}^{\delta-1}C_{n}(a)=\bigcup\limits_{a=1}^{\delta^{\prime}-1}C_{n}(a)$
if and only if all the integers in $\left[\delta,\delta^{{}^{\prime}}-1\right]$ are not coset leaders. Therefore, the Bose distance of $\mathcal{C}_{(q,n,\delta)}$ is equal to the smallest coset leader within the range $\delta,n-1]$.
We have a useful observation regarding $q$-cyclotomic cosets, as presented in the following lemma. This result originated from the proof of [40, Lemma 6]. For completeness, we include a proof below.
Lemma 1.
Suppose that $n$ and $\lambda$ are two integers such that $\mathrm{gcd}(n,q)=1$ and $\mathrm{gcd}(\lambda,q)=1$. Let $a\in[0,n-1]$ be an integer. Then
  • $a$ is the coset leader of $C_{n}(a)$ if and only if $\lambda a$ is the coset leader of $C_{\lambda n}(\lambda a)$;
  • $|C_{n}(a)|=|C_{\lambda n}(\lambda a)|.$
Proof.
By definition, the integer $a$ is the coset leader of $C_{n}(a)$ if and only if
$a\bmod n\leq aq^{i}\bmod n\quad\hbox{for any integer }i\geq 0.$ (3)
Noticing that $a\bmod n=a-n\cdot\lfloor\frac{a}{n}\rfloor$ and $\lambda a\bmod\lambda n=\lambda a-\lambda n\cdot\lfloor\frac{\lambda a}{\lambda n}\rfloor$, we can assert that
$\lambda\cdot(a\bmod n)=\lambda a-\lambda n\cdot\lfloor\frac{a}{n}\rfloor=\lambda a-\lambda n\cdot\lfloor\frac{\lambda a}{\lambda n}\rfloor=\lambda a\bmod\lambda n.$
Similarly, we also have
$\lambda\cdot(aq^{i}\bmod n)=\lambda aq^{i}\bmod\lambda n.$ (4)
Therefore, the inequality in (3) is equivalent to
$\lambda a\bmod\lambda n\leq\lambda aq^{i}\bmod\lambda n\quad\hbox{for any integer }i\geq 0.$
By definition, this holds if and only if $\lambda a$ is the coset leader of $C_{\lambda n}(\lambda a)$. Therefore, the first statment follows.
In addition, the equality in (4) also implies that
$aq^{i}\bmod n=a\quad\hbox{if and only if}\quad\lambda aq^{i}\bmod\lambda n=\lambda a.$
It follows that the smallest integer $\ell_{a}$ such that $aq^{\ell_{a}}\bmod n=a$ is equal to the smallest integer $\ell_{\lambda a}$ such that $\lambda aq^{\ell_{\lambda a}}\bmod\lambda n=\lambda a$. Therefore, we have $|C_{n}(a)|=|C_{\lambda n}(\lambda a)|.$ This completes the proof. ∎
Let $\mathbb{Z}$ be the set of all integers. For each integer $\lambda$, we denote by $\mathcal{D}_{\lambda}$ the set of integers that are divisible by $\lambda$, that is,
$\mathcal{D}_{\lambda}=\{a\in\mathbb{Z}:\lambda\mid a\}.$
Then one can directly derive the following corollary from the above lemma.
Corollary 1.
Let $n$ and $\lambda$ be two integers such that $\gcd(n,q)=1$ and $\mathrm{gcd}(\lambda,q)=1$. Then
$\left|\mathcal{L}_{\ell}^{n}(b,c)\right|=\left|\mathcal{L}^{\lambda n}_{\ell}(\lambda b,\lambda c)\cap\mathcal{D}_{\lambda}\right|$
for any positive integer $\ell$, and integers $b,c$ with $0\leq b\leq c\leq n-1$.
III Some known results on $q$-cyclotomic cosets modulo $q^{m}-1$
Throughout the rest of the paper, we always assume that $m$ is a positive integer, $h=\lfloor\frac{m}{2}\rfloor$, and $n=\frac{q^{m}-1}{\lambda}$, where $\lambda$ divides q1q-1. This implies that $\lambda n=q^{m}-1$. Additional, for a positive real number $a$, we define $N(a)$ as the number of integers in the interval [1,a1][1,a-1] that are not divisible by $q$, i.e., $N(a)=\lfloor a-1\rfloor-\lfloor(a-1)/{q}\rfloor$.
In this section, we briefly review some known results on the sizes and coset leaders of the $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$. These foundational results will be essential for deriving the main contributions of this paper. To maintain consistency, we follow the notation and terminology used in [39], which we now introduce for completeness. Let $Z_{q}$ denote the set of all non-negative integers less than $q$. Each integer $a\in[0,\lambda n]$ can be uniquely represented by its $q$-adic expansion as $a=\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell},$ where $a_{\ell}\in Z_{q}$ for all $\ell=0,1,\dots,m-1$. Let $Z_{q}^{m}$ be the set of all length-$m$ sequences of integers in $Z_{q}$. For simplicity, denote by $\mathbf{0}_{m}$ the sequence in $Z_{q}^{m}$ whose elements are all zero. We define an order on $Z_{q}^{m}$ using lexicographic order. Specifically, for any two sequences $U=(u_{m-1},\ldots,u_{1},u_{0})$ and $W=(w_{m-1},\ldots,w_{1},w_{0})$ in $Z_{q}^{m}$,
  • 1.
    $U$ and $W$ are said to be equal, denoted by $U=W$, if $u_{\ell}=w_{\ell}$ for $\ell=0,\ldots,m-1,$
  • 2.
    $U$ is less than $W$, denoted by $U<W$, if either $u_{m-1}<w_{m-1}$ or there exists an integer $i\in[0,m-2]$ such that $u_{i}<v_{i}$ and $u_{\ell}=w_{\ell}$ for all $\ell=i+1,\ldots,m-1,$ and
  • 3.
    $U\leq W$ is denoted if $U=W$ or $U<W$.
The map $V$ from the set of all the integers in $0,\lambda n]$ to $Z_{q}^{m}$ is defined as
$V(a)=(a_{m-1},\ldots,a_{1},a_{0}),$
where $\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell}$ forms the $q$-adic expansion of the integer $a\in[0,\lambda n].$
We define the following sets
[…]
Let $m$ and $k$ be two integers such that $m\geq 4$ and $m-2h\leq k\leq\lfloor{(2m-1)}/{3}\rfloor-h$. When $m$ is an odd integer, for each integer $i\in[-k+1,k]$, we define $\mathcal{A}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ that satisfies:
$a_{h+i}>0;$ (5)
$(a_{k+i-1},\ldots,a_{0})\leq(a_{h+k},\ldots,a_{h-i+1})\hbox{ and }a_{0}>0;$ (6)
$V(a)=(\mathbf{0}_{h-k},a_{h+k},\ldots,a_{h+i},\mathbf{0}_{h-k},a_{k+i-1},\ldots,a_{0}).$ (7)
When $m$ is an even integer, for each $i\in[-k,k]$, we define $\mathcal{B}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ that satisfies the condition in (5) and the following:
$(a_{k+i},\ldots,a_{0})\leq(a_{h+k},\ldots,a_{h-i})\hbox{ and }a_{0}>0;$ (8)
$V(a)=(\mathbf{0}_{h-k-1},a_{h+k},\ldots,a_{h+i},\mathbf{0}_{h-k-1},a_{k+i},\ldots,a_{0}).$ (9)
Additionally, we define $t(a)=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}$ and $\alpha(a)=\sum\limits_{\ell=0}^{k+i-1}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{A}_{k}(i)$, and define $t(a)=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}$ and $\alpha(a)=\sum\limits_{\ell=0}^{k+i}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{B}_{k}(i)$. Then we make the following remarks.
Remark 1.
The condition in (5) can be equivalently represented as $q\nmid t(a)$.
Remark 2.
The condition in (7) implies that each integer $a\in\mathcal{A}_{k}(i)$ can be uniquely decomposed as
$a=t(a)q^{h+i}+\alpha(a),$
Since $\lambda\mid q-1$, it follows that for any integer $a\in\mathcal{A}_{k}(i)$,
$\lambda\mid a\quad\hbox{ if and only if }\quad\lambda\mid t(a)+\alpha(a).$
Similarly, each integer $a\in\mathcal{B}_{k}(i)$ can be uniquely decomposed as
$a=t(a)q^{h+i}+\alpha(a),$
and
$\lambda\mid a\quad\hbox{ if and only if }\quad\lambda\mid t(a)+\alpha(a).$
Remark 3.
The condition in (6) can be equivalently expressed as
$1\leq\alpha(a)\leq\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}\quad\hbox{and}\quad q\nmid\alpha(a).$ (10)
Moreover, the form of $V(a)$ in (7) implies that
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}=t(a)\cdot q^{2i-1}\quad\hbox{for }i\in[1,k],$
and
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}+\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}$ $=$ $t(a)\cdot q^{2i-1}\quad\hbox{for }i\in[-k+1,0].$
Noticing that $0\leq\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}<1$ for $i\in[-k+1,0]$, we can conclude that (10) is equivalent to
$1\leq\alpha(a)\leq\lfloor t(a)\cdot q^{2i-1}\rfloor\quad\hbox{and}\quad q\nmid\alpha(a).$
Similarly, the condition in (8) can also be written as
$1\leq\alpha(a)\leq\sum\limits_{\ell=h-i}^{h+k}a_{\ell}q^{\ell-(h-i)}\quad\hbox{and}\quad q\nmid\alpha(a),$
which is further equivalent to
$1\leq\alpha(a)\leq\lfloor t(a)\cdot q^{2i}\rfloor\quad\hbox{and}\quad q\nmid\alpha(a).$
Remark 4.
If $a\in\mathcal{A}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell},$ then $\sum\limits_{\ell=0}^{h-k}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}$. Similarly, if $a\in\mathcal{B}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$, then $\sum\limits_{\ell=0}^{h-k-1}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-i}^{h+k}a_{\ell}q^{\ell-(h-i)}$.
Remark 5.
If $a\in\mathcal{A}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell},$ then $i$ is the smallest integer in [k+1,k][-k+1,k] such that $a_{h+i}>0$. Consequently, $\mathcal{A}_{k}(i)\cap\mathcal{A}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k+1,k]$.
Similarly, if $a\in\mathcal{B}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$, then $i$ is the smallest integer in [k,k][-k,k] such that $a_{h+i}>0$. As a result, $\mathcal{B}_{k}(i)\cap\mathcal{B}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k,k]$.
Theorem 1.
[39, Theorem 2] Let $m$ and $k$ be two integers such that $m\geq 4$ and $m-2h\leq k\leq\lfloor(2m-1)/{3}\rfloor-h.$
  • If $m$ is odd, then
    $\mathcal{S}\cap[q^{h+k},q^{h+k+1})=\bigsqcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i).$
  • If $m$ is even, then
    $(\mathcal{S}\cup\mathcal{H})\cap[q^{h+k},q^{h+k+1})=\bigsqcup\limits_{i=-k}^{k}\mathcal{B}_{k}(i).$
Lemma 2.
[39, Corollary 2] Let $m$ be an even integer and $k\in\left[0,\lfloor(2m-1)/{3}\rfloor-h\right]$ be an integer. Suppose that $a\in\left[q^{h+k},q^{h+k+1}\right)$ is an integer. Then $a\in\mathcal{H}$ if and only if $V(a)$ has the form
$(\mathbf{0}_{h-k-1},a_{k},\ldots,a_{0},\mathbf{0}_{h-k-1},a_{k},\ldots,a_{0})$
with $a_{0}>0$ and $a_{k}>0$.
Theorem 2.
[39, Theorem 1] When $m$ is an odd integer, for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
$|C_{\lambda n}(a)|=m.$
When $m$ is an even integer, for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
$|C_{\lambda n}(a)|=\begin{cases}\frac{m}{2}&\hbox{if }aq^{h}\mathrm{\ mod\ }\lambda n=a,\\ m&\hbox{if }aq^{h}\mathrm{\ mod\ }\lambda n\neq a.\end{cases}$
IV Auxillary Lemmas
The following lemmas are needed to establish the main theorems on the dimension and Bose distance of BCH codes in the subsequent sections. Their proofs are given in Appendices A–H.
Lemma 3.
Suppose that $x$ and $y$ are two positive integers. Then
$\left|\{\alpha\in[1,x]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+y\}\right|=\lfloor\frac{x+y}{\lambda}\rfloor-\lfloor\frac{\lfloor{x}/{q}\rfloor+y}{\lambda}\rfloor.$
Lemma 4.
Suppose that $x$ and $y$ are two integers such that $x\leq y$. Then
[…]
Lemma 5.
Suppose that $x$ and $y$ are two integers. Then
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+x}{\lambda}\rfloor-\lfloor\frac{t+y}{\lambda}\rfloor\right]=\frac{(q-1)(x-y)}{\lambda}.$
Lemma 6.
Suppose that $a$ is a positive integer. Then
$\sum\limits_{t=q,q\nmid t}^{aq}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\frac{a(a-1)(q-1)}{2\lambda}.$
Lemma 7.
Suppose that $x$ is an even integer. Then
[…]
Lemma 8.
Let $k$ and $a\leq q$ be two positive integers. Then
$\sum\limits_{t=q^{k},q\nmid t}^{aq^{k}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor\right]={\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}}+\textstyle{\frac{(3+(-1)^{\lambda})}{4\lambda}(a-1)(q-1)q^{k-1}}.$
Lemma 9.
Let $k$ be an integer, and let $a\leq q$ be a positive integer. Then
[…]
Lemma 10.
Let $k$ and $a\leq q$ be two positive integers. Then
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-3}&\hbox{if }-k+2\leq i\leq k-1,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-2}&\hbox{if }i=k\hbox{ or }-k+1,\end{cases}$ (11)
and
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}&\hbox{if }-k+1\leq i\leq k-1\hbox{ and }i\neq 0,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-1}&\hbox{if }i=k\hbox{ or }-k.\end{cases}$ (12)
V The dimension of $\mathcal{C}_{(q,n,\delta)}$for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$
Let $m\geq 4$ be an integer and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, let $k_{\delta}$ be the integer such that $q^{h+k_{\delta}}\leq\lambda(\delta-1)<q^{h+k_{\delta}+1}$, and let $\sum\limits_{\ell=0}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda(\delta-1)$. If $k_{\delta}\geq m-2h$, let $s_{\delta}$ be the smallest integer in $m-2h-k_{\delta},k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$ and define $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$.
If $m$ is odd, for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$, we define the function $f(\delta)$ by
$f(\delta)=\begin{cases}0,&\hbox{if }\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ {\left(\begin{aligned} &\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]\end{aligned}\right)}&\hbox{if }\delta>\frac{q^{h+1}-1}{\lambda}+1,\end{cases}$ (13)
where $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}$ and $\mathcal{T}_{i}(\delta)=\left\{t\in\mathbb{Z}:q^{k_{\delta}-i}\leq t<\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h-i}\hbox{ and }q\nmid t\right\}.$
If $m$ is even, for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$, we define the function $\widetilde{f}(\delta)$ by
$\widetilde{f}(\delta)=\begin{cases}0&\hbox{if }\delta\leq\frac{q^{h}-1}{\lambda}+1,\\ \lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&\hbox{if }\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ \left({\begin{aligned} &q^{2k_{\delta}-2}(k_{\delta}-\textstyle\frac{1}{2})\textstyle\frac{(q-1)^{2}}{\lambda}+\frac{q-1}{2\lambda}\left(q^{k_{\delta}-1}+\textstyle\frac{1+(-1)^{\lambda}}{2}\right)\\ &+\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\textstyle\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}\left[\lfloor\textstyle\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\textstyle\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]\end{aligned}}\right)&\hbox{if }\delta>\frac{q^{h+1}-1}{\lambda}+1,\end{cases}$ (14)
where $\widetilde{\mu}(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}-1}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta})}\right\}.$ We define the function $\tau(\delta)$ for each integer $\delta\in[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1]$ by
$\tau(\delta)=\begin{cases}1&\hbox{if }\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ &\delta_{h}>0\hbox{ and }\lambda\mid 2\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell},\\ 0&\hbox{otherwise.}\end{cases}$ (15)
Additionally, we define the function $g(\delta)$ for each integer $\delta\in[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1]$ by
$g(\delta)=\begin{cases}0&\hbox{if }\delta\leq\frac{q^{h}-1}{\lambda}+1,\\ \lfloor\frac{(\delta_{h}-1)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor+\tau(\delta)&\hbox{if }\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ \lfloor\frac{\phi(\delta)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor-\lfloor\frac{\lfloor\phi(\delta)/q\rfloor\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor+\tau(\delta)&\hbox{if }\frac{q^{h+1}-1}{\lambda}+1<\delta,\end{cases}$ (16)
where $\phi(\delta)=\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$.
Theorem 3.
Let $m$ and $\delta$ be integers with $m\geq 4$ and $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$.
  • If $m$ is odd, then
    $\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=n-m\left[N(\delta)-f(\delta)\right].$ (17)
  • If $m$ is even, then
    $\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=n-m\left[N(\delta)-\widetilde{f}(\delta)\right]-\frac{m}{2}g(\delta).$ (18)
Assertion 1.
Let $m$ and $\delta$ be integers with $m\geq 4$ and $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
  • If $m$ is odd, then
    $\left|\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor.$ (19)
  • If $m$ is even, then
    $\left|\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor$ (20)
    and
    $\left|\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\tau(\delta).$ (21)
Assertion 2.
Let $m$ and $\delta$ be integers with $m\geq 4$ and $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
  • If $m$ is odd, then
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].$ (22)
  • If $m$ is even, then
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]$ (23)
    and
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\begin{cases}\lfloor\frac{(\delta_{h}-1)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor&\hbox{if }k_{\delta}=0,\\ \lfloor\frac{\phi(\delta)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor-\lfloor\frac{\lfloor\phi(\delta)/q\rfloor\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor-\frac{q^{k_{\delta}-1}(q-1)\left(3+(-1)^{\lambda}\right)}{2\lambda}&\hbox{if }k_{\delta}\geq 1,\end{cases}$ (24)
where $\phi(\delta)=\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$.
Assertion 3.
Let $m$ and $k$ be two integers with $m\geq 4$ and $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$.
  • If $m$ is odd, then
    $|\mathcal{A}_{k}(i)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{1}{2\lambda}(q-1)^{3}(q+1)q^{2k-3}&\hbox{if }-k+2\leq i\leq k-1,\\ \frac{1}{2\lambda}(q-1)^{2}q^{2k-1}&\hbox{if }i=-k+1\hbox{ or }k.\end{cases}$ (25)
  • If $m$ is even, then
    $|\mathcal{B}_{k}(i)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{1}{2\lambda}(q-1)^{3}(q+1)q^{2k-2}&\hbox{if }-k+1\leq i\leq k-1\hbox{ and }i\neq 0,\\ \frac{1}{2\lambda}(q-1)^{2}q^{2k}&\hbox{if }i=-k\hbox{ or }k.\end{cases}$ (26)
Assertion 4.
Let $m\geq 4$ be an even integer, and let $k$ be an integer with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
  • If $\lambda$ is odd, then
    $|B_{k}(0)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{q(q-1)}{2\lambda}&\hbox{if $k=0$,}\\ \frac{(q-1)^{2}}{2\lambda}(q^{2k}-q^{2k-2}+q^{k-1})&\hbox{if $k\geq 1$.}\end{cases}$
  • If $\lambda$ is even, then
    $|B_{k}(0)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{(q+1)(q-1)}{2\lambda}&\hbox{if $k=0$,}\\ \frac{(q-1)^{2}}{2\lambda}(q^{2k}-q^{2k-2}+2q^{k-1})&\hbox{if $k\geq 1$.}\end{cases}$
Assertion 5.
Let $m\geq 4$ be an even integer, and let $k$ be an integer with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
  • If $\lambda$ is odd, then
    $\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\begin{cases}\frac{q-1}{\lambda}&\hbox{if }k=0,\\ \frac{(q-1)^{2}q^{k-1}}{\lambda}&\hbox{if }k\geq 1.\end{cases}$ (27)
  • If $\lambda$ is even, then
    $\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\begin{cases}\frac{2(q-1)}{\lambda}&\hbox{if }k=0,\\ \frac{2(q-1)^{2}q^{k-1}}{\lambda}&\hbox{if }k\geq 1.\end{cases}$ (28)
Proof of Theorem 3. Suppose that $m$ is odd. We first aim to show that
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=f(\delta)\quad\hbox{for }2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ (29)
If $2\leq\delta\leq\frac{q^{h+1}-1}{\lambda}+1$, then $\lambda\leq\lambda(\delta-1)\leq q^{h+1}-1$. By applying [35, Theorem 2.3], we can obtain
$1,\lambda(\delta-1)]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}=0.$
In particular,
$\left|\left[1,q^{h+1}-1\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=0.$ (30)
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. This implies $1\leq k_{\delta}\leq\lfloor(2m-1)/{3}\rfloor-h$. By applying Theorem 1, we can conclude from Assertion 3 that for any integer $k\in\left[1,\lfloor({2m-1})/{3}\rfloor-h\right],$
$\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$ $=$ $\sum\limits_{i=-k+1}^{k}\left|\mathcal{A}_{k}(i)\cap\mathcal{D}_{\lambda}\right|$
$=$ $\frac{1}{\lambda}\left[(k-1)q^{2k-3}(q-1)^{3}(q+1)+q^{2k-1}(q-1)^{2}\right].$
Therefore,
$\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$ $=$ $\sum\limits_{{k}=1}^{k_{\delta}-1}\left|\left[q^{h+{k}},q^{h+{k}+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$
$=$ $\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{(q-1)^{3}}{\lambda}(q+1)({k}-1)q^{2{k}-3}+\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{q^{2{k}-1}(q-1)^{2}}{\lambda}.$
It is straightforward to verify that
$\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{q^{2{k}-1}(q-1)^{2}}{\lambda}=\frac{(q^{2{k_{\delta}}-1}-q)(q-1)}{\lambda(q+1)}$
and
$\sum\limits_{k=1}^{k_{\delta}-1}\frac{(q-1)^{3}}{\lambda}(q+1)({k}-1)q^{2{k}-3}=\frac{\left[q-(k_{\delta}-1)q^{2k_{\delta}-3}+(k_{\delta}-2)q^{2k_{\delta}-1}\right](q-1)}{\lambda(q+1)}.$
By adding these two sums, we can obtain
$\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}.$ (31)
Combining equations (19), (22), (30) and (31) we obtain
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$ $=\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor$
$+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].$
Recalling the definition of $f(\delta)$, we can now claim that equation (29) holds.
Next, we establish equation (17). Noticing that $\lfloor({2m-1})/{3}\rfloor+1\leq m-\lfloor{m}/{3}\rfloor$, we have $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. Therefore, we can apply Theorem 2 and Lemma 1 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\quad\hbox{for any integer }a\in[1,\delta-1]$ (32)
and
$\mathcal{L}_{m}^{\lambda n}(1,\lambda(\delta-1))=\mathcal{L}^{\lambda n}(1,\lambda(\delta-1)).$
Recalling the definition of $\mathcal{S}$ and noting that an integer $a$ cannot be the coset leader of $C_{n}(a)$ if $q\mid a$, it follows that
$\left|\mathcal{L}_{m}^{\lambda n}(1,\lambda(\delta-1))\cap\mathcal{D}_{\lambda}\right|$ $=\left|\mathcal{L}^{\lambda n}(1,\lambda(\delta-1))\cap\mathcal{D}_{\lambda}\right|$
$=\left|\left\{a\in[1,\lambda(\delta-1)]:q\nmid a\right\}\cap\mathcal{D}_{\lambda}\right|-\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|.$
It can be easily verified that
$\left|\left\{a\in[1,\lambda(\delta-1)]:q\nmid a\right\}\cap\mathcal{D}_{\lambda}\right|=N(\delta).$
Utilizing Corollary 1 and equation (29), we have
$\left|\mathcal{L}_{m}^{n}(1,\delta-1)\right|=\left|\mathcal{L}_{m}^{\lambda n}(1,\lambda(\delta-1))\cap\mathcal{D}_{\lambda}\right|=N(\delta)-f(\delta).$ (33)
Recalling the equality in (2), we can now conclude from (32) and (33) that (17) holds.
Suppose that $m$ is even. Our first goal is to show that
$\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\widetilde{f}(\delta)$ (34)
and
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=g(\delta)$ (35)
for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then $2\lambda\leq\lambda(\delta-1)\leq q^{h}-1$. By applying [35, Theorem 2.3], we can conclude that
$\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=0$
and
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=0.$
In particular, we have
$\left|\left[1,q^{h}-1\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=0$ (36)
and
$\left|\left[1,q^{h}-1\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=0.$ (37)
If $\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ then we have $q^{h}\leq\lambda(\delta-1)\leq q^{h+1}-1$. Since $m$ is even, we have
  • $k_{\delta}=0$;
  • $\mathcal{T}_{0}(\delta)=\{t\in\mathbb{Z}:1\leq t\leq\delta_{h}-1\}.$
By substituting $k_{\delta}$ and $\mathcal{T}_{0}(\delta)$ as above into equations (20) and (23), we obtain
$\left|\left[\delta_{h}q^{h},\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor$
and
$\left|\left[q^{h},\delta_{h}q^{h}\right)\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right].$
Combining the above two equalities with (36), we obtain
$\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right].$ (38)
Additionally, by substituting $k_{\delta}=0$ into (21) and (24), we have
$\left|\left[\delta_{h}q^{h},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\tau(\delta)$
and
$\left|\left[q^{h},\delta_{h}q^{h}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor.$
Combining these two equalities with (37), we obtain
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor+\tau(\delta).$ (39)
Notice that $\lambda(\delta-1)=q^{h+1}-1=\sum\limits_{\ell=0}^{h}(q-1)q^{\ell}$ when $\delta=\frac{q^{h+1}-1}{\lambda}+1$. Therefore, for $\delta=\frac{q^{h+1}-1}{\lambda}+1$, we have
  • $\widetilde{\mu}(\delta)=q-1$;
  • $w_{\delta}=(q-1)q^{h}$;
  • $\delta_{h}=q-1.$
Consequently, by substituting $\delta=\frac{q^{h+1}-1}{\lambda}+1$ into equation (38) and applying Lemma 7, we obtain
$\begin{split}\left|\left[1,q^{h+1}-1\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|&=\lfloor\frac{(q-1)(q^{h}+1)}{\lambda}\rfloor-\lfloor\frac{(q-1)q^{h}}{\lambda}\rfloor+\sum\limits_{t=1}^{q-2}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]\\ &=\frac{q-1}{2\lambda}\left(q+\frac{1+(-1)^{\lambda}}{2}\right).\end{split}$ (40)
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then we have $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. This implies that $1\leq k_{\delta}\leq\lfloor({2m-1})/{3}\rfloor-h.$ By applying Theorem 1, we can conclude from Assertions 3 and 4 that
$\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{i=-k}^{k}|\mathcal{B}_{k}(i)|=\frac{(q-1)^{3}}{\lambda}(k-\frac{1}{2})q^{2k-2}(q+1)+)\frac{(q-1)^{2}}{\lambda}(\frac{1}{2}q^{k-1}+q^{2k})$
for each integer $k\in[1,\lfloor({2m-1})/{3}\rfloor-h]$. It follows that
$\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$ $=$ $\sum\limits_{k=1}^{k_{\delta}-1}\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$
$=$ $\frac{(q-1)^{2}}{\lambda}(k_{\delta}-\frac{1}{2})q^{2k_{\delta}-2}+\frac{q-1}{2\lambda}(q^{k_{\delta}-1}-q).$
Combining (20), (23), (40) and (V), we derive
$\begin{split}\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=&\frac{q-1}{2\lambda}\left(q^{k_{\delta}-1}+\frac{1+(-1)^{\lambda}}{2}\right)+\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\frac{(q-1)^{2}}{\lambda}(k_{\delta}-\frac{1}{2})q^{2k_{\delta}-2}+\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]\end{split}$
On the other hand, we conclude from Assertion 5 that
$\left|\left[q^{h},q^{h+k_{\delta}}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{k=0}^{k_{\delta}-1}\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\frac{(3+(-1)^{\lambda})(q-1)q^{k_{\delta}-1}}{2\lambda}.$
With (21) and (24), it follows that
$\left|\left[1,\lambda(\delta-1)\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\phi(\delta)\left(3+(-1)^{\lambda}\right)}{\lambda}\rfloor-\lfloor\frac{\lfloor\phi(\delta)/q\rfloor\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor+\tau(\delta),$
where $\phi(\delta)=\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$.
By now, we have already demonstrated that both (34) and (35) hold for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ Next, we show that equation (18) holds. Notice that $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. We can apply Theorem 2 and Lemma 1 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\hbox{ or }\frac{m}{2}\quad\hbox{for any integer }a\in[1,\delta-1].$ (42)
By applying Corollary 1, we can derive from (34) and (35) that
$\left|\mathcal{L}^{n}_{m}(1,\delta-1)\right|=\left|\mathcal{L}^{\lambda n}_{m}\left(1,\lambda(\delta-1)\right)\cap\mathcal{D}_{\lambda}\right|=N(\delta)-\widetilde{f}(\delta)$ (43)
and
$\left|\mathcal{L}^{n}_{\frac{m}{2}}(1,\delta-1)\right|=\left|\mathcal{L}^{\lambda n}_{\frac{m}{2}}\left(1,\lambda(\delta-1)\right)\cap\mathcal{D}_{\lambda}\right|=g(\delta).$ (44)
Recalling the equality in (2), we can now conclude from (42), (43) and (44) that (18) holds. This completes the proof. ∎
As examples of Theorem 3, we present the dimension of $\mathcal{C}_{(q,n,\delta)}$ for $2\leq\delta\leq\frac{q^{\lceil\frac{m}{2}\rceil+1}-1}{\lambda}+1$ by providing the following two Corollaries. Notably, the formulas given in Corollary 2 provide the dimension of the same narrow-sense BCH codes as those studied in [40, Theorem 3], but they are presented in a simpler form.
Corollary 2.
Let $m\geq 4$ be an even integer, and let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{h+1}-1}{\lambda}+1$. Then
[…]
Proof.
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then we have $\widetilde{f}(\delta)=0$ and $g(\delta)=0$. If $\delta>\frac{q^{h}-1}{\lambda}+1$, then we distinguish the following two cases.
Case 1. Suppose that $\delta_{h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Then we have $\widetilde{\mu}(\delta)=\delta_{h}$ and $w_{\delta}=\delta_{h}q^{h}$. By substituting them into equation (14), we obtain
$\begin{split}\widetilde{f}(\delta)&=\lfloor\frac{\delta_{h}+\delta_{h}q^{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h}q^{h}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]\\ &=\sum\limits_{t=1}^{\delta_{h}}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right].\end{split}$
In addition, it is straightforward to verify that
$\tau(\delta)=\begin{cases}1&\hbox{if }\lambda\mid 2\delta_{h},\\ 0&\hbox{if }\lambda\nmid 2\delta_{h}.\end{cases}$
Recalling equation (16), it follows that
$g(\delta)=\begin{cases}\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor+1&\hbox{if }\lambda\mid 2\delta_{h},\\ \lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor&\hbox{if }\lambda\nmid 2\delta_{h}.\end{cases}$
Case 2. Suppose that $\delta_{h}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Then $\widetilde{\mu}(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{0}$, $w_{\delta}=\delta_{h}q^{h}$ and $\tau(\delta)=0$. By substituting these values into equations (14) and (16), we obtain
$\widetilde{f}(\delta)=\lfloor\frac{\delta_{0}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{2\delta_{h}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]$
and
$g(\delta)=\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor.$
Finally, substituting the values of $g(\delta)$ and $\widetilde{f}(\delta)$ for corresponding cases into equation (18), we derive the desired equation (45). ∎
Corollary 3.
Let $m\geq 5$ be an odd integer, and let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{h+2}-1}{\lambda}+1.$ Then
[…]
Proof.
If $\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ we can directly have $f(\delta)=0$. If $\delta>\frac{q^{h+1}-1}{\lambda}+1,$ then we have $k_{\delta}=1$. We distinguish following cases.
Case 1. Suppose that $\delta_{h}>0$. Then we have $s_{\delta}=0$. This leads to
  • $\mathcal{T}_{0}(\delta)=\left\{t\in\mathbb{Z}:q\leq t\leq\delta_{h+1}q+\delta_{h}-1\hbox{ and }q\nmid t\right\}$;
  • $\mathcal{T}_{1}(\delta)=\left\{t\in\mathbb{Z}:1\leq t\leq\delta_{h+1}\right\}$;
  • $w_{\delta}=\delta_{h+1}q^{h+1}+\delta_{h}q^{h}$;
  • $\mu(\delta)=\min\left\{\delta_{h+1},\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}\right\}$.
Recall equation (13). It follows that
$\begin{split}f(\delta)=&\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{{\lfloor\mu(\delta})/q\rfloor+w_{\delta}}{\lambda}\rfloor+\sum\limits_{t=q,q\nmid t}^{\delta_{h+1}q+\delta_{h}-1}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]+\sum\limits_{t=1}^{\delta_{h+1}}\left[\lfloor\frac{tq+t}{\lambda}\rfloor-\lfloor\frac{2t}{\lambda}\rfloor\right].\end{split}$ (47)
Note that $\mu(\delta)=\delta_{h+1}$ if $\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, and $\mu(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{0}$ if $\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Thus,
$\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{{\lfloor\mu(\delta})/q\rfloor+w_{\delta}}{\lambda}\rfloor=\begin{cases}\lfloor\frac{\delta_{0}+\delta_{h+1}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+\delta_{h}}{\lambda}\rfloor&\hbox{if }\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ \lfloor\frac{2\delta_{h+1}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+\delta_{h}}{\lambda}\rfloor&\hbox{if }\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$ (48)
By applying Lemma 6, we derive
$\sum\limits_{t=q,q\nmid t}^{\delta_{h+1}q}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\frac{\delta_{h+1}(\delta_{h+1}-1)(q-1)}{2\lambda}.$
Since each integer $t\in\left[\delta_{h+1}q+1,\delta_{h+1}q+\delta_{h}-1\right]$ can be uniquely expressed as $t=\delta_{h+1}q+i$ with $i\in\left[1,\delta_{h}-1\right]$, we have
$\sum\limits_{t=\delta_{h+1}q+1}^{\delta_{h+1}q+\delta_{h}-1}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\sum\limits_{i=1}^{\delta_{h}-1}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right].$
Adding above two sums, we get
$\sum\limits_{t=q,q\nmid t}^{\delta_{h+1}q+\delta_{h}-1}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\frac{\delta_{h+1}(\delta_{h+1}-1)(q-1)}{2\lambda}+\sum\limits_{i=1}^{\delta_{h}-1}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right].$ (49)
Furthermore, it is straightforward to verify that
$\begin{split}\sum\limits_{t=1}^{\delta_{h+1}}\left[\lfloor\frac{tq+t}{\lambda}\rfloor-\lfloor\frac{2t}{\lambda}\rfloor\right]&=\sum\limits_{t=1}^{\delta_{h+1}}\frac{t(q-1)}{\lambda}\\ &=\frac{\delta_{h+1}(\delta_{h+1}+1)(q-1)}{2\lambda}.\end{split}$ (50)
We can now conclude from (47) – (50) that
$f(\delta)=\begin{cases}\lfloor\frac{\delta_{0}+\delta_{h+1}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+\delta_{h}}{\lambda}\rfloor+\frac{\delta_{h+1}^{2}(q-1)}{\lambda}+\sum\limits_{i=1}^{\delta_{h}-1}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right]&\hbox{if }\delta_{h}>0\hbox{ and }\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ \frac{\delta_{h+1}^{2}(q-1)}{\lambda}+\sum\limits_{i=1}^{\delta_{h}}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right]&\hbox{if }\delta_{h}>0\hbox{ and }\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
Case 2. Suppose that $\delta_{h}=0$. Then we have $s_{\delta}=1$. This leads to
  • $\mathcal{T}_{0}(\delta)=\left\{t\in\mathbb{Z}:q\leq t\leq\delta_{h+1}q-1\hbox{ and }q\nmid t\right\}$;
  • $\mathcal{T}_{1}(\delta)=\{t\in\mathbb{Z}:1\leq t\leq\delta_{h+1}-1\}$;
  • $w_{\delta}=\delta_{h+1}q^{h+1}$;
  • $\mu(\delta)=\min\left\{\delta_{h+1}q,\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}\right\}.$
Note that $\mu(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{1}q+\delta_{0}$ if $\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, and $\mu(\delta)=\delta_{h+1}q$ if $\delta_{h+1}q\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Thus, by substituting these terms into equation (13), we can use an analogous argument as in Case 1 to obtain
$f(\delta)=\begin{cases}\frac{(\delta_{h+1}^{2}-\delta_{h+1}+\delta_{1})(q-1)}{\lambda}+\lfloor\frac{\delta_{1}+\delta_{0}+\delta_{h+1}}{\lambda}\rfloor-\lfloor\frac{\delta_{1}+\delta_{h+1}}{\lambda}\rfloor&\hbox{if }\delta_{h}=0\hbox{ and }\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ \frac{\delta_{h+1}^{2}(q-1)}{\lambda}&\hbox{if }\delta_{h}=0\hbox{ and }\delta_{h+1}q\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
Finally, substituting the value of $f(\delta)$ into equation (17) for each corresponding case, we can conclude that (46) holds. This completes the proof. ∎
VI The Bose distance of $\mathcal{C}_{(q,n,\delta)}$
In this section, we investigate the Bose distance of BCH codes $\mathcal{C}_{(q,n,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$. Note that if $q\mid\delta$, then $\delta$ is not a coset leader modulo $n$. Therefore, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=d_{B}(\mathcal{C}_{(q,n,\delta+1)})$ for $q\mid\delta$. Given this property, it is sufficient to focus on BCH codes $\mathcal{C}_{(q,n,\delta)}$ with $q\nmid\delta$. Their Bose distances are established in the following three theorems.
Theorem 4.
Let $m$ and $\delta$ be two positive integers with $2\leq\delta\leq\frac{q^{m-h}-1}{\lambda}$ and $q\nmid\delta$. Then $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
Proof.
Note that $\lambda\mid q-1$ implies that $\mathrm{gcd}(q,\lambda)=1$. Since $q\nmid\delta$, it follows that $q\nmid\lambda\delta$. Additionally, it is clear that $2\leq\lambda\delta<q^{m-h}$. By [35, Theorem 2.3], it follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Then applying Lemma 1, we conclude that $\delta$ is a coset leader modulo $n$. Consequently, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
Theorem 5.
Let $m\geq 5$ be an odd integer, and $\delta$ be an integer with $\frac{q^{h+1}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Let $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$ and $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda\delta$. Let $r_{\delta}$ be the smallest integer in $j_{\delta}+1,\,j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$ Then
$d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(m-h-r_{\delta})},\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\hbox{ and }\delta_{h-r_{\delta}+1}+\lambda-q\neq\hat{\delta}\bmod\lambda,\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\hbox{ and }\delta_{h-r_{\delta}+1}+\lambda-q=\hat{\delta}\bmod\lambda.\end{cases}$ (51)
Proof.
We show that equation (51) holds through the following cases.
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. By Remarks 4 and 5, the inequality implies that $\lambda\delta\not\in\mathcal{A}_{j_{\delta}}$ for every integer $i\in[-j_{\delta}+1,j_{\delta}]$. Therefore, by Theorem 1, we conclude $\lambda\delta\not\in\mathcal{S}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. It follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Applying Lemma 1, we further obtain that $\delta$ is a coset leader modulo $n$. Consequently, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
Case 2. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. With an argument similar to that in the proof of [39, Theorem 6], we can show that each integer in $\lambda\delta,\hat{\delta}$] is not a coset leader modulo $\lambda n$. Applying Lemma 1, it follows that each integer in $\delta,\lfloor\frac{\hat{\delta}}{\lambda}\rfloor]$ is not a coset leader modulo $n$. Notice that when $m$ is odd,
$\hat{\delta}=\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\hat{\delta}\bmod\lambda=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$
Thus, we have
$\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+2}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}+\delta_{h-r_{\delta}+1}-\hat{\delta}\bmod\lambda+\lambda.$
This implies that
$q\mid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda\quad\hbox{if and only if}\quad\delta_{h-r_{\delta}+1}+\lambda-q=\hat{\delta}\bmod\lambda.$
Then we distinguish the following two subcases:
Subcase 1. If $\delta_{h-r_{\delta}+1}+\lambda-q\neq\hat{\delta}\bmod\lambda$, then $q\nmid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$. Applying a similar argument as in Case 1, we can show that $\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$ is a coset leader modulo $\lambda n=q^{m}-1$. By applying lemma 1, it follows that $\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1$ is a coset leader modulo $n$. Consequently, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1.$
Subcase 2. If $\delta_{h-r_{\delta}+1}+\lambda-q=\hat{\delta}\bmod\lambda$, then $q\mid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$ and $q\nmid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2\lambda$. This implies that $\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$ is not a coset leader of $\lambda n$. Furthermore, using a similar argument as in Case 1 again, we can conclude that while $\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2\lambda$ is a coset leader modulo $\lambda n$. By applying Lemma 1, it follows that $\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1$ is not a coset leader modulo $n$, while $\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2$ is a coset leader modulo $n$. Therefore, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2.$
Theorem 6.
Let $m\geq 4$ be an even integer, and let $\delta$ be an integer with $\frac{q^{h}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Let $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$ and $\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda\delta$. Let $r_{\delta}$ be the smallest integer in $j_{\delta},j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and define $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}.$
  • If $r_{\delta}\neq 0$, then
    $d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})},\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\text{ and }\delta_{h-r_{\delta}}+\lambda-q\neq\hat{\delta}\bmod\lambda,\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\text{ and }\delta_{h-r_{\delta}}+\lambda-q=\hat{\delta}\bmod\lambda.\end{cases}$ (52)
  • If $r_{\delta}=0$, then
[…]
Proof.
We can use an argument analous to that in the proof of Theorem 5 to conclude that 52) hlods if $r_{\delta}\neq 0$. Thus, we now only demonstrate equation (53) by considering the following cases.
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\geq\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}$. If the inequality is strict, by Theorem 1 together and Remarks 45, we have $\lambda\delta\notin\mathcal{S}\cup\mathcal{H}$; If equality holds, by Lemma 2, we $\lambda\delta\in\mathcal{H}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. Therefore, in either case, $\lambda\delta$ is a coset leader modulo $\lambda n$. By Lemma 1, it follows that $\delta$ is a coset leader of modulo $n$. Therefore, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
[…]
Subcase 1. Assume that $\lambda\mid\hat{\delta}$. Using Lemma 2, it is straightforward to verify that $\hat{\delta}\in\mathcal{H}$, and hence $\hat{\delta}$ is a coset leader modulo $\lambda n.$ It follows that $\frac{\hat{\delta}}{\lambda}$ is a coset leader modulo $n$. Therefore, we can obtain $d_{B}(\mathcal{C}_{(q,n,\delta)})=\frac{\hat{\delta}}{{\lambda}}$.
[…]
An $[n,k,d]$ linear code over $\mathbb{F}_{q}$ is called optimal if there does not exist an $n,k^{\prime},d]$ linear code with $k^{\prime}>k$, or an $n,k,d^{\prime}]$ linear code with $d^{\prime}>d$ over $\mathbb{F}_{q}$. Using the formulas provided in Theorems 3 and 6, we identify several narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ that are either optimal or have parameters matching those of the best known linear codes in the database. Most of these codes occur in the primitive case $\lambda=1$. See Tables II – III in our previous work [39] for examples of these primitive BCH codes. In addition, we list further examples of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ with $\lambda\neq 1$ in Table I. The parameters of these codes are nearly optimal or best known. All parameters were verified by Magma, and for all BCH codes in the table, the Bose distance $d_{B}$ coincides with the minimum distance $d$.
Table I: Examples of BCH code $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ with $\lambda\neq 1$
$q$ $m$ $\lambda$ $n$ $\delta$ $k$ $d_{B}$ Optimality
3 4 2 40 2 36 2 $d_{\text{optimal}}=3$
3 4 2 40 3 – 4 32 4 $d_{\text{optimal}}=5$
3 4 2 40 5 28 5 $d_{\text{best}}=6$
3 4 2 40 6 – 7 26 7 Best known
3 4 2 40 8 22 8 $d_{\text{best}}=9$
3 5 2 121 6 – 7 101 77 $d_{\text{best}}=8$
3 5 2 121 9 – 10 91 1010 $d_{\text{best}}=11$
4 4 3 85 4 – 5 73 55 $d_{\text{best}}=6$
VII BCH codes $\mathcal{C}_{(q,n,\delta)}$with $\lambda\delta=aq^{h+k}+b$.
As an illustration of our main results, this section presents the dimension and Bose distance of the BCH code $\mathcal{C}_{(q,n,\delta)}$ with $\lambda\delta=aq^{h+k}+b$, where $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$ and $q\nmid b$.
Corollary 4.
Let $m\geq 5$ be an odd integer, and let $k$ be an integer with $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$. If $\lambda\delta=aq^{h+k}+b$ for some integers $a\in[1,q-1]$ and $b\in[\lambda,q^{h-k+1}]$ such that $q\nmid b$, then
$d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }b>aq^{2k-1},\\ \lfloor\frac{aq^{h+k}+aq^{2k-1}}{\lambda}\rfloor+1&\hbox{if }b\leq aq^{2k-1}\end{cases}$ (54)
and
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}n-mN(\delta)+\frac{q-1}{\lambda}ma^{2}q^{2k-3}\left[(q-1)k+1\right]&\hbox{if }b\geq aq^{2k-1}+\lambda,\\ n-\frac{q-1}{\lambda}maq^{h+k-1}+\frac{q-1}{\lambda}maq^{2k-3}\left[a(q-1)k+a-q\right]&\hbox{if }b<aq^{2k-1}+\lambda.\end{cases}$ (55)
Proof.
Note that $q\nmid b$ implies that $q\nmid\lambda\delta$. Since $\lambda$ is a factor q1q-1, it follows that $q\nmid\delta$. Then by applying Theorem 5, we can directly derive equation (54).
We now demonstrate that (55) holds. It is clear that $V(\lambda(\delta-1))$ has the form
$(\mathbf{0}_{h-k},a,\mathbf{0}_{2k-1},\delta_{h-k},\ldots,\delta_{0})$
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. By definition, it follows that
  • $s_{\delta}=k_{\delta}=k$;
  • $w_{\delta}=aq^{h+k}$;
  • $\mu(\delta)=\min\left\{b-\lambda,aq^{2k-1}\right\}$;
  • $\mathcal{T}_{i}(\delta)=\left[q^{k-i},aq^{k-i}\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ for each integer $i\in[-k+1,k]$.
Then applying Lemma 10, we can obtain
$\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-3}&\hbox{if }-k+2\leq i\leq k-1,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-2}&\hbox{if }i=k\hbox{ or }-k+1.\end{cases}$
It follows that
$\begin{split}\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)&=\sum\limits_{i=-k+1}^{k}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)\\ &=\frac{1}{\lambda}\left[a(a-1)(q-1)q^{2k-2}+(a^{2}-1)(k-1)(q-1)^{2}q^{2k-3}\right].\end{split}$
By substituting the values of $k_{\delta}$ and $\mu(\delta)$, and the above expression into (13), we obtain
$f(\delta)=\begin{cases}\frac{q-1}{\lambda}a^{2}q^{2k-3}\left[k(q-1)+1\right]&\hbox{if }b\geq aq^{2k-1}+\lambda,\\ \frac{q-1}{\lambda}aq^{2k-3}\left[ak(q-1)+a-q\right]+\frac{b+a-\lambda}{\lambda}-\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor&\hbox{if }b<aq^{2k-1}+\lambda.\end{cases}$
Furthermore, notice that
$\begin{split}N(\delta)-\left[\frac{b+a-\lambda}{\lambda}-\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor\right]&=\frac{aq^{h+k}+b-\lambda}{\lambda}-\lfloor\frac{aq^{h+k}+b-\lambda}{\lambda q}\rfloor-\frac{b+a-\lambda}{\lambda}+\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor\\ &=\frac{aq^{h+k-1}(q-1)}{\lambda}-\lfloor\frac{aq+b-\lambda}{\lambda q}\rfloor+\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor\\ &=\frac{aq^{h+k-1}(q-1)}{\lambda}.\end{split}$
Therefore, we can apply Theorem 3 to derive the desired equality in (55). ∎
Corollary 5.
Let $m\geq 4$ be an even integer, and let $k$ be an integer with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h$. Suppose that $\lambda\delta=aq^{h+k}+b$ for some integers $a\in[1,q-1]$ and $b\in[\lambda,q^{h-k}]$ such that $q\nmid b$.
  • If $k=0$, then
[…]
    and
    $\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})\!=\!\begin{cases}n\!-\!mN(\delta)\!+\!m\left[\frac{a+b-\lambda}{\lambda}\!-\!\lfloor\frac{2a}{\lambda}\rfloor+\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor\!-\![\lfloor\frac{t}{\lambda}\rfloor\right]\right]\!-\!\frac{m(a-1)\left(3+(-1)^{\lambda}\right)}{4\lambda}&\hskip-6.0pt\hbox{if }b<a+\lambda,\\ n\!-\!mN(\delta)+m\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor\!-\!\lfloor\frac{t}{\lambda}\rfloor\right]-\frac{m(a-1)\left(3+(-1)^{\lambda}\right)}{4\lambda}&\hskip-6.0pt\hbox{if }b\geq a+\lambda\hbox{ and }\lambda\nmid 2a,\\ n\!-\!mN(\delta)\!+\!m\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor\!-\!\lfloor\frac{t}{\lambda}\rfloor\right]-\frac{m(a-1)\left(3+(-1)^{\lambda}\right)}{4\lambda}\!-\!\frac{m}{2}&\hskip-6.0pt\hbox{if }b\geq a+\lambda\hbox{ and }\lambda\mid 2a.\end{cases}$ (57)
  • If $k\geq 1$, then
[…]
    and
[…]
Proof.
By Theorem 6, it is easy to conclude that (56) holds if $k=0$, and (58) holds if $k\geq 1$. Therefore, we only demonstrate that (57) holds if $k=0$, and (59) holds if $k\geq 1$.
If $k=0$, then we have $\lambda(\delta-1)=aq^{h}+b-\lambda$ with $0\leq b-\lambda<q^{h-1}$. It follows that
  • $\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1$;
  • $\delta_{h}=a$;
  • $\widetilde{\mu}(\delta)=\min\{b-\lambda,a\}$;
  • $w_{\delta}=aq^{h}$;
  • $\tau(\delta)=1$ if $a\leq b-\lambda$ and $\lambda\mid 2a$, and $\tau(\delta)=0$ if otherwise.
By substituting the above values for the corresponding case into equations (14) and (16), we derive
$\widetilde{f}(\delta)=\begin{cases}\frac{a+b-\lambda}{\lambda}-\lfloor\frac{2a}{\lambda}\rfloor+\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&\hbox{if }b<a+\lambda,\\ \sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&\hbox{if }b\geq a+\lambda,\end{cases}$
and
$g(\delta)=\begin{cases}\frac{(a-1)(3+(-1)^{\lambda})}{2\lambda}+1&\hbox{if }b\geq a+\lambda\hbox{ and }\lambda\mid 2a,\\ \frac{(a-1)(3+(-1)^{\lambda})}{2\lambda}&\hbox{otherwise}.\end{cases}$
Then by Theorem 3, it follows that (57) holds.
If $k\geq 1$, then we have $k_{\delta}=k$ and $V(\lambda(\delta-1))$ has the form
$(\mathbf{0}_{h-k-1},a,\mathbf{0}_{2k},\delta_{h-k-1},\ldots,\delta_{0})$
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. By definition, it follows that
  • $w_{\delta}=aq^{h}$;
  • $s_{\delta}=k_{\delta}=k$;
  • $\tau(\delta)=0$;
  • $\widetilde{\mu}(\delta)=\min\left\{b-\lambda,aq^{2k}\right\}$;
  • $\phi(\delta)=aq^{k}-1$;
  • $\mathcal{T}_{i}(\delta)=\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ for each integer $i\in[-k,k]$.
Then applying Lemmas 8 and 10, we obtain
$\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}&\hbox{if }-k+1\leq i\leq k-1\hbox{ and }i\neq 0,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-1}&\hbox{if }i=k\hbox{ or }-k,\\ {\left(\begin{aligned} &\textstyle{\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}}\\ &+\textstyle{\frac{(3+(-1)^{\lambda})}{4\lambda}(a-1)(q-1)q^{k-1}}\end{aligned}\right)}&\hbox{if }i=0.\end{cases}$
It follows that
$\begin{split}\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=&\frac{1}{\lambda}(k-\frac{1}{2})(a^{2}-1)(q-1)^{2}q^{2k-2}+\frac{1}{\lambda}a(a-1)(q-1)q^{2k-1}\\ &+\frac{(3+(-1)^{\lambda})(a-1)q^{k-1}(q-1)}{4\lambda}.\end{split}$
By substituting $k_{\delta}$, $\tau(\delta)$ $w_{\delta}$, $\widetilde{\mu}(\delta)$, $\phi(\delta)$ and the above expression into (14) and (16), we obtain
[…]
and
$g(\delta)=\begin{cases}\frac{a(q-1)q^{k-1}}{\lambda},&\hbox{if $\lambda$ is odd,}\\ \frac{2a(q-1)q^{k-1}}{\lambda},&\hbox{if $\lambda$ is even.}\end{cases}$
By Theorem 3, it follows that (59) holds. ∎
VIII Some non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$
In this section, we determine the Bose distance and dimension of certain non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$.
Theorem 7.
Let $m\geq 4$ and $b$ be two integers with $\frac{q^{m-h}-1}{\lambda}+1\leq b<\frac{q^{\lfloor({2m-1})/{3}\rfloor+1}-1}{\lambda}$. Let $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$, let $\sum\limits_{\ell=0}^{h+j_{b}}b_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda b$, and let $r_{b}$ be the smallest integer in $m-2h-j_{b},j_{b}]$ such that $b_{h+r_{b}}>0$. If $\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$, then
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta,b)})=n-m(\delta-1)$ (60)
and
$d_{B}(\mathcal{C}_{(q,n,\delta,b)})=\delta$ (61)
for any integer $\delta$ such that $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor+1$.
Proof.
We demonstrate the proof only for the case where $m$ is even, as the case $m$ is odd follows similarly. First, observe that
$\begin{split}V\left(\sum\limits_{\ell=h+r_{b}}^{h+j_{b}}b_{\ell}q^{\ell}+q^{h-j_{b}}-1\right)=(\mathbf{0}_{h-1-j_{b}},b_{h+j_{b}},\ldots,b_{h+r_{b}},\mathbf{0}_{2r_{b}},q-1,\ldots,q-1),\end{split}$
and
$V(\lambda b)\leq V(\lambda a)\leq V\left(\sum\limits_{\ell=h+r_{b}}^{h+j_{b}}b_{\ell}q^{\ell}+q^{h-j_{b}}-1\right)$
for all integers $a\in\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lceil\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rceil\right]$. Since $\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$, it follows that $V(\lambda a)$ must have the form
$V(\lambda a)=(\mathbf{0}_{h-1-j_{b}},b_{h+j_{b}},\ldots,b_{h+r_{b}},\mathbf{0}_{2r_{b}},a_{h-j_{b}-1},\ldots,a_{0})$ (62)
with $\sum\limits_{\ell=0}^{h-j_{b}-1}a_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$ for all integers $a\in\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor\right]$.
Next, we define a function $f$ for the integers in $\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor\right]$ as
$f(a)=\frac{a}{q^{t_{a}}},$
where $t_{a}$ is largest integer such that $q^{t_{a}}\mid a$. We now show that
$f(a)\in\mathcal{L}^{n}_{m}\quad\hbox{for all }a\in\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor\right]$ (63)
by considering the following cases.
Case 1. Suppose that $t_{a}=0$, i.e., $q\nmid a$. Then $f(a)=a$. It follows that $V(\lambda f(a))$ has the form given in (62). By Remarks 4 and 5, this implies that $\lambda f(a)\not\in\mathcal{B}_{j_{b}}(i)$ for any $i\in[-j_{b},j_{b}]$. Applying Theorem 1, we conclude that $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. Therefore, $\lambda f(a)\in\mathcal{L}_{m}^{\lambda n}$. By Lemma 1, this implies that $f(a)\in\mathcal{L}^{n}_{m}$.
Case 2. Suppose that $1\leq t_{a}\leq j_{b}$. In this case, we first have $f(a)\leq b-1$. Since $V(\lambda a)$ has the form given in (62), we derive
$V\left(\lambda f(a)\right)=(\mathbf{0}_{h-1-j_{b}+t_{a}},b_{h+j_{b}},\ldots,b_{h+r_{b}},\mathbf{0}_{2r_{b}},a_{h-j_{b}-1},\ldots,a_{t_{a}}).$
Furthermore, the inequality $\sum\limits_{\ell=0}^{h-j_{b}-1}a_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$ implies that $a_{j_{b}+r_{b}}>0$. It follows that $\lambda f(a)\not\in\mathcal{B}_{j_{b}-t_{a}}(i)$ for any $i\in[t_{a}-j_{b},j_{b}-t_{a}]$, and hence $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. This implies that $\lambda f(a)\in\mathcal{L}^{\lambda n}_{m}$, and hence $f(a)\in\mathcal{L}^{n}_{m}$.
Case 3. Suppose $t_{a}\geq j_{b}+1$. In this case, we have $\lambda f(a)<q^{h}$ and $\lambda f(a)q^{h}\bmod\lambda n\neq\lambda f(a)$. By Lemma 2, we have $|C_{\lambda n}(\lambda f(a))|=m$. Moreover, by applying [35, Theorem 2.3], we conclude that $\lambda f(a)$ is a coset leader modulo $\lambda n$. It follows that $f(a)\in\mathcal{L}^{n}_{m}$.
By now we have demonstrated that (63) holds. Noticing that $C_{n}(a)=C_{n}(f(a))$ and $f$ is injective, it follows that
$\begin{split}\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)\right|=\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(f(a))\right|=\sum\limits_{a=b}^{b+\delta-2}|C_{n}(f(a))|=m(\delta-1)\end{split}$
and
$\begin{split}\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)=\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(f(a))\neq\bigcup\limits_{a=b}^{b+\delta-1}C_{n}(f(a))=\bigcup\limits_{a=b}^{b+\delta-1}C_{n}(a)\end{split}$
for any integer $\delta$ such that $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor+1$. Therefore, both (60) and (61) hold. ∎
Example 1.
Applying the above theorem, we find the following optimal BCH codes.
  • Let $q=3$, $m=4$ and $\lambda=1$. We have $n=80$. The BCH code $\mathcal{C}_{(3,80,2,b)}$ has parameters $[80,76,2]$ for all integers $b\in[11,17]\cup[21,25]$.
  • Let $q=4$, $m=4$ and $\lambda=1$. We have $n=255$. The BCH code $\mathcal{C}_{(4,255,2,b)}$ has parameters $[255,251,2]$ for all integers $b\in[18,30]\cup[35,46]\cup[52,62]$.
IX Conclusion
In this paper, we investigate the dimension and Bose distance of BCH codes of length $\frac{q^{m}-1}{\lambda}$, where $\lambda$ is a positive divisor of q1q-1. Our main contribution is to provide explicit formulas for the dimension and the narrow-sense Bose distance of BCH codes of length $\frac{q^{m}-1}{\lambda}$ for a much larger range of designed distances than previously known. In addition, we extend these results to some non-narrow-sense BCH codes of length $\frac{q^{m}-1}{\lambda}$. Applying our results, we find some BCH codes with good parameters.
Appendix A Proof of Lemma 3
Proof.
It is clear that
$\{a\in[1,x]:q\nmid a\hbox{ and }\lambda\mid a+y\}=\{a\in[1,x]:\lambda\mid a+y\}-\{a\in[1,x]:q\mid a\hbox{ and }\lambda\mid a+y\}.$ (64)
Next, we show that
$\left|\{a\in[1,x]:q\mid a\hbox{ and }\lambda\mid a+y\}\right|=\left|\left\{a\in\left[1,\lfloor{x}/{q}\rfloor\right]:\lambda\mid a+y\right\}\right|.$ (65)
Notice that there exists a bijective $a\mapsto{a}/{q}$ between the integers in $[1,x]$ that are divisible by $q$ and the integers in $\left[1,\lfloor x/q\rfloor\right]$. Furthermore, for any integer $a$ such that $q\mid a$, we have
$a+y=\frac{a}{q}(q-1)+\frac{a}{q}+y.$
Since $\lambda\mid q-1$, it follows that $\lambda\mid a+y$ if and only if $\lambda\mid\frac{a}{q}+y$ for any integer $a$ such that $q\mid a$. Therefore, we can conclude that there exists a one-to-one correspondence between the two sets in (65) by mapping $a$ to a/q{a}/{q}, and hence the equality in (65) holds. With (64), it follows that
$\begin{split}\left|\{a\in[1,x]:q\nmid a\hbox{ and }\lambda\mid a+y\}\right|&=\left|\{a\in\left[\lfloor x/q\rfloor+1,x\right]:\lambda\mid a+y\}\right|\\ &=\lfloor\frac{x+y}{\lambda}\rfloor-\lfloor\frac{\lfloor{x}/{q}\rfloor+y}{\lambda}\rfloor.\end{split}$
This completes the proof. ∎
Appendix B Proof of Lemma 4
Proof.
First, we can apply a similar argument as utilized in the proof of Lemma 3 to conclude that
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\hbox{ and }q\mid\alpha\}\right|=\left|\left\{\alpha\in\big[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor\big]:\lambda\mid 2\alpha\right\}\right|.$
It follows that
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\hbox{ and }q\nmid\alpha\}\right|=\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\}\right|-\left|\left\{\alpha\in\big[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor\big]:\lambda\mid 2\alpha\right\}\right|.$ (66)
Then we distinguish the following two cases:
Case 1. Suppose that $\lambda$ is odd. Then $\lambda\mid 2\alpha$ holds if and only if $\lambda\mid\alpha.$ Therefore,
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\}\right|=\left|\{\alpha\in[x,y]:\lambda\mid\alpha\}\right|=\lfloor\frac{y}{\lambda}\rfloor-\lfloor\frac{x-1}{\lambda}\rfloor.$
Similarly, we also have
$\left|\{\alpha\in[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor]:\lambda\mid 2\alpha\}\right|=\lfloor\frac{\lfloor y/q\rfloor}{\lambda}\rfloor-\lfloor\frac{\lceil{x}/{q}\rceil-1}{\lambda}\rfloor.$
Case 2. Suppose that $\lambda$ is even. Then $\lambda\mid 2\alpha$ holds if and only if $\frac{\lambda}{2}\mid\alpha$. Consequently,
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\}\right|=\left|\left\{\alpha\in[x,y]:\frac{\lambda}{2}\mid\alpha\right\}\right|=\lfloor\frac{2y}{\lambda}\rfloor-\lfloor\frac{2x-2}{\lambda}\rfloor.$
Similarly, we also have
$\left|\{\alpha\in\big[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor\big]:\lambda\mid 2\alpha\}\right|=\lfloor\frac{2\lfloor y/q\rfloor}{\lambda}\rfloor-\lfloor\frac{2\lceil{x}/{q}\rceil-2}{\lambda}\rfloor.$
We now can derive the desired equation from equation (66) and the discussion for the above two cases. This completes the proof. ∎
Appendix C Proof of Lemma 5
Proof.
For simplicity, we set $\beta_{1}=x-\lambda\cdot\lfloor\frac{x}{\lambda}\rfloor$ and $\beta_{2}=y-\lambda\cdot\lfloor\frac{y}{\lambda}\rfloor$. Then
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+x}{\lambda}\rfloor-\lfloor\frac{t+y}{\lambda}\rfloor\right]=\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{x}{\lambda}\rfloor-\lfloor\frac{y}{\lambda}\rfloor+\lfloor\frac{t+\beta_{1}}{\lambda}\rfloor-\lfloor\frac{t+\beta_{2}}{\lambda}\rfloor\right].$ (67)
It is straightforward to verify that
[…]
where $i$ can be any integer in $1,\frac{q-1}{\lambda}-1].$ Therefore, we can obtain
$\sum\limits_{t=1}^{q-1}\lfloor\frac{t+\beta_{1}}{\lambda}\rfloor=\frac{(q-1)(\beta_{1}+1)}{\lambda}+\sum\limits_{i=1}^{\frac{q-1}{\lambda}-1}i\lambda.$
Similarly, we can also derive
$\sum\limits_{t=1}^{q-1}\lfloor\frac{t+\beta_{2}}{\lambda}\rfloor=\frac{(q-1)(\beta_{2}+1)}{\lambda}+\sum\limits_{i=1}^{\frac{q-1}{\lambda}-1}i\lambda.$
Combining the above two equalities, we obtain
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+\beta_{1}}{\lambda}\rfloor-\lfloor\frac{t+\beta_{2}}{\lambda}\rfloor\right]=\frac{(q-1)(\beta_{1}-\beta_{2})}{\lambda}.$
With the equality in (67), it follows that
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+x}{\lambda}\rfloor-\lfloor\frac{t+y}{\lambda}\rfloor\right]=\frac{q-1}{\lambda}\left(\lambda\cdot\lfloor\frac{x}{\lambda}\rfloor+\beta_{1}\right)-\frac{q-1}{\lambda}\left(\lambda\cdot\lfloor\frac{y}{\lambda}\rfloor+\beta_{2}\right)=\frac{(q-1)(x-y)}{\lambda}.$
This completes the proof. ∎
Appendix D Proof of Lemma 6
Proof.
Notice that each integer $t\in[q,aq]$ such that $q\nmid t$ admits a unique decomposition
$t=iq+j\quad\hbox{with }i\in[1,a-1]\hbox{ and }j\in[1,q-1].$
Substituting this decomposition, we have
$\begin{split}\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor&=\lfloor\frac{i+iq+j}{\lambda}\rfloor-\lfloor\frac{iq+j}{\lambda}\rfloor\\ &=\lfloor\frac{j+2i}{\lambda}\rfloor-\lfloor\frac{j+i}{\lambda}\rfloor.\end{split}$
Then by applying Lemma 5, we obtain
$\begin{split}\sum\limits_{t=q,q\nmid t}^{aq}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&=\sum\limits_{i=1}^{a-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{j+2i}{\lambda}\rfloor-\lfloor\frac{j+i}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=1}^{a-1}\frac{(q-1)i}{\lambda}\\ &=\frac{a(a-1)(q-1)}{2\lambda}.\end{split}$
This completes the proof. ∎
Appendix E Proof of Lemma 7
Proof.
Notice that each integer $t\in[0,q-2]$ admits a unique decomposition
$t=i\lambda+j\quad\hbox{with }i\in\left[0,\frac{q-1}{\lambda}-1\right]\hbox{ and }j\in[0,\lambda-1].$
Therefore,
$\begin{split}\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{2t+x}{\lambda}\rfloor-\lfloor\frac{t+x}{\lambda}\rfloor\right]&=\sum\limits_{t=0}^{q-1}\left[\lfloor\frac{2t+x}{\lambda}\rfloor-\lfloor\frac{t+x}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=0}^{\frac{q-1}{\lambda}-1}\sum\limits_{j=0}^{\lambda-1}\left[\lfloor\frac{2i\lambda+2j+x}{\lambda}\rfloor-\lfloor\frac{i\lambda+j+x}{\lambda}\rfloor\right]+\lfloor\frac{2(q-1)+x}{\lambda}\rfloor-\lfloor\frac{q-1+x}{\lambda}\rfloor\\ &=\sum\limits_{i=0}^{\frac{q-1}{\lambda}-1}\sum\limits_{j=0}^{\lambda-1}\left[i+\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor\right]+\frac{q-1}{\lambda}\\ &=\frac{(q-1)(q+1)}{2\lambda}-\frac{q-1}{2}+\frac{q-1}{\lambda}\sum\limits_{j=0}^{\lambda-1}\left[\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor\right].\end{split}$ (68)
Let $y=x-\lambda\cdot\lfloor\frac{x}{\lambda}\rfloor$. Then we have $0\leq y\leq\lambda-1$ and
$\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor=\lfloor\frac{2j+y}{\lambda}\rfloor-\lfloor\frac{j+y}{\lambda}\rfloor.$ (69)
Noting that $\lfloor\frac{j+y}{\lambda}\rfloor=0$ for each integer $j\in[0,\lambda-y-1]$, and $\lfloor\frac{j+y}{\lambda}\rfloor=1$ for each integer $j\in[\lambda-y,\lambda-1]$, we can derive
$\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{j+y}{\lambda}\rfloor=y.$ (70)
Next, we determine the value of $\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{2j+y}{\lambda}\rfloor$ through the following two cases:
Case 1. Suppose that $\lambda$ is even. Since $x$ is even, it follows that $y$ is also even. We can obtain
$\lfloor\frac{2j+y}{\lambda}\rfloor=\begin{cases}0&\hbox{if }0\leq j\leq\frac{\lambda-y}{2}-1,\\ 1&\hbox{if }\frac{\lambda-y}{2}\leq j\leq\frac{2\lambda-y}{2}-1,\\ 2&\hbox{if }\frac{2\lambda-y}{2}\leq j\leq\lambda-1.\end{cases}$
It follows that
$\begin{split}\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{2j+y}{\lambda}\rfloor&=\left(\frac{2\lambda-y}{2}-\frac{\lambda-y}{2}\right)+2\left(\lambda-\frac{2\lambda-y}{2}\right)\\ &=\frac{\lambda}{2}+y.\end{split}$
Case 2. Suppose that $\lambda$ is odd. Then we can obtain
$\lfloor\frac{2j+y}{\lambda}\rfloor=\begin{cases}0&\hbox{if }0\leq j\leq\frac{\lambda-y}{2}-1,\\ 1&\hbox{if }\frac{\lambda+1}{2}-\lfloor\frac{y+1}{2}\rfloor\leq j\leq\lambda-\lceil\frac{y+1}{2}\rceil,\\ 2&\hbox{if }\lambda-\lceil\frac{y+1}{2}\rceil+1\leq j\leq\lambda-1.\end{cases}$
Consequently, we have
$\begin{split}\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{2j+y}{\lambda}\rfloor&=\left(\lambda-\lceil\frac{y+1}{2}\rceil-\frac{\lambda+1}{2}+\lfloor\frac{y+1}{2}\rfloor+1\right)+2\left(\lceil\frac{y+1}{2}\rceil-1\right)\\ &=\frac{\lambda-1}{2}-1+\lceil\frac{y+1}{2}\rceil+\lfloor\frac{y+1}{2}\rfloor+1\\ &=\frac{\lambda-1}{2}+y.\end{split}$
From equations (69) and (70), and the discussion for the above two cases, we can conclude that
$\sum\limits_{j=0}^{\lambda-1}\left[\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor\right]=\begin{cases}\frac{\lambda}{2},&\hbox{if }\lambda\hbox{ is even,}\\ \frac{\lambda-1}{2},&\hbox{if }\lambda\hbox{ is odd.}\end{cases}$
Then by applying (68), we obtain the desired equation. This completes the proof. ∎
Appendix F Proof of Lemma 8
Proof.
Note that each integer $t\in\left[q^{k},aq^{k}\right]$ such that $q\nmid t$ can be uniquely decomposed as
$t=iq+j\quad\hbox{with }i\in[q^{k-1},aq^{k-1}-1]\hbox{ and }j\in[1,q-1].$
Therefore, by substituting this decomposition, we obtain
$\begin{split}\sum\limits_{t=q^{k},q\nmid t}^{aq^{k}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor\right]&=\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{2iq+2j}{\lambda}\rfloor-\lfloor\frac{i+iq+j}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\frac{i(q-1)}{\lambda}+\lfloor\frac{2j+2i}{\lambda}\rfloor-\lfloor\frac{j+2i}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\frac{i(q-1)^{2}}{\lambda}+\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{2j+2i}{\lambda}\rfloor-\lfloor\frac{j+2i}{\lambda}\rfloor\right].\end{split}$
It is straightforward to obtain
$\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\frac{i(q-1)^{2}}{\lambda}=\frac{q^{k-1}(q-1)^{2}(aq^{k-1}+q^{k-1}-1)}{2\lambda}.$
In addition, by applying Lemma 7, we have
$\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{2j+2i}{\lambda}\rfloor-\lfloor\frac{j+2i}{\lambda}\rfloor\right]=\begin{cases}\frac{q^{k}(q-1)(a-1)}{2\lambda}&\hbox{if }\lambda\hbox{ is odd,}\\ \frac{q^{k-1}(q-1)(q+1)(a-1)}{2\lambda}&\hbox{if }\lambda\hbox{ is even.}\end{cases}$
Combining the above three equations, we obtain the desired equation. This completes the proof. ∎
Appendix G Proof of Lemma 9
Proof.
By definition, we first have
$\sum\limits_{t=q^{k}}^{aq^{k}-1}N(t+1)=\sum\limits_{t=q^{k}}^{aq^{k}-1}t-\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor.$ (71)
Through direct computation, we obtain
$\sum\limits_{t=q^{k}}^{aq^{k}-1}t=\begin{cases}\frac{1}{2}a(a-1)&\hbox{ if }k=0,\\ \frac{1}{2}(a^{2}-1)q^{2k}-\frac{1}{2}(a-1)q^{k}&\hbox{ if }k\geq 1.\end{cases}$ (72)
We now determine the value of $\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor.$ First, note that $\lfloor t/q\rfloor=0$ for all $t\in\left[q^{k},aq^{k}-1\right]$ when $k=0$. This implies that $\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor=0$ if $k=0$. When $k\geq 1,$ the interval $\left[q^{k},q^{k+1}-1\right]$ can be partitioned as
$\left[q^{k},q^{k+1}-1\right]=\bigsqcup\limits_{i=0}^{q^{k-1}(a-1)-1}\left[q^{k}+iq,q^{k}+(i+1)q-1\right].$
Moreover, for each integer $i\in\left[0,q^{k-1}(a-1)-1\right]$ and each integer $t\in\left[q^{k}+iq,q^{k}+(i+1)q-1\right]$, we have $\lfloor t/q\rfloor=q^{k-1}+i$. Therefore,
$\begin{split}\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor&=\sum\limits_{i=0}^{q^{k-1}(a-1)-1}\sum\limits_{t=q^{k}+iq}^{q^{k}+(i+1)q-1}(q^{k-1}+i)\\ &=\frac{1}{2}(a^{2}-1)q^{2k}-\frac{1}{2}(a-1)q^{k}\end{split}$
if $k\geq 1.$ Combining (71), (72) and the value of $\sum\limits_{t=q^{k}}^{q^{k+1}-1}\lfloor t/q\rfloor$ given as above, we obtain the desired equality. ∎
Appendix H Proof of Lemma 10
Proof.
The arguments used to derive (11) and (12) are analogus, so we only demonstrate equation (11) holds through the following two cases.
Case 1. Suppose that $1\leq i\leq k$. In this case, we first have
$\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor=\frac{t(q^{2i-1}-1)}{\lambda}+\lfloor\frac{2t}{\lambda}\rfloor\quad\hbox{and}\quad\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor=\frac{t(q^{2i-2}-1)}{\lambda}+\lfloor\frac{2t}{\lambda}\rfloor.$
It follows that
$\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor=\frac{tq^{2i-2}(q-1)}{\lambda}.$
This leads to
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]=\sum\limits_{t=q^{k-i}}^{aq^{k-i}-1}\frac{tq^{2i-2}(q-1)}{\lambda}-\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}$ (73)
with $\mathcal{W}_{i}=\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\mid t\}$.
Note that $\mathcal{W}_{i}=\varnothing$ if $i=k$. Thus,
$\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}=0\quad\hbox{if }i=k.$ (74)
On the other hand, it can be easily verified that $\mathcal{W}_{i}=\{q^{k-i}+jq\mid j=0,\ldots,q^{k-i-1}(a-1)-1\}$ if $1\leq i\leq k-1.$ Therefore,
$\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}$ $=$ $\sum\limits_{j=0}^{q^{k-i-1}(a-1)-1}\frac{(q^{k-i}+qj)q^{2i-2}(q-1)}{\lambda}$ (75)
$=$ $\frac{1}{2\lambda}(q-1)(a-1)\left[(a+1)q^{2k-3}+q^{k+i-2}\right]$
if $1\leq i\leq k-1$. Additionally, by straightforward computation, we can obtain
[…]
By substituting equations (74)-(76) into equation (73), we conclude that equation (11) holds for $1\leq i\leq k$.
Case 2. Suppose that $k+1\leq i\leq 0$. Then for any integer $t\in\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ with $q$-adic expansion $\sum\limits_{\ell=0}^{k-i}t_{\ell}q^{\ell}$, we have
$\lfloor tq^{2i-1}\rfloor=\lfloor tq^{2i-2}\rfloor\cdot q+t_{-2i+1}.$ (77)
This leads to
$\begin{split}\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor=\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}(q^{\ell}-1)}{\lambda}+\frac{\lfloor tq^{2i-2}\rfloor(q-1)}{\lambda}+\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}}{\lambda}\rfloor.\end{split}$
We also have
$\begin{split}\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor=\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}(q^{\ell}-1)}{\lambda}+\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor.\end{split}$
Combining the above two equalities, we can obtain
$\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor-\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor=\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}}{\lambda}\rfloor-\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor+\frac{\lfloor tq^{2i-2}\rfloor(q-1)}{\lambda}.$
By applying Lemma 5 with $x=\sum\limits_{\ell=1}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}$ and $y=\sum\limits_{\ell=1}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor$, we have
$\sum\limits_{t_{0}=1}^{q-1}\left[\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}}{\lambda}\rfloor-\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor\right]=\frac{(q-1)t_{-2i+1}}{\lambda}.$
Noticing that the value of $\lfloor tq^{2i-2}\rfloor$ is independent of $t_{0}$, we get
$\sum\limits_{t_{0}=1}^{q-1}\frac{\lfloor tq^{2i-2}\rfloor(q-1)}{\lambda}=\frac{\lfloor tq^{2i-2}\rfloor(q-1)^{2}}{\lambda}.$
Noting that $\lfloor tq^{2i-2}\rfloor=\lfloor\frac{\lfloor tq^{2i-1}\rfloor}{q}\rfloor$ and recalling equation (77), we can add the above two sums to obtain
$\begin{split}\sum\limits_{t_{0}=1}^{q-1}\left[\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor-\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor\right]&=\frac{\lfloor tq^{2i-2}\rfloor(q-1)^{2}}{\lambda}+\frac{(q-1)t_{-2i+1}}{\lambda}\\ &=\frac{q-1}{\lambda}N(\lfloor tq^{2i-1}\rfloor+1).\end{split}$ (78)
In addition, each integer $t\in\left[q^{k-i},aq^{k-i+1}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ with $q$-adic expansion $\sum\limits_{\ell=0}^{k-i}t_{\ell}q^{\ell}$ can be uniquely decomposed as
$t=\lfloor tq^{2i-1}\rfloor\cdot q^{-2i+1}+\sum\limits_{\ell=1}^{-2i}t_{\ell}q^{\ell}+t_{0}.$
Furthermore, as $t$ ranges over integers from $q^{k-i}$ to $q^{k-i+1}-1$ that are not divisible by $q$, the value of $\lfloor tq^{2i-1}\rfloor$ ranges from $q^{k+i-1}$ to $aq^{k+i-1}-1$. Additionally, for each fixed value of $\lfloor tq^{2i-1}\rfloor$, the sum $\sum\limits_{\ell=1}^{-2i}t_{\ell}q^{\ell}$ ranges over integers from 0 to $q^{-2i}-1$, while $t_{0}$ varies from $1$ to q1q-1. Therefore, we can conclude that
$\begin{split}\sum\limits_{\begin{subarray}{c}t=q^{k-i},q\nmid t\end{subarray}}^{q^{k-i+1}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]&=\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}\sum\limits_{\sum_{\ell=1}^{-2i}t_{\ell}q^{\ell}=0}^{q^{-2i}-1}\sum\limits_{t_{0}=1}^{q-1}\left[\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor-\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor\right]\\ &=\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}\frac{q^{-2i}(q-1)}{\lambda}N(\lfloor tq^{2i-1}\rfloor+1),\end{split}$
where the second equality follows from equation (78). We can now apply Lemma 9 to conclude that (11) holds for $k+1\leq i\leq 0$.
By now, we have established equation (11) ∎
Appendix I Proof of Assertion 1
Proof.
Suppose that $m$ is odd. We first aim to show that
$\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}=\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}.$ (79)
For any integer $a\in\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k_{\delta}}a_{\ell}q^{\ell}$, we have
$V(\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell})\leq V(a)\leq V(\lambda(\delta-1)).$
Noting that
$V(\lambda(\delta-1))=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{0})$
and
$V(\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell})=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},\mathbf{0}_{h-k_{\delta}+1}),$
it follows that
$V(a)=(\mathbf{0}_{h-k_{\delta}},a_{h+k_{\delta}},\ldots,a_{0})=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},a_{h-k_{\delta}},\ldots,a_{0}).$ (80)
Recalling the definition of $s_{\delta}$, it follows that $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $a_{h+s_{\delta}}>0$. By Remark 5, this implies $a\not\in\mathcal{A}_{k_{\delta}}(i)$ for any integer $i\neq s_{\delta}$. Therefore, $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(i)=\varnothing$ for any integer $i\neq s_{\delta}.$ Then applying Theorem 1, we can conclude that (79) holds.
Now, let us count the number of integers in the set $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}.$ Recall that $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$, $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}$, and $\alpha(a)=\sum\limits_{\ell=0}^{k_{\delta}+s_{\delta}-1}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{A}_{k_{\delta}}(s_{\delta})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k_{\delta}}a_{\ell}q^{\ell}$. We conclude from equation (80) and the definition of $\mathcal{A}_{k_{\delta}}(s_{\delta})$ that an integer $a\in\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition
$a=w_{\delta}+\alpha(a)$
with $\alpha(a)\in\{\alpha\in\left[1,\mu(\delta)\right]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+w_{\delta}\}.$ Consequently, we have
$\left|\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{A}_{k_{\delta}}\cap\mathcal{D}_{\lambda}\right|=\left|\{\alpha\in\left[1,\mu(\delta)\right]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+w_{\delta}\}\right|.$
Then by applying Lemma 3, equation (19) follows.
Suppose that $m$ is even. The equality in (20) can be obtained by emplying a similar argument as above. It remains to establish equation (21). By applying Lemma 2, we can conclude that $a\in\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $V(a)$ has the form
$(\mathbf{0}_{h-k_{\delta}-1},\delta_{h+k_{\delta}},\ldots,\delta_{h},\mathbf{0}_{h-k_{\delta}-1},\delta_{h+k_{\delta}},\ldots,\delta_{h})$
with $\delta_{h}>0$, $\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$ and $\lambda\mid\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}+\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$. Since $\lambda\mid q-1$, the condition $\lambda\mid\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}+\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ is equivalent $\lambda\mid 2\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}.$ Therefore, an integer satisfying the above condition exists and is unique if and only if $\delta_{h}>0$, $\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$ and $\lambda\mid 2\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$. It follows that equation (21) holds. ∎
Appendix J Proof of Assertion 2
Proof.
Suppose that $m$ is odd. By definition, an integer $a\in\left[q^{h+k_{\delta}},\;\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)$ if and only if the following conditions are satisfied:
  • (i)
    $V(a)$ is of the form given in (7) and satisfies (5) and (6) for $k=k_{\delta}$;
  • (ii)
    $V(q^{h+k_{\delta}})\leq V(a)<V\left(\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)$.
Notice that
$V(q^{h+k_{\delta}})=(\mathbf{0}_{h-k_{\delta}},1,\mathbf{0}_{h+k_{\delta}})\quad\text{and}\quad V\left(\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},\mathbf{0}_{h-k_{\delta}+1}).$
Therefore, the form of $V(a)$ in (7) together with the inequality $a_{0}>0$ in (6) imply that condition (ii) is satisfied if and only if
$(1,\mathbf{0}_{2k_{\delta}-1})\leq(a_{h+k_{\delta}},\ldots,a_{h+i},\mathbf{0}_{k_{\delta}+i-1})<(\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1}).$
Since $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$, this is further equivalent to
$q^{k_{\delta}-i}\leq\sum_{\ell=h+i}^{h+k_{\delta}}a_{\ell}q^{\ell-h-i}<\sum_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h-i}.$
Recall the definition of the set $\mathcal{T}_{i}(\delta)$ and Remarks 13. We now can conclude that for each integer $i\in\left[-k_{\delta}+1,k_{\delta}\right]$, an integer $a\in\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition
$a=t(a)\cdot q^{h+i}+\alpha(a)$
with
$\begin{cases}\lambda\mid t(a)+\alpha(a),\\ t(a)\in\mathcal{T}_{i}(\delta),\\ 1\leq\alpha(a)\leq\lfloor t(a)\cdot q^{2i-1}\rfloor\hbox{ and }q\nmid\alpha(a).\end{cases}$
Consequently, we can apply Lemma 3 to conclude that
$\begin{split}\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left|\left\{\alpha\in[1,\lfloor tq^{2i-1}\rfloor]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+t\right\}\right|\\ &=\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].\end{split}$
Then applying Theorem 1, it follows that (22) holds.
Suppose that $m$ is even. We can first utilize a similar argument as above to obtain (23). Next, we demonstrate that equation (24) holds. By applying Lemma 2, we can conclude that an integer $a\in\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $a$ can be decomposed as decomposition
$a=\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}\cdot q^{h}+\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}$ (81)
with $q\nmid\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}$, $q^{h+k_{\delta}}\leq a<\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ and $\lambda\mid a$. It is easy to see that $q^{h+k_{\delta}}\leq a<\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ holds if and only if $q^{k_{\delta}}\leq\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$. Furthermore, since $\lambda\mid q-1$, the condition $\lambda\mid a$ is satisfied if and only if $\lambda\mid 2\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}$. Therefore, an integer $a\in\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition given in (81) with
$\begin{cases}q^{k_{\delta}}\leq\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1;\\ q\nmid\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell};\\ \lambda\mid 2\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}.\end{cases}$
Consequently,
$\begin{split}\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\left|\left\{\alpha\in\left[q^{k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1\right]:\lambda\mid 2\alpha\hbox{ and }q\nmid\alpha\right\}\right|\end{split}$
By applying Lemma 4, it follows that equation (24) holds. ∎
Appendix K Proof of Assertion 3
Proof.
The arguments used to establish (25) and (26) are analogous. Therefore, we only demonstrate that (25) holds when $m$ is odd. Using reasoning similar to that in the proof of Assertion 2, we can conclude that an integer $a\in\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition
$a=t(a)\cdot q^{h+i}+\alpha(a)$
with
$\begin{cases}\lambda\mid t(a)+\alpha(a);\\[4.30554pt] t(a)\in\left[q^{k-i},\,q^{k-i+1}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\};\\[4.30554pt] 1\leq\alpha(a)\leq\left\lfloor t(a)\cdot q^{2i-1}\right\rfloor\text{ and }q\nmid\alpha(a).\end{cases}$
Consequently, by applying Lemma 3, we derive
$\begin{split}\left|\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{t=q^{k-i},q\nmid t}^{q^{k-i+1}-1}\left|\left\{\alpha\in\left[1,\lfloor tq^{2i-1}\rfloor\right]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+t\right\}\right|\\ &=\sum\limits_{t=q^{k-i},q\nmid t}^{q^{k-i+1}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].\end{split}$
Then applying Lemma 10, we obtain equation (25).
Appendix L Proof of Assertion 4
Proof.
Following a similar approach to the proof of Assertion 3, we can obtain
$|B_{k}(0)\cap\mathcal{D}_{\lambda}|=\sum_{t=q^{k},q\nmid t}^{q^{k+1}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{\left\lfloor{t}{q}^{-1}\right\rfloor+t}{\lambda}\rfloor\right].$
Then applying Lemmas 7 and 8, the assertion follows. ∎
Appendix M Proof of Assertion 5
Proof.
By applying Lemma 2, we can conclude that an integer $a\in\mathcal{H}\cap[q^{h+k},q^{h+k+1})\cap\mathcal{D}_{\lambda}$ if and only if $a$ adimits the decomposition
$a=\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}\cdot q^{h}+\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}.$ (82)
with $a_{0}>0$, $a_{k}>0$, and $\lambda\mid a$. Since $\lambda\mid q-1$, it follows that $\lambda\mid a$ is equivalent to $\lambda\mid 2\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}$. In addition, it is straightforward to verify that the conditions $a_{0}>0$ and $a_{k}>0$ are satisfied if and only if $q^{k}\leq\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}\leq q^{k+1}-1$ and $q\nmid\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}$. Therefore, an integer $a\in\mathcal{H}\cap[q^{h+k},q^{h+k+1})\cap\mathcal{D}_{\lambda}$ if and only if $a$ has the decomposition as given in (82) with
$\begin{cases}q\nmid\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell};\\ q^{k}\leq\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}\leq q^{k+1}-1;\\ \lambda\mid 2\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}.\end{cases}$
Consequently,
$\left|\mathcal{H}\cap\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{D}_{\lambda}\right|=\left|\left\{\alpha\in\left[q^{k},q^{k+1}-1\right]:\lambda\mid 2\alpha\hbox{ and }q\nmid\alpha\right\}\right|.$
Then by applying Lemma 4, the assertion follows. ∎