### v2

# The dimension and Bose distance of some BCH codes of length $\frac{q^{m}-1}{\lambda}$

Abstract
BCH codes are important error correction codes, widely utilized due to their robust algebraic structure, multi-error correcting capability, and efficient decoding algorithms. Despite their practical importance and extensive study, their parameters, including dimension, minimum distance and Bose distance, remain largely unknown in general. This paper addresses this challenge by investigating the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$ over the finite field $\mathbb{F}_{q}$, where $\lambda$ is a positive divisor of q1q-1. Specifically, for narrow-sense BCH codes of this length with $m\geq 4$, we derive explicit formulas for their dimension for designed distance $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}+1$. We also provide explicit formulas for their Bose distance in the range $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}$. These ranges for $\delta$ are notably larger than the previously known results for this class of BCH codes. Furthermore, we extend these findings to determine the dimension and Bose distance for certain non-narrow-sense BCH codes of the same length. Several optimal linear codes can be obtained from these BCH codes.
March 10, 2026
I Introduction
Throughout this paper, let $q$ be a prime power and $\mathbb{F}_{q}$ be the finite field of order $q$. Let $\mathbb{F}_{q}^{n}$ denote the $n$-dimensional linear space over $\mathbb{F}_{q}.$ A code of length $n$ over $\mathbb{F}_{q}$ is defined as a nonempty subset of $\mathbb{F}_{q}^{n}$. In particular, an $[n,k,d]$ linear code $\mathcal{C}$ over $\mathbb{F}_{q}$ is defined as a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum distance $d$. A linear code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is said to be cyclic if $(c_{0},c_{1},\ldots,c_{n-1})\in\mathcal{C}$ implies that $(c_{n-1},c_{0},\ldots,c_{n-2})\in\mathcal{C}.$ Identify each vector $(c_{0},c_{1},\ldots,c_{n-1})\in\mathbb{F}_{q}^{n}$ with its polynomial representation
$c_{0}+c_{1}x+\cdots+c_{n-1}x^{n-1}\in\mathbb{F}_{q}[x]/(x^{n}-1),$
and each code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ with a subset of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ In this way, a code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is a cyclic code if and only if it is an ideal of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ Note that each ideal of $\mathbb{F}_{q}[x]/(x^{n}-1)$ is principal. Therefore, a cyclic code $\mathcal{C}\subseteq\mathbb{F}_{q}[x]/(x^{n}-1)$ can be generated by a monic polynomial $g(x)$, denoted as $\mathcal{C}=\left\langle g(x)\right\rangle.$ Moreover, the polynomial $g(x)$ is a divisor of $x^{n}-1$. The polynomial $g(x)$ is called the generator polynomial of $\mathcal{C}$, and $h(x)=(x^{n}-1)/{g(x)}$ is called the parity-check polynomial of $\mathcal{C}$.
Suppose that $n$ is a positive integer such that $\mathrm{gcd}(n,q)=1$. Let $m=\mathrm{ord}_{n}(q)$, i.e., the smallest integer such that $q^{m}\equiv 1\ (\mathrm{mod}\ n).$ Let $\alpha$ be a primitive element of the field $\mathbb{F}_{q^{m}}$. Then $\beta=\alpha^{\frac{q^{m}-1}{n}}$ is a primitive $n$-th root of unity. This leads to the factorization $x^{n}-1=\prod\limits_{i=0}^{n-1}(x-\beta^{i}).$ For each integer $i\in[0,n-1]$, we denote by $m_{i}(x)$ the minimal polynomial of $\beta^{i}$ over $\mathbb{F}_{q}$. A cyclic code of length $n$ over $\mathbb{F}_{q}$ is called a BCH code with designed distance $\delta$ if its generator polynomial takes the form
$\mathrm{lcm}(m_{b}(x),m_{b+1}(x),\ldots,m_{b+\delta-2}(x)),$
where $b$ and $\delta$ are integers with $2\leq\delta\leq n$, and $\mathrm{lcm}$ denotes the least common multiple of the polynomials. We denote by $\mathcal{C}_{(q,n,\delta,b)}$ such a BCH code. If $b=1$, it is called a narrow-sense BCH code, simply denoted by $\mathcal{C}_{(q,n,\delta)}$. If $n=q^{m}-1$, then it is called a primitive BCH code. Note that $\mathcal{C}_{(q,n,\delta,b)}$ and $\mathcal{C}_{(q,n,\delta^{{}^{\prime}},b)}$ may be identical even for distinct $\delta$ and $\delta^{{}^{\prime}}$. The Bose distance of $\mathcal{C}_{(q,n,\delta,b)}$, denoted by $d_{B}$ or $d_{B}(\mathcal{C}_{(q,n,\delta,b)})$, is the largest integer such that $\mathcal{C}_{(q,n,\delta,b)}=\mathcal{C}_{(q,n,d_{B},b)}$.
BCH codes were first independently discovered by Hocquenghem [16] and by Bose and Ray-Chaudhuri [1, 2]. They occupy a central place in coding theory due to their remarkable properties. First, they offer great flexibility in the choice of code parameters, enabling error correction capabilities to be tailored to specific applications. In addition, for block lengths up to a few hundred bits, many BCH codes are among the most powerful codes known for given length and dimension. In addition, efficient encoding and decoding algorithms have been developed, which make BCH codes highly practical for real-world applications.
Despite their widespread use and extensive study in the literature [4, 9, 6, 13, 15, 21, 20, 18, 19, 22, 23, 27, 30, 38, 3, 34, 12, 33, 35, 11, 32, 39], several open problems persist regarding BCH codes, especially concerning the precise determination of their dimension, minimum distance, and Bose distance. These parameters are crucial indicators of a BCH code’s performance. Specifically, a BCH code over $\mathbb{F}_{q}$ with dimension $k$ and minimum distance $d$ can transmit $k$ $q$-ary information symbols and correct up to $\left\lfloor\frac{d-1}{2}\right\rfloor$ $q$-ary symbol errors. Furthermore, the Bose distance $d_{B}$ provides a fundamental lower bound on the minimum distance, as established by the BCH bound [16, 2]. Notably, Charpin [3] conjectured that for a narrow-sense primitive BCH code, $d\leq d_{B}+4$. Hence, determining the Bose distance is also invaluable for a deeper understanding of BCH codes and their capabilities. However, as noted by Charpin [5] and Ding [10], the general determination of these parameters remains a challenging problem.
This paper is dedicated to the investigation of the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$, where $\lambda$ is a positive divisor of q1q-1. This class of BCH codes includes primitive BCH codes when $\lambda=1$. To date, these fundamental parameters are precisely known only for limited cases. Most existing results focus on the case where $\lambda=1$ and $b=1$, which corresponds to narrow-sense primitive BCH codes. For a comprehensive overview of the parameters of such codes, readers are referred to [40, 26, 28, 14, 29, 7, 8].
In contrast, for cases when $\lambda\neq 1$, the understanding is much more limited. Even for some specific cases, such as $\lambda=2$ and $\lambda=q-1$, the dimension and Bose distance are known only for a few designed distances. For general positive divisors $\lambda$ of q1q-1, Zhu et al. [41] determined the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$. Recently, Sun [31] determined the dimension and minimum distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for other specific designed distances. For convenience, we summarize in Table I some known results on the parameters of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for divisors $\lambda$ of q1q-1 with $\lambda\neq 1$. A check mark ✓indicates that the corresponding parameter is known, while a blank entry indicates that it is unknown. For non-narrow-sense BCH codes of length $(q^{m}-1)/\lambda$, the available results in the literature are even more scarce; see [26] for some related work. Readers may also refer to [10] for an excellent survey on known results regarding the parameters of BCH codes.
Table I: Known results on the parameters of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for divisors $\lambda$ of q1q-1 with $\lambda\neq 1$
$q$ $\lambda$ $m$ $\delta$ $k$ $_{B} dd$ Reference
prime power $\lambda=q-1$ $m\geq 4$ is even $\leq\delta\leq q^{{m}/{2}} [22$]
prime power $\lambda=q-1$ $m\geq 5$ is odd $\delta=aq^{(m-1)/2}+1$ with $1\leq a\leq q-1$ [26]
prime power $\lambda=q-1$ $m\geq 1$ $\delta=2$ [24]
$q\geq 3$ $\lambda=q-1$ $m\geq 1$ $\delta=3$ partially solved [24]
$q=3$ $\lambda=2$ $m\geq 3$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2}$, $i=1,2$ [24]
$q$ is odd $\lambda=2$ $m\geq 3$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2}$, $i=1,2$ [41]
$q$ is odd $\lambda=2$ $m\geq 2$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2}$, $1\leq i\leq\lfloor\frac{(m+11)}{6}\rfloor$ [25]
$q$ is odd $\lambda=2$ $m\geq 4$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lceil\frac{3m}{4}\rceil-1}+q^{\lceil\frac{m}{2}\rceil-1}+q^{\lceil\frac{m}{4}\rceil-1}+1}{2}$ [37]
$q$ is odd $\lambda=2$ $m\geq 7$, $m\equiv 3(\bmod 4)$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\frac{3m-1}{4}}+q^{\frac{m+1}{2}}+q^{\frac{m-11}{4}+t}+1}{2}$, $t=2,3,4$ [37]
$q$ is odd $\lambda=2$ $m\geq 7$, $m\not\equiv 3(\bmod 4)$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lceil\frac{3m}{4}\rceil}+1}{2}$ [37]
$q$ is odd $\lambda=2$ $m=2$ $\delta=\frac{i_{1}q+i_{0}}{2}$, $0\leq i_{1}\leq q-2$, $\max\{i_{0},1\}\leq i_{0}\leq q-1$ and $i_{0}\equiv i_{1}(\bmod 2)$ [37]
$q$ is odd $\lambda=2$ $m=3$ $\delta=\frac{i_{2}q^{2}+i_{1}q+i_{0}}{2}$, $0\leq i_{2}\leq q-2$, $i_{2}\leq i_{1}\leq q-1$, $\max\{i_{2},1\}\leq i_{0}\leq q-1$, $i_{1}+i_{0}\equiv i_{2}(\bmod 2)$ [37]
$q$ is odd $\lambda=2$ $m\geq 4$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2},1\leq i\leq\lfloor\frac{m+6}{4}\rfloor$ [37]
prime power $\lambda\mid q-1$ $m\geq 3$ is odd $2\leq\delta\leq\frac{q^{(m+1)/2}-1}{\lambda}+1$ [41]
prime power $\lambda\mid q-1$ $m\geq 4$ is even $2\leq\delta\leq\frac{q^{m/2+1}-1}{\lambda}+1$ [41]
prime power $\lambda\mid q-1$ $m\geq 2$ $\delta=\frac{(q-\lambda\ell_{0})q^{m-1-\ell_{1}}-1}{\lambda}$, $0\leq\ell_{1}\leq m-1$ and $0\leq\ell_{0}<\frac{q-1}{\lambda}$ [31]
A deep understanding of $q$-cyclotomic cosets, particularly their sizes and coset leaders, is crucial for determining the dimension and Bose distance of BCH codes. Indeed, the main challenge in determining these parameters of BCH codes often stems from the irregular distribution of coset leaders. Our previous work [40] explored the distribution of coset leaders modulo $q^{m}-1$ within the range $1,q^{\lfloor(2m-1)/3\rfloor+1}]$ and the sizes of corresponding $q$-cyclotomic cosets, enabling us to determine the Bose distance and dimension of narrow-sense primitive BCH codes $\mathcal{C}_{(q,q^{m}-1,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}$.
To extend these results to BCH codes of length $(q^{m}-1)/\lambda$, we need to investigate $q$-cyclotomic cosets modulo $(q^{m}-1)/\lambda$. A key and useful observation in this regard is that an integer $a$ is a coset leader modulo $(q^{m}-1)/\lambda$ if and only if $\lambda a$ is a coset leader modulo $q^{m}-1$, and the size of the $q$-cyclotomic coset modulo $(q^{m}-1)/\lambda$ of $a$ is equal to the size of the $q$-cyclotomic coset modulo $q^{m}-1$ of $\lambda a$. Consequently, the problem of finding coset leaders and sizes of $q$-cyclotomic coset modulo $(q^{m}-1)/\lambda$ can be reduced to identifying integers divisible by $\lambda$ that are coset leaders modulo $q^{m}-1$ and determining the sizes of their corresponding $q$-cyclotomic coset modulo $q^{m}-1$.
Building upon this crucial observation and our prior analysis of $q$-cyclotomic cosets modulo $q^{m}-1$, we successfully generalize our results on primitive BCH codes in [40] to BCH codes of length $(q^{m}-1)/\lambda$. Specifically, for any positive divisor $\lambda$ of q1q-1 and positive integer $m\geq 4$, this paper determines:
  • the dimension of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$;
  • the Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$.
It is important to note that the existing knowledge of the dimension for narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ only covers designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$ and some specific cases. Our results significantly extend this range, as evidenced by the inequality $\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}\geq\frac{q^{\lfloor(m+1)/2\rfloor}-1}{\lambda}\cdot q^{\lceil(m-4)/6\rceil}$. This implies that the range of $\delta$ for which we provide the dimension of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ is substantially larger than previously established. Additionally, we extend these results to certain non-narrow-sense BCH codes of the same length.
This paper is organized as follows. In Sections II and III, we provide essential preliminaries and review our previous results concerning $q$-cyclotomic cosets modulo $q^{m}-1$. Section IV presents several auxiliary lemmas that will be employed in subsequent sections. Based on the theoretical foundations established in Sections II through IV, we then determine the dimension of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$ in Section V and the Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ in Section VI, both by providing explicit formulas. Utilizing these formulas, we present some examples of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ and compare them with the tables of the best known linear codes maintained by Markus Grassl at http://www.codetables.de, which is called Database later in this paper. Furthermore, as an illustration of our main results, Section VII applies these formulas to compute the dimension and Bose distance of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ specifically when $\lambda\delta$ takes the form $aq^{h+k}+b$, where $h=\lfloor m/2\rfloor$, and $k,a,b$ are integers such that $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$ and $q\nmid b$. Following this, Section VIII extends our analysis to determine the dimension and Bose distance of certain non-narrow-sense BCH codes. Moreover, we identify some non-narrow-sense BCH codes that possess optimal parameters. Finally, Section IX concludes the paper.
II Preliminaries
Let $n$ be an integer such that $\gcd(n,q)=1$. For each integer $a\in[0,n-1]$, the $q$-cyclotomic coset of $a$ modulo $n$ is defined as
$C_{n}(a)=\{aq^{i}\bmod n\mid i=0,\ldots,\ell_{a}-1\},$ (1)
where $\ell_{a}$ is the smallest positive integer such that $aq^{\ell_{a}}\equiv a\pmod{n}$. It is clear that $|C_{n}(a)|=\ell_{a}$, where $|\cdot|$ denotes the size of a set. Moreover, it is well known that $|C_{n}(a)|=\ell_{a}$ is a divisor of $m=\mathrm{ord}_{n}(q)$. The smallest integer in $C_{n}(a)$ is called the coset leader of $C_{n}(a)$. For convenience, we occasionally refer to it simply as a coset leader modulo $n$. Let $\ell$ be a positive integer, and let $c,d\in[0,n-1]$ be two integers such that $c\leq d$. We define
$\mathcal{L}^{n}(c,d)=\{a\in[c,d]:a\text{ is the coset leader of }C_{n}(a)\}$
and
[…]
In particular, we simply denote $\mathcal{L}^{n}_{\ell}(0,n-1)$ by $\mathcal{L}^{n}_{\ell}$.
It is widely known that the minimal polynomial $m_{a}(x)$ of $\beta^{a}$ over $\mathbb{F}_{q}$ is given by
$m_{a}(x)=\prod\limits_{i\in C_{n}(a)}(x-\beta^{i});$
see, for example, [17, Theorem 4.1.1]. Hence, the generator polynomial $g(x)$ of the BCH code $\mathcal{C}_{(q,n,\delta,b)}$ can be expressed as
$g(x)=\prod\limits_{i\in\mathcal{G}}(x-\beta^{i}),\quad\mathcal{G}=\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a).$
Consequently, the dimension of $\mathcal{C}_{(q,n,\delta,b)}$ is
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta,b)})=n-\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)\right|,$ (2)
and the Bose distance $d_{B}$ of $\mathcal{C}_{(q,n,\delta,b)}$ is exactly the largest integer such that
$\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)=\bigcup\limits_{a=b}^{b+d_{B}-2}C_{n}(a).$ (3)
In particular, the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta)}$ can be given by
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=n-\left|\bigcup\limits_{a=1}^{\delta-1}C_{n}(a)\right|=n-\sum\limits_{a\in\mathcal{L}^{n}(1,\delta-1)}|C_{n}(a)|.$ (4)
Thus, determining the coset leaders modulo $n$ within $1,\delta-1]$ and the sizes of corresponding $q$-cyclotomic cosets allows us to compute the dimension of $\mathcal{C}_{(q,n,\delta)}$. Additionally, for $\delta^{\prime}>\delta$, one can easily verify that
$\bigcup_{a=1}^{\delta-1}C_{n}(a)=\bigcup_{a=1}^{\delta^{\prime}-1}C_{n}(a)$
if and only if none of the integers from $\delta$ to $\delta^{\prime}-1$ is a coset leader. Therefore, the Bose distance of $\mathcal{C}_{(q,n,\delta)}$ equals the smallest coset leader that is not less than $\delta$.
We have a useful observation regarding $q$-cyclotomic cosets, as presented in the following lemma. This result originated from the proof of [41, Lemma 6]. For completeness, we include a proof below.
Lemma 1.
Suppose that $n$ and $\lambda$ are two integers such that $\mathrm{gcd}(n,q)=1$ and $\mathrm{gcd}(\lambda,q)=1$. Let $a\in[0,n-1]$ be an integer. Then
  • $a$ is the coset leader of $C_{n}(a)$ if and only if $\lambda a$ is the coset leader of $C_{\lambda n}(\lambda a)$;
  • $|C_{n}(a)|=|C_{\lambda n}(\lambda a)|.$
Proof.
By definition, the integer $a$ is the coset leader of $C_{n}(a)$ if and only if
$a\bmod n\leq aq^{i}\bmod n\quad\hbox{for any integer }i\geq 0.$ (5)
Noticing that $a\bmod n=a-n\cdot\lfloor\displaystyle\frac{a}{n}\rfloor$ and $\lambda a\bmod\lambda n=\lambda a-\lambda n\cdot\lfloor\displaystyle\frac{\lambda a}{\lambda n}\rfloor$, we can assert that
$\lambda\cdot(a\bmod n)=\lambda a-\lambda n\cdot\lfloor\displaystyle\frac{a}{n}\rfloor=\lambda a-\lambda n\cdot\lfloor\displaystyle\frac{\lambda a}{\lambda n}\rfloor=\lambda a\bmod\lambda n.$
Similarly, we also have
$\lambda\cdot(aq^{i}\bmod n)=\lambda aq^{i}\bmod\lambda n.$ (6)
Therefore, the inequality in (5) is equivalent to
$\lambda a\bmod\lambda n\leq\lambda aq^{i}\bmod\lambda n\quad\hbox{for any integer }i\geq 0.$
By definition, this holds if and only if $\lambda a$ is the coset leader of $C_{\lambda n}(\lambda a)$. Therefore, the first statement follows.
In addition, the equality in (6) also implies that
$aq^{i}\bmod n=a\quad\hbox{if and only if}\quad\lambda aq^{i}\bmod\lambda n=\lambda a.$
It follows that the smallest integer $\ell_{a}$ such that $aq^{\ell_{a}}\bmod n=a$ is equal to the smallest integer $\ell_{\lambda a}$ such that $\lambda aq^{\ell_{\lambda a}}\bmod\lambda n=\lambda a$. Therefore, we have $|C_{n}(a)|=|C_{\lambda n}(\lambda a)|.$ This completes the proof. ∎
Let $\mathbb{Z}$ be the set of all integers. For each integer $\lambda$, we denote by $\mathcal{D}_{\lambda}$ the set of all the integers that are divisible by $\lambda$, that is,
$\mathcal{D}_{\lambda}=\{a\in\mathbb{Z}:\lambda\mid a\}.$
Then one can directly derive the following corollary from Lemma 1.
Corollary 1.
Let $n$ and $\lambda$ be two integers such that $\gcd(n,q)=1$ and $\mathrm{gcd}(\lambda,q)=1$. Then
$\left|\mathcal{L}_{\ell}^{n}(b,c)\right|=\left|\mathcal{L}^{\lambda n}_{\ell}(\lambda b,\lambda c)\cap\mathcal{D}_{\lambda}\right|$
for any positive integer $\ell$, and integers $b,c$ with $0\leq b\leq c\leq n-1$.
III Some known results on $q$-cyclotomic cosets modulo $q^{m}-1$
Throughout the rest of the paper, let $m$ be a positive integer, let $\lambda$ be a positive divisor of q1q-1, and set $h=\lfloor\frac{m}{2}\rfloor$ and $n=(q^{m}-1)/\lambda$. For a positive real number $a$, we denote by $N(a)$ the number of integers in the range [1,a1][1,a-1] that are not divisible by $q$. Equivalently, $N(a)=\lfloor a-1\rfloor-\lfloor(a-1)/{q}\rfloor.$
In this section, we briefly review some known results on $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$. These foundational results will be essential for deriving the main contributions of this paper. To maintain consistency, we follow the notation and terminology used in [40], which we now introduce for completeness.
Let $Z_{q}=\{0,1,2,\ldots,q-1\}$ be the set of nonnegative integers strictly less than $q$, and let $Z_{q}^{m}=\{(a_{m-1},\ldots,a_{0}):a_{i}\in Z_{q}\}$ be the set of all length-$m$ sequences of integers in $Z_{q}$. For simplicity, we denote by $\mathbf{0}_{m}$ the sequence in $Z_{q}^{m}$ whose elements are all zero. We define an order on $Z_{q}^{m}$ by lexicographic order. Specifically, for any two sequences $U=(u_{m-1},\ldots,u_{1},u_{0})$ and $W=(w_{m-1},\ldots,w_{1},w_{0})$ in $Z_{q}^{m}$, we define:
  • 1.
    $U$ is equal to $W$, denoted by $U=W$, if $u_{\ell}=w_{\ell}$ for $\ell=0,\ldots,m-1.$
  • 2.
    $U$ is less than $W$, denoted by $U<W$, if either $u_{m-1}<w_{m-1}$ or there exists an integer $i\in[0,m-2]$ such that $u_{i}<v_{i}$ and $u_{\ell}=w_{\ell}$ for all $\ell=i+1,\ldots,m-1.$
  • 3.
    $U\leq W$ is denoted if $U=W$ or $U<W$.
Note that each integer $a\in[0,\lambda n]$ can be uniquely represented by its $q$-adic expansion as $a=\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell},$ where $a_{\ell}\in Z_{q}$ for all $\ell=0,1,\dots,m-1$. We define a map $V$ from the set of integers in $0,\lambda n]$ to $Z_{q}^{m}$ as
$V(a)=(a_{m-1},\ldots,a_{1},a_{0}),$
where $\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell}$ forms the $q$-adic expansion of the integer $a\in[0,\lambda n].$
Assume that $m\geq 4$. Let $k$ be an integer such that $m-2h\leq k\leq\lfloor{(2m-1)}/{3}\rfloor-h$. If $m$ is an odd integer, for each integer $i\in[-k+1,k]$, we define $\mathcal{A}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ satisfying the following conditions:
$a_{h+i}>0;$ (7)
$(a_{k+i-1},\ldots,a_{0})\leq(a_{h+k},\ldots,a_{h-i+1})\hbox{ and }a_{0}>0;$ (8)
$V(a)=(\mathbf{0}_{h-k},a_{h+k},\ldots,a_{h+i},\mathbf{0}_{h-k},a_{k+i-1},\ldots,a_{0}).$ (9)
If $m$ is an even integer, for each integer $i\in[-k,k]$, we define $\mathcal{B}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ satisfying the condition in (7) and the following:
$(a_{k+i},\ldots,a_{0})\leq(a_{h+k},\ldots,a_{h-i})\hbox{ and }a_{0}>0;$ (10)
$V(a)=(\mathbf{0}_{h-k-1},a_{h+k},\ldots,a_{h+i},\mathbf{0}_{h-k-1},a_{k+i},\ldots,a_{0}).$ (11)
With the above preparations, we now present several known results on $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$ as follows.
Lemma 2.
[36, Theorem 2.3] If $m$ is odd, then for any integer $\leq q^{h+1}, aa$ is the coset leader of $C_{\lambda n}(a)$ if and only if $q\nmid a$. If $m$ is even, then for any integer $\leq 2q^{h}, aa$ is the coset leader of $C_{\lambda n}(a)$ if and only if $q\nmid a$.
Clearly, if an integer $a$ is divisible by $q$, then $a/q\in C_{\lambda n}(a)$. Therefore, $q\nmid a$ is a necessary condition for $a$ to be a coset leader modulo $\lambda n=q^{m}-1$. The above lemma shows that this condition is also sufficient for sufficiently small $a$. However, for larger values of $a$, there exist integers $a$ that are neither divisible by $q$ nor coset leaders modulo $\lambda n=q^{m}-1$. We denote by $\mathcal{S}$ the set such integers, that is,
[…]
In addition, we define $\mathcal{H}$ to be the set of integers in $1,\lambda n-1]$ that are coset leaders modulo $\lambda n$ and whose corresponding $q$-cyclotomic cosets have size $m/2$, i.e.,
[…]
Lemma 3.
[40, Theorem 2] Let $m$ and $k$ be two integers such that $m\geq 4$ and $m-2h\leq k\leq\lfloor(2m-1)/{3}\rfloor-h.$
  • If $m$ is odd, then
    $\mathcal{S}\cap[q^{h+k},q^{h+k+1})=\bigsqcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i).$
  • If $m$ is even, then
    $(\mathcal{S}\cup\mathcal{H})\cap[q^{h+k},q^{h+k+1})=\bigsqcup\limits_{i=-k}^{k}\mathcal{B}_{k}(i).$
In [40, Theorem 1], we determined the sizes of certain $q$-cyclotomic cosets modulo $q^{m}-1$. There is a typo in the even-$m$ case: the upper bound $q^{m-\lfloor m/4\rfloor}$ should be $q^{m-\lfloor m/3\rfloor}$. For convenience, we restate this result below with the corrected range.
Lemma 4.
[40, Theorem 1] If $m$ is an odd integer, then for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
$|C_{\lambda n}(a)|=m.$
If $m$ is an even integer, then for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
$|C_{\lambda n}(a)|=\begin{cases}\frac{m}{2}&\hbox{if }aq^{h}\mathrm{\ mod\ }\lambda n=a,\\ m&\hbox{if }aq^{h}\mathrm{\ mod\ }\lambda n\neq a.\end{cases}$
The following result characterizes the integers $a$ within certain ranges for which the $q$-cyclotomic coset $C_{\lambda n}(a)$ has size $m/2$ when $m$ is even.
Lemma 5.
[40, Corollary 2] Let $m$ be an even integer and $k\in\left[0,\lfloor(2m-1)/{3}\rfloor-h\right]$ be an integer. Suppose that $a\in\left[q^{h+k},q^{h+k+1}\right)$ is an integer. Then $a\in\mathcal{H}$ if and only if $V(a)$ has the form
$(\mathbf{0}_{h-k-1},a_{k},\ldots,a_{0},\mathbf{0}_{h-k-1},a_{k},\ldots,a_{0})$
with $a_{0}>0$ and $a_{k}>0$.
Although the above results only describe the sizes and coset leaders of $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$, by applying Lemma 1, they also yield information on $q$-cyclotomic cosets modulo $n=(q^{m}-1)/\lambda.$ These facts will be used later to determine the dimension and Bose distance of BCH codes of length $n=(q^{m}-1)/\lambda$.
To facilitate the use of these results in subsequent sections, we introduce additional notation and make several remarks concerning the definitions of the sets $\mathcal{A}_{k}(i)$ and $\mathcal{B}_{k}(i)$. For each integer $a\in\mathcal{A}_{k}(i)$, we define
$t(a)=\sum_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}\quad\text{and}\quad\alpha(a)=\sum_{\ell=0}^{k+i-1}a_{\ell}q^{\ell}.$
Similarly, for each integer $a\in\mathcal{B}_{k}(i)$, we define
$t(a)=\sum_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}\quad\text{and}\quad\alpha(a)=\sum_{\ell=0}^{k+i}a_{\ell}q^{\ell}.$
We then make the following remarks.
Remark 1.
The condition in (7) can be equivalently represented as $q\nmid t(a)$.
Remark 2.
The condition in (9) implies that each integer $a\in\mathcal{A}_{k}(i)$ can be uniquely decomposed as
$a=t(a)q^{h+i}+\alpha(a).$
Since $\lambda$ is a divisor q1q-1, it follows that for any integer $a\in\mathcal{A}_{k}(i)$,
$\lambda\mid a\quad\hbox{ if and only if }\quad\lambda\mid t(a)+\alpha(a).$
Similarly, each integer $a\in\mathcal{B}_{k}(i)$ can be uniquely decomposed as
$a=t(a)q^{h+i}+\alpha(a),$
and
$\lambda\mid a\quad\hbox{ if and only if }\quad\lambda\mid t(a)+\alpha(a).$
Remark 3.
The condition in (8) can be equivalently expressed as
$1\leq\alpha(a)\leq\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}\quad\hbox{and}\quad q\nmid\alpha(a).$ (12)
Moreover, the form of $V(a)$ in (9) implies that
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}=t(a)\cdot q^{2i-1}\quad\hbox{for all integers }i\in[1,k],$
and
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}+\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}=t(a)\cdot q^{2i-1}\quad\hbox{for all integers }i\in[-k+1,0].$
Noticing that $0\leq\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}<1$ for all integers $i\in[-k+1,0]$, we conclude that (12) is equivalent to
$1\leq\alpha(a)\leq\lfloor t(a)\cdot q^{2i-1}\rfloor\quad\hbox{and}\quad q\nmid\alpha(a).$
Similarly, the condition in (10) can also be written as
$1\leq\alpha(a)\leq\sum\limits_{\ell=h-i}^{h+k}a_{\ell}q^{\ell-(h-i)}\quad\hbox{and}\quad q\nmid\alpha(a),$
which is further equivalent to
$1\leq\alpha(a)\leq\lfloor t(a)\cdot q^{2i}\rfloor\quad\hbox{and}\quad q\nmid\alpha(a).$
Remark 4.
If $a\in\mathcal{A}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell},$ then $\sum\limits_{\ell=0}^{h-k}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}$. Similarly, if $a\in\mathcal{B}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$, then $\sum\limits_{\ell=0}^{h-k-1}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-i}^{h+k}a_{\ell}q^{\ell-(h-i)}$.
Remark 5.
If $a\in\mathcal{A}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell},$ then $i$ is the smallest integer in [k+1,k][-k+1,k] such that $a_{h+i}>0$. Consequently, $\mathcal{A}_{k}(i)\cap\mathcal{A}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k+1,k]$.
Similarly, if $a\in\mathcal{B}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$, then $i$ is the smallest integer in [k,k][-k,k] such that $a_{h+i}>0$. As a result, $\mathcal{B}_{k}(i)\cap\mathcal{B}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k,k]$.
IV Auxiliary Lemmas
The following lemmas are needed to establish the main theorems on the dimension and Bose distance of BCH codes in the subsequent sections. Their proofs are given in Appendices A–H.
Lemma 6.
Suppose that $x$ and $y$ are two positive integers. Then
$\left|\{\alpha\in[1,x]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+y\}\right|=\lfloor\frac{x+y}{\lambda}\rfloor-\lfloor\frac{\lfloor{x}/{q}\rfloor+y}{\lambda}\rfloor.$
Lemma 7.
Suppose that $x$ and $y$ are two integers such that $x\leq y$. Then
[…]
Lemma 8.
Suppose that $x$ and $y$ are two integers. Then
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+x}{\lambda}\rfloor-\lfloor\frac{t+y}{\lambda}\rfloor\right]=\frac{(q-1)(x-y)}{\lambda}.$
Lemma 9.
Suppose that $a$ is a positive integer. Then
$\sum\limits_{t=q,q\nmid t}^{aq}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\frac{a(a-1)(q-1)}{2\lambda}.$
Lemma 10.
Suppose that $x$ is an even integer. Then
[…]
Lemma 11.
Let $k$ and $a$ be two positive integers with $a\leq q$. Then
$\sum\limits_{t=q^{k},q\nmid t}^{aq^{k}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor\right]={\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}}+\textstyle{\frac{(3+(-1)^{\lambda})}{4\lambda}(a-1)(q-1)q^{k-1}}.$
Lemma 12.
Let $k$ be an integer, and let $a$ be a positive integer with $a\leq q$. Then
[…]
Lemma 13.
Let $k$ and $a$ be two positive integers with $a\leq q$. Then
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-3}&\hbox{if }-k+2\leq i\leq k-1,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-2}&\hbox{if }i=k\hbox{ or }-k+1,\end{cases}$ (13)
and
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}&\hbox{if }-k+1\leq i\leq k-1\hbox{ and }i\neq 0,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-1}&\hbox{if }i=k\hbox{ or }-k.\end{cases}$ (14)
V The dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
Suppose that $m\geq 4$. Let $\delta$ be an integer satisfying $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$. Define $k_{\delta}=\lfloor\log_{q}\lambda(\delta-1)\rfloor-h$. This implies that $q^{h+k_{\delta}}\leq\lambda(\delta-1)<q^{h+k_{\delta}+1}$. Write the $q$-adic expansion of $\lambda(\delta-1)$ as $\lambda(\delta-1)=\sum\limits_{\ell=0}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$. If $k_{\delta}\geq m-2h$, let $s_{\delta}$ denote the smallest integer in $m-2h-k_{\delta},k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$ and let $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$.
If $m$ is odd, for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$, we define the function $f(\delta)$ by
$f(\delta)=\begin{cases}0&\hbox{if }\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ {\left(\begin{aligned} &\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]\end{aligned}\right)}&\hbox{if }\delta>\frac{q^{h+1}-1}{\lambda}+1,\end{cases}$ (15)
where $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}$ and $\mathcal{T}_{i}(\delta)=\left\{t\in\mathbb{Z}:q^{k_{\delta}-i}\leq t<\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h-i}\hbox{ and }q\nmid t\right\}.$
If $m$ is even, for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$, we define the function $\widetilde{f}(\delta)$ by
$\widetilde{f}(\delta)=\begin{cases}0&\hbox{if }\delta\leq\frac{q^{h}-1}{\lambda}+1,\\ \lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&\hbox{if }\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ \left({\begin{aligned} &q^{2k_{\delta}-2}(k_{\delta}-\textstyle\frac{1}{2})\textstyle\frac{(q-1)^{2}}{\lambda}+\frac{q-1}{2\lambda}\left(q^{k_{\delta}-1}+\textstyle\frac{1+(-1)^{\lambda}}{2}\right)\\ &+\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\textstyle\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}\left[\lfloor\textstyle\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\textstyle\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]\end{aligned}}\right)&\hbox{if }\delta>\frac{q^{h+1}-1}{\lambda}+1,\end{cases}$ (16)
where $\widetilde{\mu}(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}-1}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta})}\right\}.$ Define the function $\tau(\delta)$ for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$ by
$\tau(\delta)=\begin{cases}1&\hbox{if }\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ &\delta_{h}>0\hbox{ and }\lambda\mid 2\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell},\\ 0&\hbox{otherwise.}\end{cases}$ (17)
Additionally, we define the function $g(\delta)$ for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$ by
$g(\delta)=\begin{cases}0&\hbox{if }\delta\leq\frac{q^{h}-1}{\lambda}+1,\\ \lfloor\frac{(\delta_{h}-1)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor+\tau(\delta)&\hbox{if }\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ \lfloor\frac{\phi(\delta)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor-\lfloor\frac{\lfloor\phi(\delta)/q\rfloor\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor+\tau(\delta)&\hbox{if }\frac{q^{h+1}-1}{\lambda}+1<\delta,\end{cases}$ (18)
where $\phi(\delta)=\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$.
Theorem 1.
Let $m\geq 4$ be an integer, let $\lambda$ be a positive divisor of q1q-1, and let $\delta$ be an integer such that $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ Set $n=(q^{m}-1)/\lambda$.
  • If $m$ is odd, then
    $\mathrm{dim}\bigl(\mathcal{C}_{(q,n,\delta)}\bigr)=n-m\bigl[N(\delta)-f(\delta)\bigr].$ (19)
  • If $m$ is even, then
    $\mathrm{dim}\bigl(\mathcal{C}_{(q,n,\delta)}\bigr)=n-m\bigl[N(\delta)-\widetilde{f}(\delta)\bigr]-\frac{m}{2}g(\delta).$ (20)
    Here $N(\delta)=\delta-1-\lfloor(\delta-1)/q\rfloor$ denotes the number of integers in $1,\delta-1]$ not divisible by $q$, and $f(\delta)$, $\widetilde{f}(\delta)$ and $g(\delta)$ are defined in (15), (16) and (18) respectively.
The proof of Theorem 1 proceeds via Assertions 15. Readers may find the proofs for these assertions in Appendices IM.
Assertion 1.
Suppose that $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
  • If $m$ is odd, then
    $\left|\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor.$ (21)
  • If $m$ is even, then
    $\left|\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor$ (22)
    and
    $\left|\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\tau(\delta).$ (23)
Assertion 2.
Suppose that $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
  • If $m$ is odd, then
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].$ (24)
  • If $m$ is even, then
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]$ (25)
    and
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\begin{cases}\lfloor\frac{(\delta_{h}-1)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor&\hbox{if }k_{\delta}=0,\\ \lfloor\frac{\phi(\delta)\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor-\lfloor\frac{\lfloor\phi(\delta)/q\rfloor\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor-\frac{q^{k_{\delta}-1}(q-1)\left(3+(-1)^{\lambda}\right)}{2\lambda}&\hbox{if }k_{\delta}\geq 1,\end{cases}$ (26)
where $\phi(\delta)=\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$.
Assertion 3.
Let $k$ be an integer such that $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$.
  • If $m$ is odd, then
    $|\mathcal{A}_{k}(i)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{1}{2\lambda}(q-1)^{3}(q+1)q^{2k-3}&\hbox{if }-k+2\leq i\leq k-1,\\ \frac{1}{2\lambda}(q-1)^{2}q^{2k-1}&\hbox{if }i=-k+1\hbox{ or }k.\end{cases}$ (27)
  • If $m$ is even, then
    $|\mathcal{B}_{k}(i)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{1}{2\lambda}(q-1)^{3}(q+1)q^{2k-2}&\hbox{if }-k+1\leq i\leq k-1\hbox{ and }i\neq 0,\\ \frac{1}{2\lambda}(q-1)^{2}q^{2k}&\hbox{if }i=-k\hbox{ or }k.\end{cases}$ (28)
Assertion 4.
Suppose that $m$ is even. Let $k$ be an integer such that $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
  • If $\lambda$ is odd, then
    $|B_{k}(0)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{q(q-1)}{2\lambda}&\hbox{if $k=0$,}\\ \frac{(q-1)^{2}}{2\lambda}(q^{2k}-q^{2k-2}+q^{k-1})&\hbox{if $k\geq 1$.}\end{cases}$
  • If $\lambda$ is even, then
    $|B_{k}(0)\cap\mathcal{D}_{\lambda}|=\begin{cases}\frac{(q+1)(q-1)}{2\lambda}&\hbox{if $k=0$,}\\ \frac{(q-1)^{2}}{2\lambda}(q^{2k}-q^{2k-2}+2q^{k-1})&\hbox{if $k\geq 1$.}\end{cases}$
Assertion 5.
Suppose that $m$ is even. Let $k$ be an integer such that $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
  • If $\lambda$ is odd, then
    $\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\begin{cases}\frac{q-1}{\lambda}&\hbox{if }k=0,\\ \frac{(q-1)^{2}q^{k-1}}{\lambda}&\hbox{if }k\geq 1.\end{cases}$ (29)
  • If $\lambda$ is even, then
    $\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\begin{cases}\frac{2(q-1)}{\lambda}&\hbox{if }k=0,\\ \frac{2(q-1)^{2}q^{k-1}}{\lambda}&\hbox{if }k\geq 1.\end{cases}$ (30)
Proof of Theorem 1. Suppose that $m$ is odd. We first aim to show that
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=f(\delta)\quad\hbox{for }2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ (31)
If $2\leq\delta\leq\frac{q^{h+1}-1}{\lambda}+1$, then $\lambda\leq\lambda(\delta-1)\leq q^{h+1}-1$. By applying Lemma 2, we obtain
$1,\lambda(\delta-1)]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}=0.$
In particular,
$\left|\left[1,q^{h+1}-1\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=0.$ (32)
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. This implies $1\leq k_{\delta}\leq\lfloor(2m-1)/{3}\rfloor-h$. By applying Lemma 3, we can conclude from Assertion 3 that for any integer $k\in\left[1,\lfloor({2m-1})/{3}\rfloor-h\right],$
$\begin{split}\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{i=-k+1}^{k}\left|\mathcal{A}_{k}(i)\cap\mathcal{D}_{\lambda}\right|\\ &=\frac{1}{\lambda}\left[(k-1)q^{2k-3}(q-1)^{3}(q+1)+q^{2k-1}(q-1)^{2}\right].\end{split}$
Therefore,
$\begin{split}\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{{k}=1}^{k_{\delta}-1}\left|\left[q^{h+{k}},q^{h+{k}+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|\\ &=\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{(q-1)^{3}}{\lambda}(q+1)({k}-1)q^{2{k}-3}+\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{q^{2{k}-1}(q-1)^{2}}{\lambda}.\end{split}$
It is straightforward to verify that
$\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{q^{2{k}-1}(q-1)^{2}}{\lambda}=\frac{(q^{2{k_{\delta}}-1}-q)(q-1)}{\lambda(q+1)}$
and
$\sum\limits_{k=1}^{k_{\delta}-1}\frac{(q-1)^{3}}{\lambda}(q+1)({k}-1)q^{2{k}-3}=\frac{\left[q-(k_{\delta}-1)q^{2k_{\delta}-3}+(k_{\delta}-2)q^{2k_{\delta}-1}\right](q-1)}{\lambda(q+1)}.$
By adding these two sums, we can obtain
$\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}.$ (33)
Combining equations (21), (24), (32) and (33) we obtain
$\begin{split}\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=&\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].\end{split}$
Recalling the definition of $f(\delta)$, we can now claim that equation (31) holds.
Next, we establish equation (19). Noticing that $\lfloor({2m-1})/{3}\rfloor+1\leq m-\lfloor{m}/{3}\rfloor$, we have $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. Therefore, we can apply Lemmas 1 and 4 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\quad\hbox{for all integers }a\in[1,\delta-1]$ (34)
and
$\mathcal{L}_{m}^{\lambda n}(1,\lambda(\delta-1))=\mathcal{L}^{\lambda n}(1,\lambda(\delta-1)).$
Recalling the definition of $\mathcal{S}$ and noting that an integer $a$ cannot be the coset leader of $C_{n}(a)$ if $q\mid a$, it follows that
$\left|\mathcal{L}_{m}^{\lambda n}(1,\lambda(\delta-1))\cap\mathcal{D}_{\lambda}\right|$ $=\left|\mathcal{L}^{\lambda n}(1,\lambda(\delta-1))\cap\mathcal{D}_{\lambda}\right|$
$=\left|\left\{a\in[1,\lambda(\delta-1)]:q\nmid a\right\}\cap\mathcal{D}_{\lambda}\right|-\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|.$
It can be easily verified that
$\left|\left\{a\in[1,\lambda(\delta-1)]:q\nmid a\right\}\cap\mathcal{D}_{\lambda}\right|=N(\delta).$
Utilizing Corollary 1 and equation (31), we have
$\left|\mathcal{L}_{m}^{n}(1,\delta-1)\right|=\left|\mathcal{L}_{m}^{\lambda n}(1,\lambda(\delta-1))\cap\mathcal{D}_{\lambda}\right|=N(\delta)-f(\delta).$ (35)
Recalling the equality in (4), we can now conclude from (34) and (35) that (19) holds.
Suppose that $m$ is even. Our first goal is to show that
$\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\widetilde{f}(\delta)$ (36)
and
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=g(\delta)$ (37)
for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then $\lambda\leq\lambda(\delta-1)\leq q^{h}-1$. By applying Lemma 2, we conclude that
$\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=0$
and
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=0.$
In particular, we have
$\left|\left[1,q^{h}-1\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=0$ (38)
and
$\left|\left[1,q^{h}-1\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=0.$ (39)
If $\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ then we have $q^{h}\leq\lambda(\delta-1)\leq q^{h+1}-1$. In this case, we have
  • $k_{\delta}=s_{\delta}=0$;
  • $\mathcal{T}_{0}(\delta)=\{t\in\mathbb{Z}:1\leq t\leq\delta_{h}-1\}.$
By substituting $k_{\delta}$, $s_{\delta}$ and $\mathcal{T}_{0}(\delta)$ as above into equations (22) and (25), we obtain
$\left|\left[\delta_{h}q^{h},\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor$
and
$\left|\left[q^{h},\delta_{h}q^{h}\right)\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right].$
Combining the above two equalities with (38), we obtain
$\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right].$ (40)
Additionally, by substituting $k_{\delta}=0$ into (23) and (26), we have
$\left|\left[\delta_{h}q^{h},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\tau(\delta)$
and
$\left|\left[q^{h},\delta_{h}q^{h}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor.$
Combining these two equalities with (39), we obtain
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor+\tau(\delta).$ (41)
Notice that $\lambda(\delta-1)=q^{h+1}-1=\sum\limits_{\ell=0}^{h}(q-1)q^{\ell}$ when $\delta=\frac{q^{h+1}-1}{\lambda}+1$. Therefore, for $\delta=\frac{q^{h+1}-1}{\lambda}+1$, we have
  • $\widetilde{\mu}(\delta)=q-1$;
  • $w_{\delta}=(q-1)q^{h}$;
  • $\delta_{h}=q-1.$
Consequently, by substituting $\delta=\frac{q^{h+1}-1}{\lambda}+1$ into equation (40) and applying Lemma 10, we obtain
$\begin{split}\left|\left[1,q^{h+1}-1\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|&=\lfloor\frac{(q-1)(q^{h}+1)}{\lambda}\rfloor-\lfloor\frac{(q-1)q^{h}}{\lambda}\rfloor+\sum\limits_{t=1}^{q-2}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]\\ &=\frac{q-1}{2\lambda}\left(q+\frac{1+(-1)^{\lambda}}{2}\right).\end{split}$ (42)
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then we have $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. In this case, we have $1\leq k_{\delta}\leq\lfloor({2m-1})/{3}\rfloor-h.$ By applying Lemma 3, we can conclude from Assertions 3 and 4 that
$\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{i=-k}^{k}|\mathcal{B}_{k}(i)|=\frac{(q-1)^{3}}{\lambda}(k-\frac{1}{2})q^{2k-2}(q+1)+)\frac{(q-1)^{2}}{\lambda}(\frac{1}{2}q^{k-1}+q^{2k})$
for each integer $k\in[1,\lfloor({2m-1})/{3}\rfloor-h]$. It follows that
$\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$ $=$ $\sum\limits_{k=1}^{k_{\delta}-1}\left|\left[q^{h+k},q^{h+k+1}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|$
$=$ $\frac{(q-1)^{2}}{\lambda}(k_{\delta}-\frac{1}{2})q^{2k_{\delta}-2}+\frac{q-1}{2\lambda}(q^{k_{\delta}-1}-q).$
Combining (22), (25), (42) and (V), we derive
$\begin{split}\left|\left[1,\lambda(\delta-1)\right]\cap(\mathcal{S}\cup\mathcal{H})\cap\mathcal{D}_{\lambda}\right|=&\frac{q-1}{2\lambda}\left(q^{k_{\delta}-1}+\frac{1+(-1)^{\lambda}}{2}\right)+\lfloor\frac{\widetilde{\mu}(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\widetilde{\mu}(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\frac{(q-1)^{2}}{\lambda}(k_{\delta}-\frac{1}{2})q^{2k_{\delta}-2}+\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]\end{split}$
On the other hand, we conclude from Assertion 5 that
$\left|\left[q^{h},q^{h+k_{\delta}}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\sum\limits_{k=0}^{k_{\delta}-1}\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\frac{(3+(-1)^{\lambda})(q-1)q^{k_{\delta}-1}}{2\lambda}.$
With (23) and (26), it follows that
$\left|\left[1,\lambda(\delta-1)\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\lfloor\frac{\phi(\delta)\left(3+(-1)^{\lambda}\right)}{\lambda}\rfloor-\lfloor\frac{\lfloor\phi(\delta)/q\rfloor\left(3+(-1)^{\lambda}\right)}{2\lambda}\rfloor+\tau(\delta),$
where $\phi(\delta)=\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$.
By now, we have already demonstrated that both (36) and (37) hold for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ Next, we show that equation (20) holds. Notice that $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. We can apply Lemmas 1 and 4 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\hbox{ or }\frac{m}{2}\quad\hbox{for all integers }a\in[1,\delta-1].$ (44)
By applying Corollary 1, we can derive from (36) and (37) that
$\left|\mathcal{L}^{n}_{m}(1,\delta-1)\right|=\left|\mathcal{L}^{\lambda n}_{m}\left(1,\lambda(\delta-1)\right)\cap\mathcal{D}_{\lambda}\right|=N(\delta)-\widetilde{f}(\delta)$ (45)
and
$\left|\mathcal{L}^{n}_{\frac{m}{2}}(1,\delta-1)\right|=\left|\mathcal{L}^{\lambda n}_{\frac{m}{2}}\left(1,\lambda(\delta-1)\right)\cap\mathcal{D}_{\lambda}\right|=g(\delta).$ (46)
Recalling the equality in (4), we can now conclude from (44), (45) and (46) that (20) holds. This completes the proof. ∎
Zhu et al. determined the dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for designed distance $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$ in [41, Theorem 1] for odd $m$ and in [41, Theorem 3] for even $m$, respectively. As an illustration of Theorem 1 in this paper, we present the dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for designed distances $2\leq\delta\leq\frac{q^{\lceil m/2\rceil+1}-1}{\lambda}+1$ in the following two corollaries. Notably, for even $m$, although the range of designed distances considered in Corollary 2 coincides with that in [41, Theorem 3], our derivation leads to a simpler expression, involving only four cases compared to eleven cases in [41, Theorem 3].
Corollary 2.
Let $m\geq 4$ be an even integer, and let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{h+1}-1}{\lambda}+1$. Then
[…]
Proof.
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then we have $\widetilde{f}(\delta)=0$ and $g(\delta)=0$.
If $\delta>\frac{q^{h}-1}{\lambda}+1$, then we distinguish the following two cases.
Case 1. Suppose that $\delta_{h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Then we have $\widetilde{\mu}(\delta)=\delta_{h}$ and $w_{\delta}=\delta_{h}q^{h}$. By substituting them into equation (16), we obtain
$\begin{split}\widetilde{f}(\delta)&=\lfloor\frac{\delta_{h}+\delta_{h}q^{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h}q^{h}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]\\ &=\sum\limits_{t=1}^{\delta_{h}}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right].\end{split}$
In addition, it is straightforward to verify that
$\tau(\delta)=\begin{cases}1&\hbox{if }\lambda\mid 2\delta_{h},\\ 0&\hbox{if }\lambda\nmid 2\delta_{h}.\end{cases}$
Recalling equation (18), it follows that
$g(\delta)=\begin{cases}\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor+1&\hbox{if }\lambda\mid 2\delta_{h},\\ \lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor&\hbox{if }\lambda\nmid 2\delta_{h}.\end{cases}$
Case 2. Suppose that $\delta_{h}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Then $\widetilde{\mu}(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{0}$, $w_{\delta}=\delta_{h}q^{h}$ and $\tau(\delta)=0$. By substituting these values into equations (16) and (18), we obtain
$\widetilde{f}(\delta)=\lfloor\frac{\delta_{0}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{2\delta_{h}}{\lambda}\rfloor+\sum\limits_{t=1}^{\delta_{h}}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]$
and
$g(\delta)=\lfloor\frac{\left(3+(-1)^{\lambda}\right)(\delta_{h}-1)}{2\lambda}\rfloor.$
Finally, substituting the values of $g(\delta)$ and $\widetilde{f}(\delta)$ for corresponding cases into equation (20), we derive the desired equation (47). ∎
Corollary 3.
Let $m\geq 5$ be an odd integer, and let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{h+2}-1}{\lambda}+1.$ Then
[…]
Proof.
If $\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ we directly have $f(\delta)=0$.
If $\delta>\frac{q^{h+1}-1}{\lambda}+1,$ then we have $k_{\delta}=1$. We distinguish following cases.
Case 1. Suppose that $\delta_{h}>0$. In this case, we have $s_{\delta}=0$. This leads to
  • $\mathcal{T}_{0}(\delta)=\left\{t\in\mathbb{Z}:q\leq t\leq\delta_{h+1}q+\delta_{h}-1\hbox{ and }q\nmid t\right\}$;
  • $\mathcal{T}_{1}(\delta)=\left\{t\in\mathbb{Z}:1\leq t\leq\delta_{h+1}\right\}$;
  • $w_{\delta}=\delta_{h+1}q^{h+1}+\delta_{h}q^{h}$;
  • $\mu(\delta)=\min\left\{\delta_{h+1},\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}\right\}$.
By substituting these values into equation (15), we obtain
$\begin{split}f(\delta)=&\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{{\lfloor\mu(\delta})/q\rfloor+w_{\delta}}{\lambda}\rfloor+\sum\limits_{t=q,q\nmid t}^{\delta_{h+1}q+\delta_{h}-1}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]+\sum\limits_{t=1}^{\delta_{h+1}}\left[\lfloor\frac{tq+t}{\lambda}\rfloor-\lfloor\frac{2t}{\lambda}\rfloor\right].\end{split}$ (49)
Note that $\mu(\delta)=\delta_{h+1}$ if $\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, and $\mu(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{0}$ if $\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Thus,
$\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{{\lfloor\mu(\delta})/q\rfloor+w_{\delta}}{\lambda}\rfloor=\begin{cases}\lfloor\frac{\delta_{0}+\delta_{h+1}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+\delta_{h}}{\lambda}\rfloor&\hbox{if }\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ \lfloor\frac{2\delta_{h+1}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+\delta_{h}}{\lambda}\rfloor&\hbox{if }\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$ (50)
By applying Lemma 9, we derive
$\sum\limits_{t=q,q\nmid t}^{\delta_{h+1}q}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\frac{\delta_{h+1}(\delta_{h+1}-1)(q-1)}{2\lambda}.$
Since each integer $t\in\left[\delta_{h+1}q+1,\delta_{h+1}q+\delta_{h}-1\right]$ can be uniquely expressed as $t=\delta_{h+1}q+i$ with an integer $i\in\left[1,\delta_{h}-1\right]$, we have
$\sum\limits_{t=\delta_{h+1}q+1}^{\delta_{h+1}q+\delta_{h}-1}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\sum\limits_{i=1}^{\delta_{h}-1}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right].$
Adding above two sums, we get
$\sum\limits_{t=q,q\nmid t}^{\delta_{h+1}q+\delta_{h}-1}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]=\frac{\delta_{h+1}(\delta_{h+1}-1)(q-1)}{2\lambda}+\sum\limits_{i=1}^{\delta_{h}-1}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right].$ (51)
Furthermore, it is straightforward to verify that
$\begin{split}\sum\limits_{t=1}^{\delta_{h+1}}\left[\lfloor\frac{tq+t}{\lambda}\rfloor-\lfloor\frac{2t}{\lambda}\rfloor\right]&=\sum\limits_{t=1}^{\delta_{h+1}}\frac{t(q-1)}{\lambda}\\ &=\frac{\delta_{h+1}(\delta_{h+1}+1)(q-1)}{2\lambda}.\end{split}$ (52)
We can now conclude from (49) – (52) that
$f(\delta)=\begin{cases}\lfloor\frac{\delta_{0}+\delta_{h+1}+\delta_{h}}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+\delta_{h}}{\lambda}\rfloor+\frac{\delta_{h+1}^{2}(q-1)}{\lambda}+\sum\limits_{i=1}^{\delta_{h}-1}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right]&\hbox{if }\delta_{h}>0\hbox{ and }\delta_{h+1}>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ \frac{\delta_{h+1}^{2}(q-1)}{\lambda}+\sum\limits_{i=1}^{\delta_{h}}\left[\lfloor\frac{2\delta_{h+1}+i}{\lambda}\rfloor-\lfloor\frac{\delta_{h+1}+i}{\lambda}\rfloor\right]&\hbox{if }\delta_{h}>0\hbox{ and }\delta_{h+1}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
Case 2. Suppose that $\delta_{h}=0$. Then we have $s_{\delta}=1$. This leads to
  • $\mathcal{T}_{0}(\delta)=\left\{t\in\mathbb{Z}:q\leq t\leq\delta_{h+1}q-1\hbox{ and }q\nmid t\right\}$;
  • $\mathcal{T}_{1}(\delta)=\{t\in\mathbb{Z}:1\leq t\leq\delta_{h+1}-1\}$;
  • $w_{\delta}=\delta_{h+1}q^{h+1}$;
  • $\mu(\delta)=\min\left\{\delta_{h+1}q,\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}\right\}.$
Note that $\mu(\delta)=\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}=\delta_{1}q+\delta_{0}$ if $\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$, and $\mu(\delta)=\delta_{h+1}q$ if $\delta_{h+1}q\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$. Thus, by substituting these terms into equation (15), we can use an analogous argument as in Case 1 to obtain
$f(\delta)=\begin{cases}\frac{(\delta_{h+1}^{2}-\delta_{h+1}+\delta_{1})(q-1)}{\lambda}+\lfloor\frac{\delta_{1}+\delta_{0}+\delta_{h+1}}{\lambda}\rfloor-\lfloor\frac{\delta_{1}+\delta_{h+1}}{\lambda}\rfloor&\hbox{if }\delta_{h}=0\hbox{ and }\delta_{h+1}q>\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell},\\ \frac{\delta_{h+1}^{2}(q-1)}{\lambda}&\hbox{if }\delta_{h}=0\hbox{ and }\delta_{h+1}q\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}.\end{cases}$
Finally, substituting the value of $f(\delta)$ into equation (19) for each corresponding case, we can conclude that (48) holds. This completes the proof. ∎
VI The Bose distance of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
In this section, we investigate the Bose distance of narrow-sense BCH codes of length $n=(q^{m}-1)/\lambda$ for designed distances $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$. Note that if $q\mid\delta$, then $\delta$ is not a coset leader modulo $n$. Therefore, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=d_{B}(\mathcal{C}_{(q,n,\delta+1)})$ for $q\mid\delta$. Given this property, it is sufficient to focus on BCH codes $\mathcal{C}_{(q,n,\delta)}$ with $q\nmid\delta$. Their Bose distances are established in the following three theorems.
Theorem 2.
Let $m$ be a positive integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{m-h}-1}{\lambda}$ and $q\nmid\delta$. Then $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
Proof.
Note that $\lambda\mid q-1$ implies that $\mathrm{gcd}(q,\lambda)=1$. Since $q\nmid\delta$, it follows that $q\nmid\lambda\delta$. Additionally, it is clear that $2\leq\lambda\delta<q^{m-h}$. By Lemma 2, it follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Then applying Lemma 1, we conclude that $\delta$ is a coset leader modulo $n$. Consequently, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta.$
Theorem 3.
Let $m\geq 5$ be an odd integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer with $\frac{q^{h+1}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Define $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$, and write the $q$-adic expansion of $\lambda\delta$ as $\lambda\delta=\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$. Let $r_{\delta}$ be the smallest integer in $j_{\delta}+1,\,j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and let $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$ Then
$d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(m-h-r_{\delta})},\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\hbox{ and }\delta_{h-r_{\delta}+1}+\lambda-q\neq\hat{\delta}\bmod\lambda,\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}\hbox{ and }\delta_{h-r_{\delta}+1}+\lambda-q=\hat{\delta}\bmod\lambda.\end{cases}$ (53)
Proof.
We show that equation (53) holds through the following cases.
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. By Remarks 4 and 5, the inequality implies that $\lambda\delta\not\in\mathcal{A}_{j_{\delta}}(i)$ for every integer $i\in[-j_{\delta}+1,j_{\delta}]$. Therefore, by Lemma 3, we conclude $\lambda\delta\not\in\mathcal{S}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. It follows that $\lambda\delta$ is a coset leader modulo $\lambda n$. Applying Lemma 1, we further obtain that $\delta$ is a coset leader modulo $n$. Consequently, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
Case 2. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}$. With an argument similar to that in the proof of [40, Theorem 6], we can show that each integer in $\lambda\delta,\hat{\delta}$] is not a coset leader modulo $\lambda n$. Applying Lemma 1, it follows that each integer in $\left[\delta,\lfloor\frac{\hat{\delta}}{\lambda}\rfloor\right]$ is not a coset leader modulo $n$. Notice that when $m$ is odd,
$\hat{\delta}=\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\hat{\delta}\bmod\lambda=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+1}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}.$
Thus, we have
$\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}+2}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta}+1)}+\delta_{h-r_{\delta}+1}-\hat{\delta}\bmod\lambda+\lambda.$
This implies that
$q\mid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda\quad\hbox{if and only if}\quad\delta_{h-r_{\delta}+1}+\lambda-q=\hat{\delta}\bmod\lambda.$
Then we distinguish the following two subcases:
Subcase 1. If $\delta_{h-r_{\delta}+1}+\lambda-q\neq\hat{\delta}\bmod\lambda$, then $q\nmid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$. Applying a similar argument as in Case 1, we can show that $\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$ is a coset leader modulo $\lambda n=q^{m}-1$. By applying Lemma 1, it follows that $\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1$ is a coset leader modulo $n$. Consequently, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1.$
Subcase 2. If $\delta_{h-r_{\delta}+1}+\lambda-q=\hat{\delta}\bmod\lambda$, then $q\mid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$ and $q\nmid\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2\lambda$. This implies that $\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+\lambda$ is not a coset leader of $\lambda n$. Furthermore, using a similar argument as in Case 1 again, we can conclude that while $\lambda\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2\lambda$ is a coset leader modulo $\lambda n$. By applying Lemma 1, it follows that $\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1$ is not a coset leader modulo $n$, while $\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2$ is a coset leader modulo $n$. Therefore, we have $d_{B}(\mathcal{C}_{(q,n,\delta)})=\lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2.$
Theorem 4.
Let $m\geq 4$ be an even integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer such that $\frac{q^{h}-1}{\lambda}+1\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ and $q\nmid\delta$. Define $j_{\delta}=\lfloor\log_{q}(\lambda\delta)\rfloor-h$, and write the $q$-adic expansion of $\lambda\delta$ as $\lambda\delta=\sum\limits_{\ell=0}^{h+j_{\delta}}\delta_{\ell}q^{\ell}$. Let $r_{\delta}$ be the smallest integer in $j_{\delta},j_{\delta}]$ such that $\delta_{h+r_{\delta}}>0$ and let $\hat{\delta}=\sum\limits_{\ell=h+r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell}+\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}.$
  • If $r_{\delta}\neq 0$, then
    $d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})},\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+1&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\text{ and }\delta_{h-r_{\delta}}+\lambda-q\neq\hat{\delta}\bmod\lambda,\\ \lfloor\frac{\hat{\delta}}{\lambda}\rfloor+2&\hbox{if }\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\leq\sum\limits_{\ell=h-r_{\delta}}^{h+j_{\delta}}\delta_{\ell}q^{\ell-(h-r_{\delta})}\text{ and }\delta_{h-r_{\delta}}+\lambda-q=\hat{\delta}\bmod\lambda.\end{cases}$ (54)
  • If $r_{\delta}=0$, then
[…]
Proof.
We can use an argument analogous to that in the proof of Theorem 3 to conclude that 54) holds if $r_{\delta}\neq 0$. Thus, we now only demonstrate equation (55) by considering the following cases.
Case 1. Suppose that $\sum\limits_{\ell=0}^{h-j_{\delta}-1}\delta_{\ell}q^{\ell}\geq\sum\limits_{\ell=h}^{h+j_{\delta}}\delta_{\ell}q^{\ell-h}$. If the inequality is strict, by Lemma 3 and Remarks 45, we have $\lambda\delta\notin\mathcal{S}\cup\mathcal{H}$; If equality holds, by Lemma 5, we $\lambda\delta\in\mathcal{H}$. Note that the assumptions $q\nmid\delta$ and $\lambda\mid q-1$ imply $q\nmid\lambda\delta$. Therefore, in either case, $\lambda\delta$ is a coset leader modulo $\lambda n$. By Lemma 1, it follows that $\delta$ is a coset leader of modulo $n$. Therefore, $d_{B}(\mathcal{C}_{(q,n,\delta)})=\delta$.
[…]
Subcase 1. Assume that $\lambda\mid\hat{\delta}$. Using Lemma 5, it is straightforward to verify that $\hat{\delta}\in\mathcal{H}$, and hence $\hat{\delta}$ is a coset leader modulo $\lambda n.$ It follows that $\frac{\hat{\delta}}{\lambda}$ is a coset leader modulo $n$. Therefore, we can obtain $d_{B}(\mathcal{C}_{(q,n,\delta)})=\frac{\hat{\delta}}{{\lambda}}$.
[…]
An $[n,k,d]$ linear code over $\mathbb{F}_{q}$ is called optimal if there does not exist an $n,k^{\prime},d]$ linear code with $k^{\prime}>k$, or an $n,k,d^{\prime}]$ linear code with $d^{\prime}>d$ over $\mathbb{F}_{q}$. Using the formulas given in Theorems 14, we can readily determine the dimension and Bose distance of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for designed distances $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$, where $\lambda$ is a positive divisor of q1q-1. By comparing these codes with the Database, we identify several BCH codes that are either optimal or have parameters matching those of the best-known linear codes. But most of these codes arise in the primitive case $\lambda=1$. See Tables II–III in our previous work [40] for examples of such primitive BCH codes.
In addition, we list further examples of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ with $\lambda\neq 1$ in Table II. The parameters of these codes are nearly optimal or best-known in the Database. All parameters were verified by Magma, and for all BCH codes listed in Table II, the Bose distance $d_{B}$ coincides with the minimum distance $d$. The Python programs implementing these formulas, together with additional examples, are available at https://zheng-run.github.io.
Table II: Examples of BCH code $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ with $\lambda\neq 1$
$q$ $m$ $\lambda$ $n$ $\delta$ $k$ $d_{B}$ Optimality
3 4 2 40 2 36 2 $d_{\text{optimal}}=3$
3 4 2 40 3 – 4 32 4 $d_{\text{optimal}}=5$
3 4 2 40 5 28 5 $d_{\text{best}}=6$
3 4 2 40 6 – 7 26 7 Best known
3 4 2 40 8 22 8 $d_{\text{best}}=9$
3 5 2 121 6 – 7 101 77 $d_{\text{best}}=8$
3 5 2 121 9 – 10 91 1010 $d_{\text{best}}=11$
4 4 3 85 4 – 5 73 55 $d_{\text{best}}=6$
VII BCH codes with designed distance $\delta$ satisfying $\lambda\delta=aq^{h+k}+b$.
As an illustration of Theorems 14 established in the previous two sections, we derive the following two corollaries in this section. These corollaries present the dimension and Bose distance of narrow-sense BCH code $\mathcal{C}_{(q,n,\delta)}$ for $\delta$ satisfying $\lambda\delta=aq^{h+k}+b$, where $k,a,b$ are integers such that $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$ and $q\nmid b$.
Corollary 4.
Let $m\geq 5$ be an odd integer. Let $\delta$ be an integer such that $\lambda\delta=aq^{h+k}+b$ for some integers $k$, $a$, $b$ with $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{h-k+1}$ and $q\nmid b$. Then
$d_{B}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}\delta&\hbox{if }b>aq^{2k-1},\\ \lfloor\frac{aq^{h+k}+aq^{2k-1}}{\lambda}\rfloor+1&\hbox{if }b\leq aq^{2k-1},\end{cases}$ (56)
and
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=\begin{cases}n-mN(\delta)+\frac{q-1}{\lambda}ma^{2}q^{2k-3}\left[(q-1)k+1\right]&\hbox{if }b\geq aq^{2k-1}+\lambda,\\ n-\frac{q-1}{\lambda}maq^{h+k-1}+\frac{q-1}{\lambda}maq^{2k-3}\left[a(q-1)k+a-q\right]&\hbox{if }b<aq^{2k-1}+\lambda.\end{cases}$ (57)
Proof.
Note that $q\nmid b$ implies that $q\nmid\lambda\delta$. Since $\lambda$ is a divisor of q1q-1, it follows that $q\nmid\delta$. Then by applying Theorem 3, we can directly derive equation (56).
We now demonstrate that (57) holds. It is clear that $V(\lambda(\delta-1))$ has the form
$(\mathbf{0}_{h-k},a,\mathbf{0}_{2k-1},\delta_{h-k},\ldots,\delta_{0})$
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. In this case, we have
  • $s_{\delta}=k_{\delta}=k$;
  • $w_{\delta}=aq^{h+k}$;
  • $\mu(\delta)=\min\left\{b-\lambda,aq^{2k-1}\right\}$;
  • $\mathcal{T}_{i}(\delta)=\left[q^{k-i},aq^{k-i}\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ for each integer $i\in[-k+1,k]$.
Then applying Lemma 13, we can obtain
$\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-3}&\hbox{if }-k+2\leq i\leq k-1,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-2}&\hbox{if }i=k\hbox{ or }-k+1.\end{cases}$
It follows that
$\begin{split}\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)&=\sum\limits_{i=-k+1}^{k}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)\\ &=\frac{1}{\lambda}\left[a(a-1)(q-1)q^{2k-2}+(a^{2}-1)(k-1)(q-1)^{2}q^{2k-3}\right].\end{split}$
By substituting the values of $k_{\delta}$ and $\mu(\delta)$, and the above expression into (15), we obtain
$f(\delta)=\begin{cases}\frac{q-1}{\lambda}a^{2}q^{2k-3}\left[k(q-1)+1\right]&\hbox{if }b\geq aq^{2k-1}+\lambda,\\ \frac{q-1}{\lambda}aq^{2k-3}\left[ak(q-1)+a-q\right]+\frac{b+a-\lambda}{\lambda}-\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor&\hbox{if }b<aq^{2k-1}+\lambda.\end{cases}$
Furthermore, notice that
$\begin{split}N(\delta)-\left[\frac{b+a-\lambda}{\lambda}-\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor\right]&=\frac{aq^{h+k}+b-\lambda}{\lambda}-\lfloor\frac{aq^{h+k}+b-\lambda}{\lambda q}\rfloor-\frac{b+a-\lambda}{\lambda}+\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor\\ &=\frac{aq^{h+k-1}(q-1)}{\lambda}-\lfloor\frac{aq+b-\lambda}{\lambda q}\rfloor+\lfloor\frac{\lfloor(b-\lambda+aq)/q\rfloor}{\lambda}\rfloor\\ &=\frac{aq^{h+k-1}(q-1)}{\lambda}.\end{split}$
Therefore, we can apply Theorem 1 to derive the desired equality in (57). ∎
Remark 6.
In Corollary 4, we assume $q\nmid b$. Since $\lambda$ is a divisor of q1q-1, we have $\gcd(q,\lambda)=1$. From $\lambda\delta=aq^{h+k}+b$, it follows that $q\mid b$ if and only if $q\mid\lambda\delta$, which is further equivalent to $q\mid\delta$. In this case, we have $\mathcal{C}_{(q,n,\delta)}=\mathcal{C}_{(q,n,\delta+1)},$ while clearly $q\nmid(\delta+1)$. Hence it suffices to consider the case $q\nmid b$. Similarly, in Corollary 5 below we also impose the condition $q\nmid b$.
Corollary 5.
Let $m\geq 4$ be an even integer. Let $\delta$ be an integer such that $\lambda\delta=aq^{h+k}+b$ for some integers $k,a,b$ with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{h-k}$ and $q\nmid b$.
  • If $k=0$, then
[…]
    and
    $\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})\!=\!\begin{cases}n\!-\!mN(\delta)\!+\!m\left[\frac{a+b-\lambda}{\lambda}\!-\!\lfloor\frac{2a}{\lambda}\rfloor+\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor\!-\![\lfloor\frac{t}{\lambda}\rfloor\right]\right]\!-\!\frac{m(a-1)\left(3+(-1)^{\lambda}\right)}{4\lambda}&\hskip-6.0pt\hbox{if }b<a+\lambda,\\ n\!-\!mN(\delta)+m\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor\!-\!\lfloor\frac{t}{\lambda}\rfloor\right]-\frac{m(a-1)\left(3+(-1)^{\lambda}\right)}{4\lambda}&\hskip-6.0pt\hbox{if }b\geq a+\lambda\hbox{ and }\lambda\nmid 2a,\\ n\!-\!mN(\delta)\!+\!m\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor\!-\!\lfloor\frac{t}{\lambda}\rfloor\right]-\frac{m(a-1)\left(3+(-1)^{\lambda}\right)}{4\lambda}\!-\!\frac{m}{2}&\hskip-6.0pt\hbox{if }b\geq a+\lambda\hbox{ and }\lambda\mid 2a.\end{cases}$ (59)
  • If $k\geq 1$, then
[…]
    and
[…]
Proof.
By Theorem 4, it is easy to conclude that (58) holds if $k=0$, and (60) holds if $k\geq 1$. Therefore, we only demonstrate that (59) holds if $k=0$, and (61) holds if $k\geq 1$.
If $k=0$, then we have $\lambda(\delta-1)=aq^{h}+b-\lambda$ with $0\leq b-\lambda<q^{h-1}$. In this case, we have
  • $\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1$;
  • $\delta_{h}=a$;
  • $\widetilde{\mu}(\delta)=\min\{b-\lambda,a\}$;
  • $w_{\delta}=aq^{h}$;
  • $\tau(\delta)=1$ if $a\leq b-\lambda$ and $\lambda\mid 2a$, and $\tau(\delta)=0$ if otherwise.
By substituting the above values for the corresponding case into equations (16) and (18), we derive
$\widetilde{f}(\delta)=\begin{cases}\frac{a+b-\lambda}{\lambda}-\lfloor\frac{2a}{\lambda}\rfloor+\sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&\hbox{if }b<a+\lambda,\\ \sum\limits_{t=1}^{a}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&\hbox{if }b\geq a+\lambda,\end{cases}$
and
$g(\delta)=\begin{cases}\frac{(a-1)(3+(-1)^{\lambda})}{2\lambda}+1&\hbox{if }b\geq a+\lambda\hbox{ and }\lambda\mid 2a,\\ \frac{(a-1)(3+(-1)^{\lambda})}{2\lambda}&\hbox{otherwise}.\end{cases}$
Then by Theorem 1, it follows that (59) holds.
If $k\geq 1$, then we have $k_{\delta}=k$ and $V(\lambda(\delta-1))$ has the form
$(\mathbf{0}_{h-k-1},a,\mathbf{0}_{2k},\delta_{h-k-1},\ldots,\delta_{0})$
with $\sum\limits_{\ell=0}^{h-k}\delta_{\ell}q^{\ell}=b-\lambda$. In this case, we have
  • $w_{\delta}=aq^{h}$;
  • $s_{\delta}=k_{\delta}=k$;
  • $\tau(\delta)=0$;
  • $\widetilde{\mu}(\delta)=\min\left\{b-\lambda,aq^{2k}\right\}$;
  • $\phi(\delta)=aq^{k}-1$;
  • $\mathcal{T}_{i}(\delta)=\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ for each integer $i\in[-k,k]$.
Then applying Lemmas 11 and 13, we obtain
$\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=\begin{cases}\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}&\hbox{if }-k+1\leq i\leq k-1\hbox{ and }i\neq 0,\\ \frac{1}{2\lambda}a(a-1)(q-1)q^{2k-1}&\hbox{if }i=k\hbox{ or }-k,\\ {\left(\begin{aligned} &\textstyle{\frac{1}{2\lambda}(a^{2}-1)(q-1)^{2}q^{2k-2}}\\ &+\textstyle{\frac{(3+(-1)^{\lambda})}{4\lambda}(a-1)(q-1)q^{k-1}}\end{aligned}\right)}&\hbox{if }i=0.\end{cases}$
It follows that
$\begin{split}\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor\right]=&\frac{1}{\lambda}(k-\frac{1}{2})(a^{2}-1)(q-1)^{2}q^{2k-2}+\frac{1}{\lambda}a(a-1)(q-1)q^{2k-1}\\ &+\frac{(3+(-1)^{\lambda})(a-1)q^{k-1}(q-1)}{4\lambda}.\end{split}$
By substituting $k_{\delta}$, $\tau(\delta)$ $w_{\delta}$, $\widetilde{\mu}(\delta)$, $\phi(\delta)$ and the above expression into (16) and (18), we obtain
[…]
and
$g(\delta)=\begin{cases}\frac{a(q-1)q^{k-1}}{\lambda},&\hbox{if $\lambda$ is odd,}\\ \frac{2a(q-1)q^{k-1}}{\lambda},&\hbox{if $\lambda$ is even.}\end{cases}$
By Theorem 1, it follows that (61) holds. ∎
VIII Some non-narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
In the previous sections, we study the parameters of narrow-sense BCH codes of length $n=(q^{m}-1)/\lambda$. For the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta)}$, the equality in (4) allows us to compute $\dim(\mathcal{C}_{(q,n,\delta)})$ by determining the coset leaders modulo $n$ within the range $1,\delta-1]$ and the sizes of their corresponding $q$-cyclotomic cosets. However, this approach cannot be directly applied to non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$ with $b\geq 2$, since we do not have an analogue of (4) for these codes. In general, for $b\geq 2$, we do not have
$\dim(\mathcal{C}_{(q,n,\delta,b)})=n-\sum_{a\in\mathcal{L}^{n}(b,b+\delta-2)}|C_{n}(a)|,$
because the equality $\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)\right|=\sum\limits_{a\in\mathcal{L}^{n}(b,b+\delta-2)}|C_{n}(a)|$ may not hold when $b\geq 2$. Consequently, determining the dimension of non-narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta,b)}$ with $b\geq 2$ is in general more difficult, and there are very few results on the parameters of such codes in the literature.
Nevertheless, although the above equality does not hold in general, the results on the coset leaders and sizes of $q$-cyclotomic cosets given in Sections II and III enable us to determine the dimension and Bose distance of certain classes of non-narrow-sense BCH codes, as shown in the following result.
Theorem 5.
Let $m\geq 4$ be an integer, and let $\lambda$ be a positive divisor of q1q-1. Set $n=(q^{m}-1)/\lambda$ and $h=\lfloor\frac{m}{2}\rfloor$. Let $b$ be an integer with $\frac{q^{m-h}-1}{\lambda}+1\leq b<\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}.$ Define $j_{b}=\lfloor\log_{q}(\lambda b)\rfloor-h,$ and write the $q$-adic expansion of $\lambda b$ as $\lambda b=\sum\limits_{\ell=0}^{h+j_{b}}b_{\ell}q^{\ell}.$ Let $r_{b}$ be the smallest integer in $m-2h-j_{b},\;j_{b}\,]$ such that $b_{h+r_{b}}>0$. If $\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}\;>\;\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})},$ then for any integer $\delta$ with $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\left\lfloor\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}/\lambda\right\rfloor+1,$ we have
$\dim(\mathcal{C}_{(q,n,\delta,b)})=n-m(\delta-1)$ (62)
and
$d_{B}(\mathcal{C}_{(q,n,\delta,b)})=\delta.$ (63)
Proof.
We demonstrate the proof only for the case where $m$ is even, as the case $m$ is odd follows similarly. First, observe that
$\begin{split}V\left(\sum\limits_{\ell=h+r_{b}}^{h+j_{b}}b_{\ell}q^{\ell}+q^{h-j_{b}}-1\right)=(\mathbf{0}_{h-1-j_{b}},b_{h+j_{b}},\ldots,b_{h+r_{b}},\mathbf{0}_{2r_{b}},q-1,\ldots,q-1),\end{split}$
and
$V(\lambda b)\leq V(\lambda a)\leq V\left(\sum\limits_{\ell=h+r_{b}}^{h+j_{b}}b_{\ell}q^{\ell}+q^{h-j_{b}}-1\right)$
for all integers $a\in\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lceil\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rceil\right]$. Since $\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$, it follows that $V(\lambda a)$ must have the form
$V(\lambda a)=(\mathbf{0}_{h-1-j_{b}},b_{h+j_{b}},\ldots,b_{h+r_{b}},\mathbf{0}_{2r_{b}},a_{h-j_{b}-1},\ldots,a_{0})$ (64)
with $\sum\limits_{\ell=0}^{h-j_{b}-1}a_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$ for all integers $a\in\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor\right]$.
Next, we define a function $f$ for the integers in $\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor\right]$ as
$f(a)=\frac{a}{q^{t_{a}}},$
where $t_{a}$ is largest integer such that $q^{t_{a}}\mid a$. Recall that $\mathcal{L}^{n}_{\ell}(0,n-1)$ is simply denoted by $\mathcal{L}^{n}_{\ell}$. We now show that
$f(a)\in\mathcal{L}^{n}_{m}\quad\hbox{for all }a\in\left[b,b+\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor\right]$ (65)
by considering the following cases.
Case 1. Suppose that $t_{a}=0$, i.e., $q\nmid a$. Then $f(a)=a$. It follows that $V(\lambda f(a))$ has the form given in (64). By Remarks 4 and 5, this implies that $\lambda f(a)\not\in\mathcal{B}_{j_{b}}(i)$ for any $i\in[-j_{b},j_{b}]$. Applying Lemma 3, we conclude that $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. Therefore, $\lambda f(a)\in\mathcal{L}_{m}^{\lambda n}$. By Lemma 1, this implies that $f(a)\in\mathcal{L}^{n}_{m}$.
Case 2. Suppose that $1\leq t_{a}\leq j_{b}$. In this case, we first have $f(a)\leq b-1$. Since $V(\lambda a)$ has the form given in (64), we derive
$V\left(\lambda f(a)\right)=(\mathbf{0}_{h-1-j_{b}+t_{a}},b_{h+j_{b}},\ldots,b_{h+r_{b}},\mathbf{0}_{2r_{b}},a_{h-j_{b}-1},\ldots,a_{t_{a}}).$
Furthermore, the inequality $\sum\limits_{\ell=0}^{h-j_{b}-1}a_{\ell}q^{\ell}>\sum\limits_{\ell=h-r_{b}}^{h+j_{b}}b_{\ell}q^{\ell-(h-r_{b})}$ implies that $a_{j_{b}+r_{b}}>0$. It follows that $\lambda f(a)\not\in\mathcal{B}_{j_{b}-t_{a}}(i)$ for every integer $i\in[t_{a}-j_{b},j_{b}-t_{a}]$, and hence $\lambda f(a)\not\in\mathcal{H}\cup\mathcal{S}$. This implies that $\lambda f(a)\in\mathcal{L}^{\lambda n}_{m}$, and hence $f(a)\in\mathcal{L}^{n}_{m}$.
Case 3. Suppose $t_{a}\geq j_{b}+1$. In this case, we have $\lambda f(a)<q^{h}$ and $\lambda f(a)q^{h}\bmod\lambda n\neq\lambda f(a)$. By Lemma 4, we have $|C_{\lambda n}(\lambda f(a))|=m$. Moreover, by applying Lemma 2, we conclude that $\lambda f(a)$ is a coset leader modulo $\lambda n$. It follows that $f(a)\in\mathcal{L}^{n}_{m}$.
By now we have demonstrated that (65) holds. Note that $C_{n}(a)=C_{n}(f(a))$ and $f$ is injective. Therefore, we have
$\begin{split}\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)\right|=\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(f(a))\right|=\sum\limits_{a=b}^{b+\delta-2}|C_{n}(f(a))|=m(\delta-1)\end{split}$
and
$\begin{split}\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)=\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(f(a))\neq\bigcup\limits_{a=b}^{b+\delta-1}C_{n}(f(a))=\bigcup\limits_{a=b}^{b+\delta-1}C_{n}(a)\end{split}$
for any integer $\delta$ such that $2\leq\delta\leq\frac{q^{h-j_{b}}-1}{\lambda}-\lfloor\frac{\sum\limits_{\ell=0}^{h-j_{b}-1}b_{\ell}q^{\ell}}{\lambda}\rfloor+1$. Recalling equation (2) and the fact that the Bose distance $d_{B}$ of $\mathcal{C}_{(q,n,\delta,b)}$ is the largest integer for which (3) holds, we deduce that (62) and (63) both hold. ∎
Applying the above theorem, we find the following BCH codes.
Example 1.
Let $q=3$, $m=4$ and $\lambda=1$. We have $n=80$. The BCH code $\mathcal{C}_{(3,80,2,b)}$ has parameters $[80,76,2]$ for all integers $b\in[11,17]\cup[21,25]$.
Example 2.
Let $q=4$, $m=4$ and $\lambda=1$. We have $n=255$. The BCH code $\mathcal{C}_{(4,255,2,b)}$ has parameters $[255,251,2]$ for all integers $b\in[18,30]\cup[35,46]\cup[52,62]$.
By the Hamming bound, no linear code over $\mathbb{F}_{3}$ with parameters $[80,76,d]$ exists for $d>32$, nor does a linear code over $\mathbb{F}_{4}$ with parameters $[255,251,d]$ exist for $d>2$. Therefore, the BCH codes in the above examples are optimal. Readers can also find the bounds for the minimum distance of linear codes of length up to 256 in the Database.
IX Conclusion
In this paper, we investigate the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$, where $\lambda$ is a positive divisor of q1q-1. Our main contribution is to provide explicit formulas for the dimension and Bose distance of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$ for a much larger range of designed distances than previously known. In addition, we extend these results to some non-narrow-sense BCH codes of the same length. Applying our results, we find some BCH codes with good parameters.
Appendix A Proof of Lemma 6
Proof.
It is clear that
$\{a\in[1,x]:q\nmid a\hbox{ and }\lambda\mid a+y\}=\{a\in[1,x]:\lambda\mid a+y\}-\{a\in[1,x]:q\mid a\hbox{ and }\lambda\mid a+y\}.$ (66)
Next, we show that
$\left|\{a\in[1,x]:q\mid a\hbox{ and }\lambda\mid a+y\}\right|=\left|\left\{a\in\left[1,\lfloor{x}/{q}\rfloor\right]:\lambda\mid a+y\right\}\right|.$ (67)
Notice that there exists a bijective $a\mapsto{a}/{q}$ between the integers in $[1,x]$ that are divisible by $q$ and the integers in $\left[1,\lfloor x/q\rfloor\right]$. Furthermore, for any integer $a$ such that $q\mid a$, we have
$a+y=\frac{a}{q}(q-1)+\frac{a}{q}+y.$
Since $\lambda\mid q-1$, it follows that $\lambda\mid a+y$ if and only if $\lambda\mid\frac{a}{q}+y$ for any integer $a$ such that $q\mid a$. Therefore, we can conclude that there exists a one-to-one correspondence between the two sets in (67) by mapping $a$ to a/q{a}/{q}, and hence the equality in (67) holds. With (66), it follows that
$\begin{split}\left|\{a\in[1,x]:q\nmid a\hbox{ and }\lambda\mid a+y\}\right|&=\left|\{a\in\left[\lfloor x/q\rfloor+1,x\right]:\lambda\mid a+y\}\right|\\ &=\lfloor\frac{x+y}{\lambda}\rfloor-\lfloor\frac{\lfloor{x}/{q}\rfloor+y}{\lambda}\rfloor.\end{split}$
This completes the proof. ∎
Appendix B Proof of Lemma 7
Proof.
First, we can apply a similar argument as utilized in the proof of Lemma 6 to conclude that
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\hbox{ and }q\mid\alpha\}\right|=\left|\left\{\alpha\in\big[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor\big]:\lambda\mid 2\alpha\right\}\right|.$
It follows that
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\hbox{ and }q\nmid\alpha\}\right|=\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\}\right|-\left|\left\{\alpha\in\big[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor\big]:\lambda\mid 2\alpha\right\}\right|.$ (68)
Then we distinguish the following two cases:
Case 1. Suppose that $\lambda$ is odd. Then $\lambda\mid 2\alpha$ holds if and only if $\lambda\mid\alpha.$ Therefore,
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\}\right|=\left|\{\alpha\in[x,y]:\lambda\mid\alpha\}\right|=\lfloor\frac{y}{\lambda}\rfloor-\lfloor\frac{x-1}{\lambda}\rfloor.$
Similarly, we also have
$\left|\{\alpha\in[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor]:\lambda\mid 2\alpha\}\right|=\lfloor\frac{\lfloor y/q\rfloor}{\lambda}\rfloor-\lfloor\frac{\lceil{x}/{q}\rceil-1}{\lambda}\rfloor.$
Case 2. Suppose that $\lambda$ is even. Then $\lambda\mid 2\alpha$ holds if and only if $\frac{\lambda}{2}\mid\alpha$. Consequently,
$\left|\{\alpha\in[x,y]:\lambda\mid 2\alpha\}\right|=\left|\left\{\alpha\in[x,y]:\frac{\lambda}{2}\mid\alpha\right\}\right|=\lfloor\frac{2y}{\lambda}\rfloor-\lfloor\frac{2x-2}{\lambda}\rfloor.$
Similarly, we also have
$\left|\{\alpha\in\big[\lceil{x}/{q}\rceil,\lfloor{y}/{q}\rfloor\big]:\lambda\mid 2\alpha\}\right|=\lfloor\frac{2\lfloor y/q\rfloor}{\lambda}\rfloor-\lfloor\frac{2\lceil{x}/{q}\rceil-2}{\lambda}\rfloor.$
We now can derive the desired equation from equation (68) and the discussion for the above two cases. This completes the proof. ∎
Appendix C Proof of Lemma 8
Proof.
For simplicity, we set $\beta_{1}=x-\lambda\cdot\lfloor\frac{x}{\lambda}\rfloor$ and $\beta_{2}=y-\lambda\cdot\lfloor\frac{y}{\lambda}\rfloor$. Then
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+x}{\lambda}\rfloor-\lfloor\frac{t+y}{\lambda}\rfloor\right]=\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{x}{\lambda}\rfloor-\lfloor\frac{y}{\lambda}\rfloor+\lfloor\frac{t+\beta_{1}}{\lambda}\rfloor-\lfloor\frac{t+\beta_{2}}{\lambda}\rfloor\right].$ (69)
It is straightforward to verify that
[…]
where $i$ can be any integer in $1,\frac{q-1}{\lambda}-1].$ Therefore, we can obtain
$\sum\limits_{t=1}^{q-1}\lfloor\frac{t+\beta_{1}}{\lambda}\rfloor=\frac{(q-1)(\beta_{1}+1)}{\lambda}+\sum\limits_{i=1}^{\frac{q-1}{\lambda}-1}i\lambda.$
Similarly, we can also derive
$\sum\limits_{t=1}^{q-1}\lfloor\frac{t+\beta_{2}}{\lambda}\rfloor=\frac{(q-1)(\beta_{2}+1)}{\lambda}+\sum\limits_{i=1}^{\frac{q-1}{\lambda}-1}i\lambda.$
Combining the above two equalities, we obtain
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+\beta_{1}}{\lambda}\rfloor-\lfloor\frac{t+\beta_{2}}{\lambda}\rfloor\right]=\frac{(q-1)(\beta_{1}-\beta_{2})}{\lambda}.$
With the equality in (69), it follows that
$\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{t+x}{\lambda}\rfloor-\lfloor\frac{t+y}{\lambda}\rfloor\right]=\frac{q-1}{\lambda}\left(\lambda\cdot\lfloor\frac{x}{\lambda}\rfloor+\beta_{1}\right)-\frac{q-1}{\lambda}\left(\lambda\cdot\lfloor\frac{y}{\lambda}\rfloor+\beta_{2}\right)=\frac{(q-1)(x-y)}{\lambda}.$
This completes the proof. ∎
Appendix D Proof of Lemma 9
Proof.
Notice that each integer $t\in[q,aq]$ such that $q\nmid t$ admits a unique decomposition
$t=iq+j\quad\hbox{with }i\in[1,a-1]\hbox{ and }j\in[1,q-1].$
Substituting this decomposition, we have
$\begin{split}\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor&=\lfloor\frac{i+iq+j}{\lambda}\rfloor-\lfloor\frac{iq+j}{\lambda}\rfloor\\ &=\lfloor\frac{j+2i}{\lambda}\rfloor-\lfloor\frac{j+i}{\lambda}\rfloor.\end{split}$
Then by applying Lemma 8, we obtain
$\begin{split}\sum\limits_{t=q,q\nmid t}^{aq}\left[\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{t}{\lambda}\rfloor\right]&=\sum\limits_{i=1}^{a-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{j+2i}{\lambda}\rfloor-\lfloor\frac{j+i}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=1}^{a-1}\frac{(q-1)i}{\lambda}\\ &=\frac{a(a-1)(q-1)}{2\lambda}.\end{split}$
This completes the proof. ∎
Appendix E Proof of Lemma 10
Proof.
Notice that each integer $t\in[0,q-2]$ admits a unique decomposition
$t=i\lambda+j\quad\hbox{with }i\in\left[0,\frac{q-1}{\lambda}-1\right]\hbox{ and }j\in[0,\lambda-1].$
Therefore,
$\begin{split}\sum\limits_{t=1}^{q-1}\left[\lfloor\frac{2t+x}{\lambda}\rfloor-\lfloor\frac{t+x}{\lambda}\rfloor\right]&=\sum\limits_{t=0}^{q-1}\left[\lfloor\frac{2t+x}{\lambda}\rfloor-\lfloor\frac{t+x}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=0}^{\frac{q-1}{\lambda}-1}\sum\limits_{j=0}^{\lambda-1}\left[\lfloor\frac{2i\lambda+2j+x}{\lambda}\rfloor-\lfloor\frac{i\lambda+j+x}{\lambda}\rfloor\right]+\lfloor\frac{2(q-1)+x}{\lambda}\rfloor-\lfloor\frac{q-1+x}{\lambda}\rfloor\\ &=\sum\limits_{i=0}^{\frac{q-1}{\lambda}-1}\sum\limits_{j=0}^{\lambda-1}\left[i+\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor\right]+\frac{q-1}{\lambda}\\ &=\frac{(q-1)(q+1)}{2\lambda}-\frac{q-1}{2}+\frac{q-1}{\lambda}\sum\limits_{j=0}^{\lambda-1}\left[\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor\right].\end{split}$ (70)
Let $y=x-\lambda\cdot\lfloor\frac{x}{\lambda}\rfloor$. Then we have $0\leq y\leq\lambda-1$ and
$\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor=\lfloor\frac{2j+y}{\lambda}\rfloor-\lfloor\frac{j+y}{\lambda}\rfloor.$ (71)
Noting that $\lfloor\frac{j+y}{\lambda}\rfloor=0$ for each integer $j\in[0,\lambda-y-1]$, and $\lfloor\frac{j+y}{\lambda}\rfloor=1$ for each integer $j\in[\lambda-y,\lambda-1]$, we can derive
$\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{j+y}{\lambda}\rfloor=y.$ (72)
Next, we determine the value of $\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{2j+y}{\lambda}\rfloor$ through the following two cases:
Case 1. Suppose that $\lambda$ is even. Since $x$ is even, it follows that $y$ is also even. We can obtain
$\lfloor\frac{2j+y}{\lambda}\rfloor=\begin{cases}0&\hbox{if }0\leq j\leq\frac{\lambda-y}{2}-1,\\ 1&\hbox{if }\frac{\lambda-y}{2}\leq j\leq\frac{2\lambda-y}{2}-1,\\ 2&\hbox{if }\frac{2\lambda-y}{2}\leq j\leq\lambda-1.\end{cases}$
It follows that
$\begin{split}\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{2j+y}{\lambda}\rfloor&=\left(\frac{2\lambda-y}{2}-\frac{\lambda-y}{2}\right)+2\left(\lambda-\frac{2\lambda-y}{2}\right)\\ &=\frac{\lambda}{2}+y.\end{split}$
Case 2. Suppose that $\lambda$ is odd. Then we can obtain
$\lfloor\frac{2j+y}{\lambda}\rfloor=\begin{cases}0&\hbox{if }0\leq j\leq\frac{\lambda-y}{2}-1,\\ 1&\hbox{if }\frac{\lambda+1}{2}-\lfloor\frac{y+1}{2}\rfloor\leq j\leq\lambda-\lceil\frac{y+1}{2}\rceil,\\ 2&\hbox{if }\lambda-\lceil\frac{y+1}{2}\rceil+1\leq j\leq\lambda-1.\end{cases}$
Consequently, we have
$\begin{split}\sum\limits_{j=0}^{\lambda-1}\lfloor\frac{2j+y}{\lambda}\rfloor&=\left(\lambda-\lceil\frac{y+1}{2}\rceil-\frac{\lambda+1}{2}+\lfloor\frac{y+1}{2}\rfloor+1\right)+2\left(\lceil\frac{y+1}{2}\rceil-1\right)\\ &=\frac{\lambda-1}{2}-1+\lceil\frac{y+1}{2}\rceil+\lfloor\frac{y+1}{2}\rfloor+1\\ &=\frac{\lambda-1}{2}+y.\end{split}$
From equations (71) and (72), and the discussion for the above two cases, we can conclude that
$\sum\limits_{j=0}^{\lambda-1}\left[\lfloor\frac{2j+x}{\lambda}\rfloor-\lfloor\frac{j+x}{\lambda}\rfloor\right]=\begin{cases}\frac{\lambda}{2},&\hbox{if }\lambda\hbox{ is even,}\\ \frac{\lambda-1}{2},&\hbox{if }\lambda\hbox{ is odd.}\end{cases}$
Then by applying (70), we obtain the desired equation. This completes the proof. ∎
Appendix F Proof of Lemma 11
Proof.
Note that each integer $t\in\left[q^{k},aq^{k}\right]$ such that $q\nmid t$ can be uniquely decomposed as
$t=iq+j\quad\hbox{with }i\in[q^{k-1},aq^{k-1}-1]\hbox{ and }j\in[1,q-1].$
Therefore, by substituting this decomposition, we obtain
$\begin{split}\sum\limits_{t=q^{k},q\nmid t}^{aq^{k}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{-1}\rfloor+t}{\lambda}\rfloor\right]&=\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{2iq+2j}{\lambda}\rfloor-\lfloor\frac{i+iq+j}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\frac{i(q-1)}{\lambda}+\lfloor\frac{2j+2i}{\lambda}\rfloor-\lfloor\frac{j+2i}{\lambda}\rfloor\right]\\ &=\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\frac{i(q-1)^{2}}{\lambda}+\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{2j+2i}{\lambda}\rfloor-\lfloor\frac{j+2i}{\lambda}\rfloor\right].\end{split}$
It is straightforward to obtain
$\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\frac{i(q-1)^{2}}{\lambda}=\frac{q^{k-1}(q-1)^{2}(aq^{k-1}+q^{k-1}-1)}{2\lambda}.$
In addition, by applying Lemma 10, we have
$\sum\limits_{i=q^{k-1}}^{aq^{k-1}-1}\sum\limits_{j=1}^{q-1}\left[\lfloor\frac{2j+2i}{\lambda}\rfloor-\lfloor\frac{j+2i}{\lambda}\rfloor\right]=\begin{cases}\frac{q^{k}(q-1)(a-1)}{2\lambda}&\hbox{if }\lambda\hbox{ is odd,}\\ \frac{q^{k-1}(q-1)(q+1)(a-1)}{2\lambda}&\hbox{if }\lambda\hbox{ is even.}\end{cases}$
Combining the above three equations, we obtain the desired equation. This completes the proof. ∎
Appendix G Proof of Lemma 12
Proof.
By definition, we first have
$\sum\limits_{t=q^{k}}^{aq^{k}-1}N(t+1)=\sum\limits_{t=q^{k}}^{aq^{k}-1}t-\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor.$ (73)
Through direct computation, we obtain
$\sum\limits_{t=q^{k}}^{aq^{k}-1}t=\begin{cases}\frac{1}{2}a(a-1)&\hbox{ if }k=0,\\ \frac{1}{2}(a^{2}-1)q^{2k}-\frac{1}{2}(a-1)q^{k}&\hbox{ if }k\geq 1.\end{cases}$ (74)
We now determine the value of $\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor.$ First, note that $\lfloor t/q\rfloor=0$ for all $t\in\left[q^{k},aq^{k}-1\right]$ when $k=0$. This implies that $\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor=0$ if $k=0$. When $k\geq 1,$ the interval $\left[q^{k},q^{k+1}-1\right]$ can be partitioned as
$\left[q^{k},q^{k+1}-1\right]=\bigsqcup\limits_{i=0}^{q^{k-1}(a-1)-1}\left[q^{k}+iq,q^{k}+(i+1)q-1\right].$
Moreover, for each integer $i\in\left[0,q^{k-1}(a-1)-1\right]$ and each integer $t\in\left[q^{k}+iq,q^{k}+(i+1)q-1\right]$, we have $\lfloor t/q\rfloor=q^{k-1}+i$. Therefore,
$\begin{split}\sum\limits_{t=q^{k}}^{aq^{k}-1}\lfloor t/q\rfloor&=\sum\limits_{i=0}^{q^{k-1}(a-1)-1}\sum\limits_{t=q^{k}+iq}^{q^{k}+(i+1)q-1}(q^{k-1}+i)\\ &=\frac{1}{2}(a^{2}-1)q^{2k}-\frac{1}{2}(a-1)q^{k}\end{split}$
if $k\geq 1.$ Combining (73), (74) and the value of $\sum\limits_{t=q^{k}}^{q^{k+1}-1}\lfloor t/q\rfloor$ given as above, we obtain the desired equality. ∎
Appendix H Proof of Lemma 13
Proof.
The arguments used to derive (13) and (14) are analogous, so we only demonstrate equation (13) holds through the following two cases.
Case 1. Suppose that $1\leq i\leq k$. In this case, we first have
$\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor=\frac{t(q^{2i-1}-1)}{\lambda}+\lfloor\frac{2t}{\lambda}\rfloor\quad\hbox{and}\quad\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor=\frac{t(q^{2i-2}-1)}{\lambda}+\lfloor\frac{2t}{\lambda}\rfloor.$
It follows that
$\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor=\frac{tq^{2i-2}(q-1)}{\lambda}.$
This leads to
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]=\sum\limits_{t=q^{k-i}}^{aq^{k-i}-1}\frac{tq^{2i-2}(q-1)}{\lambda}-\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}$ (75)
with $\mathcal{W}_{i}=\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\mid t\}$.
Note that $\mathcal{W}_{i}=\varnothing$ if $i=k$. Thus,
$\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}=0\quad\hbox{if }i=k.$ (76)
On the other hand, it can be easily verified that $\mathcal{W}_{i}=\{q^{k-i}+jq\mid j=0,\ldots,q^{k-i-1}(a-1)-1\}$ if $1\leq i\leq k-1.$ Therefore,
$\begin{split}\sum\limits_{t\in\mathcal{W}_{i}}\frac{tq^{2i-2}(q-1)}{\lambda}&=\sum\limits_{j=0}^{q^{k-i-1}(a-1)-1}\frac{(q^{k-i}+qj)q^{2i-2}(q-1)}{\lambda}\\ &=\frac{1}{2\lambda}(q-1)(a-1)\left[(a+1)q^{2k-3}+q^{k+i-2}\right]\end{split}$ (77)
if $1\leq i\leq k-1$. Additionally, by straightforward computation, we can obtain
[…]
By substituting equations (76)-(78) into equation (75), we conclude that equation (13) holds for $1\leq i\leq k$.
Case 2. Suppose that $k+1\leq i\leq 0$. Then for any integer $t\in\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ with $q$-adic expansion $\sum\limits_{\ell=0}^{k-i}t_{\ell}q^{\ell}$, we have
$\lfloor tq^{2i-1}\rfloor=\lfloor tq^{2i-2}\rfloor\cdot q+t_{-2i+1}.$ (79)
This leads to
$\begin{split}\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor=\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}(q^{\ell}-1)}{\lambda}+\frac{\lfloor tq^{2i-2}\rfloor(q-1)}{\lambda}+\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}}{\lambda}\rfloor.\end{split}$
We also have
$\begin{split}\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor=\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}(q^{\ell}-1)}{\lambda}+\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor.\end{split}$
Combining the above two equalities, we can obtain
$\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor-\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor=\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}}{\lambda}\rfloor-\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor+\frac{\lfloor tq^{2i-2}\rfloor(q-1)}{\lambda}.$
By applying Lemma 8 with $x=\sum\limits_{\ell=1}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}$ and $y=\sum\limits_{\ell=1}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor$, we have
$\sum\limits_{t_{0}=1}^{q-1}\left[\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor+t_{-2i+1}}{\lambda}\rfloor-\lfloor\frac{\sum\limits_{\ell=0}^{k-i}t_{\ell}+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor\right]=\frac{(q-1)t_{-2i+1}}{\lambda}.$
Noticing that the value of $\lfloor tq^{2i-2}\rfloor$ is independent of $t_{0}$, we get
$\sum\limits_{t_{0}=1}^{q-1}\frac{\lfloor tq^{2i-2}\rfloor(q-1)}{\lambda}=\frac{\lfloor tq^{2i-2}\rfloor(q-1)^{2}}{\lambda}.$
Noting that $\lfloor tq^{2i-2}\rfloor=\lfloor\frac{\lfloor tq^{2i-1}\rfloor}{q}\rfloor$ and recalling equation (79), we can add the above two sums to obtain
$\begin{split}\sum\limits_{t_{0}=1}^{q-1}\left[\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor-\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor\right]&=\frac{\lfloor tq^{2i-2}\rfloor(q-1)^{2}}{\lambda}+\frac{(q-1)t_{-2i+1}}{\lambda}\\ &=\frac{q-1}{\lambda}N(\lfloor tq^{2i-1}\rfloor+1).\end{split}$ (80)
In addition, each integer $t\in\left[q^{k-i},aq^{k-i+1}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ with $q$-adic expansion $\sum\limits_{\ell=0}^{k-i}t_{\ell}q^{\ell}$ can be uniquely decomposed as
$t=\lfloor tq^{2i-1}\rfloor\cdot q^{-2i+1}+\sum\limits_{\ell=1}^{-2i}t_{\ell}q^{\ell}+t_{0}.$
Furthermore, as $t$ ranges over integers from $q^{k-i}$ to $q^{k-i+1}-1$ that are not divisible by $q$, the value of $\lfloor tq^{2i-1}\rfloor$ ranges from $q^{k+i-1}$ to $aq^{k+i-1}-1$. Additionally, for each fixed value of $\lfloor tq^{2i-1}\rfloor$, the sum $\sum\limits_{\ell=1}^{-2i}t_{\ell}q^{\ell}$ ranges over integers from 0 to $q^{-2i}-1$, while $t_{0}$ varies from $1$ to q1q-1. Therefore, we can conclude that
$\begin{split}\sum\limits_{\begin{subarray}{c}t=q^{k-i},q\nmid t\end{subarray}}^{q^{k-i+1}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]&=\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}\sum\limits_{\sum_{\ell=1}^{-2i}t_{\ell}q^{\ell}=0}^{q^{-2i}-1}\sum\limits_{t_{0}=1}^{q-1}\left[\lfloor\frac{t+\lfloor tq^{2i-1}\rfloor}{\lambda}\rfloor-\lfloor\frac{t+\lfloor tq^{2i-2}\rfloor}{\lambda}\rfloor\right]\\ &=\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}\frac{q^{-2i}(q-1)}{\lambda}N(\lfloor tq^{2i-1}\rfloor+1),\end{split}$
where the second equality follows from equation (80). We can now apply Lemma 12 to conclude that (13) holds for $k+1\leq i\leq 0$.
By now, we have established equation (13) ∎
Appendix I Proof of Assertion 1
Proof.
Suppose that $m$ is odd. We first aim to show that
$\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}=\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}.$ (81)
For any integer $a\in\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k_{\delta}}a_{\ell}q^{\ell}$, we have
$V(\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell})\leq V(a)\leq V(\lambda(\delta-1)).$
Noting that
$V(\lambda(\delta-1))=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{0})$
and
$V(\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell})=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},\mathbf{0}_{h-k_{\delta}+1}),$
it follows that
$V(a)=(\mathbf{0}_{h-k_{\delta}},a_{h+k_{\delta}},\ldots,a_{0})=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},a_{h-k_{\delta}},\ldots,a_{0}).$ (82)
Recalling the definition of $s_{\delta}$, it follows that $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $a_{h+s_{\delta}}>0$. By Remark 5, this implies $a\not\in\mathcal{A}_{k_{\delta}}(i)$ for any integer $i\neq s_{\delta}$. Therefore, $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(i)=\varnothing$ for any integer $i\neq s_{\delta}.$ Then applying Lemma 3, we can conclude that (81) holds.
Now, let us count the number of integers in the set $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}.$ Recall that $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$, $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}$, and $\alpha(a)=\sum\limits_{\ell=0}^{k_{\delta}+s_{\delta}-1}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{A}_{k_{\delta}}(s_{\delta})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k_{\delta}}a_{\ell}q^{\ell}$. We conclude from equation (82) and the definition of $\mathcal{A}_{k_{\delta}}(s_{\delta})$ that an integer $a\in\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\cap\mathcal{D}_{\lambda}$ if and only if $a$ satisfies the decomposition
$a=w_{\delta}+\alpha(a)$
with $\alpha(a)\in\{\alpha\in\left[1,\mu(\delta)\right]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+w_{\delta}\}.$ Consequently, we have
$\left|\left[\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{A}_{k_{\delta}}\cap\mathcal{D}_{\lambda}\right|=\left|\{\alpha\in\left[1,\mu(\delta)\right]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+w_{\delta}\}\right|.$
Then by applying Lemma 6, equation (21) follows.
Suppose that $m$ is even. The equality in (22) can be obtained by employing a similar argument as above. It remains to establish equation (23). By applying Lemma 5, we can conclude that $a\in\left[\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\lambda(\delta-1)\right]\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $V(a)$ has the form
$(\mathbf{0}_{h-k_{\delta}-1},\delta_{h+k_{\delta}},\ldots,\delta_{h},\mathbf{0}_{h-k_{\delta}-1},\delta_{h+k_{\delta}},\ldots,\delta_{h})$
with $\delta_{h}>0$, $\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$ and $\lambda\mid\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}+\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$. Since $\lambda\mid q-1$, the condition $\lambda\mid\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}+\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ is equivalent $\lambda\mid 2\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}.$ Therefore, an integer satisfying the above condition exists and is unique if and only if $\delta_{h}>0$, $\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\leq\sum\limits_{\ell=0}^{h-1}\delta_{\ell}q^{\ell}$ and $\lambda\mid 2\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$. It follows that equation (23) holds. ∎
Appendix J Proof of Assertion 2
Proof.
Suppose that $m$ is odd. By definition, an integer $a\in\left[q^{h+k_{\delta}},\;\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)$ if and only if the following conditions are satisfied:
  • (i)
    $V(a)$ is of the form given in (9) and satisfies (7) and (8) for $k=k_{\delta}$;
  • (ii)
    $V(q^{h+k_{\delta}})\leq V(a)<V\left(\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)$.
Notice that
$V(q^{h+k_{\delta}})=(\mathbf{0}_{h-k_{\delta}},1,\mathbf{0}_{h+k_{\delta}})\quad\text{and}\quad V\left(\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},\mathbf{0}_{h-k_{\delta}+1}).$
Therefore, the form of $V(a)$ in (9) together with the inequality $a_{0}>0$ in (8) imply that condition (ii) is satisfied if and only if
$(1,\mathbf{0}_{2k_{\delta}-1})\leq(a_{h+k_{\delta}},\ldots,a_{h+i},\mathbf{0}_{k_{\delta}+i-1})<(\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1}).$
Since $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$, this is further equivalent to
$q^{k_{\delta}-i}\leq\sum_{\ell=h+i}^{h+k_{\delta}}a_{\ell}q^{\ell-h-i}<\sum_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h-i}.$
Recall the definition of the set $\mathcal{T}_{i}(\delta)$ and Remarks 13. We now can conclude that for each integer $i\in\left[-k_{\delta}+1,k_{\delta}\right]$, an integer $a\in\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition
$a=t(a)\cdot q^{h+i}+\alpha(a)$
with
$\begin{cases}\lambda\mid t(a)+\alpha(a),\\ t(a)\in\mathcal{T}_{i}(\delta),\\ 1\leq\alpha(a)\leq\lfloor t(a)\cdot q^{2i-1}\rfloor\hbox{ and }q\nmid\alpha(a).\end{cases}$
We may now conclude that, for each integer $i\in[-k_{\delta}+1,\,k_{\delta}]$, an integer
$a\in\left[q^{h+k_{\delta}},\sum_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}$
exists if and only if $a$ admits a decomposition of the form
$a=t(a)\,q^{h+i}+\alpha(a),$
where
$\begin{cases}\lambda\mid t(a)+\alpha(a),\\ t(a)\in\mathcal{T}_{i}(\delta),\\ 1\leq\alpha(a)\leq\lfloor t(a)\,q^{2i-1}\rfloor\quad\text{and}\quad q\nmid\alpha(a).\end{cases}$
Consequently, we can apply Lemma 6 to conclude that
$\begin{split}\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left|\left\{\alpha\in[1,\lfloor tq^{2i-1}\rfloor]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+t\right\}\right|\\ &=\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].\end{split}$
Then applying Lemma 3, it follows that (24) holds.
Suppose that $m$ is even. We can first utilize a similar argument as above to obtain (25). Next, we demonstrate that equation (26) holds. By applying Lemma 5, we can conclude that an integer $a\in\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $a$ can be decomposed as decomposition
$a=\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}\cdot q^{h}+\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}$ (83)
with $q\nmid\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}$, $q^{h+k_{\delta}}\leq a<\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ and $\lambda\mid a$. It is easy to see that $q^{h+k_{\delta}}\leq a<\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ holds if and only if $q^{k_{\delta}}\leq\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1$. Furthermore, since $\lambda\mid q-1$, the condition $\lambda\mid a$ is satisfied if and only if $\lambda\mid 2\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}$. Therefore, an integer $a\in\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition given in (83) with
$\begin{cases}q^{k_{\delta}}\leq\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1;\\ q\nmid\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell};\\ \lambda\mid 2\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}.\end{cases}$
Consequently,
$\begin{split}\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\cap\mathcal{D}_{\lambda}\right|=\left|\left\{\alpha\in\left[q^{k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}-1\right]:\lambda\mid 2\alpha\hbox{ and }q\nmid\alpha\right\}\right|\end{split}$
By applying Lemma 7, it follows that equation (26) holds. ∎
Appendix K Proof of Assertion 3
Proof.
The arguments used to establish (27) and (28) are analogous. Therefore, we only demonstrate that (27) holds when $m$ is odd. Using reasoning similar to that in the proof of Assertion 2, we can conclude that an integer $a\in\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition
$a=t(a)\cdot q^{h+i}+\alpha(a)$
with
$\begin{cases}\lambda\mid t(a)+\alpha(a);\\[4.30554pt] t(a)\in\left[q^{k-i},\,q^{k-i+1}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\};\\[4.30554pt] 1\leq\alpha(a)\leq\left\lfloor t(a)\cdot q^{2i-1}\right\rfloor\text{ and }q\nmid\alpha(a).\end{cases}$
Consequently, by applying Lemma 6, we derive
$\begin{split}\left|\mathcal{A}_{k_{\delta}}(i)\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{t=q^{k-i},q\nmid t}^{q^{k-i+1}-1}\left|\left\{\alpha\in\left[1,\lfloor tq^{2i-1}\rfloor\right]:q\nmid\alpha\hbox{ and }\lambda\mid\alpha+t\right\}\right|\\ &=\sum\limits_{t=q^{k-i},q\nmid t}^{q^{k-i+1}-1}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].\end{split}$
Then applying Lemma 13, we obtain equation (27).
Appendix L Proof of Assertion 4
Proof.
Following a similar approach to the proof of Assertion 3, we can obtain
$|B_{k}(0)\cap\mathcal{D}_{\lambda}|=\sum_{t=q^{k},q\nmid t}^{q^{k+1}-1}\left[\lfloor\frac{2t}{\lambda}\rfloor-\lfloor\frac{\left\lfloor{t}{q}^{-1}\right\rfloor+t}{\lambda}\rfloor\right].$
Then applying Lemmas 10 and $1$, the assertion follows. ∎
Appendix M Proof of Assertion 5
Proof.
By applying Lemma 5, we can conclude that an integer $a\in\mathcal{H}\cap[q^{h+k},q^{h+k+1})\cap\mathcal{D}_{\lambda}$ if and only if $a$ admits the decomposition
$a=\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}\cdot q^{h}+\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}.$ (84)
with $a_{0}>0$, $a_{k}>0$, and $\lambda\mid a$. Since $\lambda\mid q-1$, it follows that $\lambda\mid a$ is equivalent to $\lambda\mid 2\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}$. In addition, it is straightforward to verify that the conditions $a_{0}>0$ and $a_{k}>0$ are satisfied if and only if $q^{k}\leq\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}\leq q^{k+1}-1$ and $q\nmid\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}$. Therefore, an integer $a\in\mathcal{H}\cap[q^{h+k},q^{h+k+1})\cap\mathcal{D}_{\lambda}$ if and only if $a$ has the decomposition as given in (84) with
$\begin{cases}q\nmid\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell};\\ q^{k}\leq\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}\leq q^{k+1}-1;\\ \lambda\mid 2\sum\limits_{\ell=0}^{k}a_{\ell}q^{\ell}.\end{cases}$
Consequently,
$\left|\mathcal{H}\cap\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{D}_{\lambda}\right|=\left|\left\{\alpha\in\left[q^{k},q^{k+1}-1\right]:\lambda\mid 2\alpha\hbox{ and }q\nmid\alpha\right\}\right|.$
Then by applying Lemma 7, the assertion follows. ∎